### v2

# Yong Ji1, Junye Li1 and Rui Yang∗2,3,4

Abstract. Saturated set and its reduced case, the set of generic points, constitute two significant types of fractal-like sets in multifractal analysis of dynamical systems. In the context of infinite entropy systems, this paper aims to give some qualitative aspects of saturated sets and the set of generic points in both topological and measure-theoretic perspectives. For systems with specification property, we establish the certain variational principles for saturated sets in terms of Bowen and packing metric mean dimensions, and show the upper capacity metric mean dimension of saturated sets have full metric mean dimension. All results are useful for understanding the topological structures of dynamical systems with infinite topological entropy.
As applications, we further exhibit some qualitative aspects of metric mean dimensions of level sets and the set of mean Li-Yorke pairs in infinite entropy systems.
Keywords and phrases: Metric mean dimension; Saturated set; Generic point; Variational principle; Multifractal analysis
1 Introduction
Let $(X,d,f)$ be a topological dynamical system (t.d.s. for short), where $(X,d)$ is a compact metric space with a metric $d$, and $f:X\to X$ is a homeomorphism. Let $M(X)$, $M_{f}(X)$, $M_{f}^{e}(x)$ denote the sets of Borel probability measures endowed with weak-topology, $f$-invariant Borel probability measures, and ergodic $f$-invariant Borel probability measures on $X$, respectively. Measure-theoretic entropy and topological entropy are two vital invariants for us to understand the topological complexity of dynamical systems, which are related by the classical variational principle:
$h_{top}(f)=\sup_{\mu\in M_{f}(X)}{h_{\mu}(f)},$
where $h_{top}(f)$ is the topological entropy of $(X,f)$, $h_{\mu}(f)$ is the measure-theoretic entropy of $\mu$, and the supremum ranges over all $f$-invariant measures of $X$. In 1973, analogous to the definition of Hausdorff dimension Bowen [Bow73] introduced Bowen topological entropy for any subset of the phase space. The pioneering work directly contributes to the development of dimension theory of dynamical systems, especially as a powerful tool to study the “size” of fractal-like sets in multifractal analysis.
The multifractal analysis, which can date back to Besicovitch, takes into account the decomposition of the whole phase space into smaller subsets consisting of the points with the similar dynamical behaviors(or known as multifractal spectrum), and then from the geometrical or topological viewpoint describes the “size” of those subsets by using fractal dimensions and dimensional entropies. Such a procedure allows us to get more information about the dynamics and even recovers the partial dynamical behaviors of the system, which is called phenomenon multifractal rigidity. We aim to study the multifractal analysis of saturated sets in infinite entropy systems.
Let us mention some relevant results about the dimensional entropies of saturated sets. Given $x\in X$, by $\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}\in M(X)$, where $\delta_{x}$ is the Dirac mass at $x$, we denote the $n$-th empirical measure of the Birkhoff average w.r.t. $x$. Then the limit-point set $V_{f}(x)$ of $\{\mathcal{E}_{n}(x)\}_{n\geq 1}$ is a nonempty compact connected subset of $M_{f}(X)$ in the sense of weak-topology (cf. [DGS76, Proposition 3.8]). Specially, when $V_{f}(x)=\{\mu\}$ for some measure $\mu$, one says that $x$ is a generic point of $\mu$. The set of all generic points of $\mu$ is denoted by $G_{\mu}$. Then the Birkhoff’s ergodic theorem implies that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic. Bowen [Bow73] proved that for any ergodic measure $\mu$, the Bowen topological entropy of $G_{\mu}$ is equal to $h_{\mu}(f)$, that is,
$h_{top}^{B}(f,G_{\mu})=h_{\mu}(f).$
For non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$ and even easily constructs the example showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$ such that Bowen’s result is not valid. It is natural to expect whether there are some proper conditions such that the equality still holds for non-ergodic measures. A subset $D$ of $X$ is called a saturated set if $x\in D$ and $V_{f}(x)=V_{f}(y)$, then $y\in D$. More generally, given a nonempty closed connected subset $K\subset M_{f}(X)$, we can obtain a saturated set $G_{K}=\{x\in X:V_{f}(x)=K\}$ associated with $K$. For systems with specification property, including hyperbolic systems, Sigmund [Sig72, Sig74] firstly proved the existence of the saturated sets, which was later extended to non-uniformly hyperbolic and non-uniformly expanding systems by several authors [LST13, TV17]. As for the estimation of Bowen topological entropy of saturated sets, for systems with $g$-almost product property that is weaker than specification property, which solely requires a point partially $\epsilon$-shadowing the given finite pieces of orbit-segments, and with uniform separation property (e.g. expansive and more generally asymptotically $h$-expansive systems), Pfister and Sullivan [PS07] established the certain variational principle for any nonempty compact connected set $K\subset M_{f}(X)$,
$h_{top}^{B}(T,G_{K})=\inf\{h_{\mu}(f):\mu\in K\}.$
Specially, only the assumption of $g$-almost product property they [PS07] showed
$h_{top}^{B}(f,G_{\mu})=h_{\mu}(f)$
holds for all $\mu\in M_{f}(X)$, which extends Bowen’s remarkable result to non-ergodic measures. Consequently, the work of [Bow73, PS07] leads to an intensive study toward the estimation of different types of dimensional entropies of saturated sets. Packing entropy of non-compact subsets was introduced by Feng and Huang [FH12] in a way resembling the definition of packing dimension. Additionally, under the (stronger) conditions of specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] showed the following variational principle for any nonempty compact connected set $K\subset M_{f}(X)$,
$h_{top}^{P}(f,G_{K})=\sup\{h_{\mu}(f):\mu\in K\},$
where $h_{top}^{P}(f,K)$ is the packing entropy of $K$. Afterwards, weakening the specification property, for systems with $g$-almost product property, Hou, Tian and Zhang [HTZ23] not only proved that the variational principle $h_{top}^{P}(f,G_{K}^{C})=\sup\{h_{\mu}(f):\mu\in K\}$ holds for (mixed) saturated sets with more complicated dynamics, but also showed the upper capacity topological entropy of the saturated sets have full topological entropy, i.e.
$h_{top}^{UC}(f,G_{K}^{C})=h_{top}(f),$
where $G_{K}^{C}:=G_{K}\cap\{x\in X:C_{f}(X)\subset\omega_{f}(x)\},$ $\omega_{f}(x):=\bigcap_{n=0}^{\infty}\overline{\bigcup_{k=n}^{\infty}\{f^{k}x\}}$ is the $\omega$-limit set of $x$, and $C_{f}(X)$ is the measure center given by $C_{f}(X):=\overline{\cup_{\mu\in M_{f}(X)}{\rm supp}{(\mu)}}$, ${\rm supp}(\mu)$ is the supported set of $\mu$. Furthermore, for systems satisfying $g$-almost product property and uniform separation property, Huang, Tian and Wang [HTW19] showed Pfister and Sullivan’s variational principle is valid for transitive-saturated set $G_{K}^{T}:=G_{K}\cap{\rm Tran}$, where ${\rm Tran}$ denotes the set of transitive points. Readers can turn to [JCZ22, LLS17] for more results involving the dimensional entropies of saturated sets in the setting of nonuniformly hyperbolic systems, and [ZC18, DZZ23, Z18, RTZ23, ZRZ23] for some results concerning the amenable Bowen and packing topological entropies of saturated sets in the context of the actions of amenable groups.
According to the possible values of topological entropy, dynamical systems can be divided into three types: zero entropy systems, finite positive entropy systems and infinite entropy systems. Infinite entropy system always means the system may possess quiet complicated topological complexity, so the notion of topological entropy no longer provides more information about the dynamics of the system. In 2000, Lindenstrauss and Weiss [LW00] introduced a refined notion for topological entropy, called metric mean dimension, which is a dynamical analogue of box dimension to measure how fast the $\epsilon$-topological entropy diverges to $\infty$ as $\epsilon$ goes to 00. It turns out that any finite entropy system has zero metric mean dimension. See some applications of metric mean dimension in the analog compression of information theory [GS20], and in estimating the upper bound of mean dimension of some complex systems [Tsu18a, Tsu18b]. The notion of measure-theoretic counterpart of metric mean dimension, called upper and lower mean Rényi information dimensions, is due to Gutman and Śpiewak[GS21]. More precisely, the lower and upper mean Rényi information dimensions of $\mu\in M_{f}(X)$ are respectively defined by
$\underline{\rm MRID}_{\mu}(f,X,d):$ $=\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
$\overline{\rm MRID}_{\mu}(f,X,d):$ $=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
which can be explained how fast the Kolmogorov-Sinai $\epsilon$-entropy $\inf_{\mathrm{diam}\xi<\epsilon}\limits h_{\mu}(f,\xi)$ diverges to $\infty$ as $\epsilon\to 0$. Some variational principles of metric mean dimension of $\mathbb{Z}$-actions can be found in [LW00, GS21, S22, YCZ23a]. Then the aforementioned work suggests the following questions for saturated sets in infinite entropy systems:
Questions:
  • (1)
    For ergodic measures, do the metric mean dimensions of the set of generic points equal to the mean Rényi information dimensions?
  • (2)
    For non-ergodic measures, under the assumption of specification-like property, are there some certain variational principles for metric mean dimensions of saturated sets? If not, can metric mean dimensions of saturated sets have the full metric mean dimension?
For (1), the authors in [YCZ24] showed the packing metric mean dimension of the set of generic points equals to the mean Rényi information dimension, however it holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula in infinite entropy systems (cf.[YCZ24, §5, Question,(1)]); for (2), noting the uniform separation property implies the system has finite topological entropy, so extending the work of [PS07] to the framework of infinite entropy systems encounters some nontrivial challenges. Only the assumption of $g$-almost product property, Yuan [Y23] investigated the Bowen metric mean dimension of $G_{K}$, and showed that for any $\{\mu_{1},\mu_{2},...,\mu_{k}\}\subset M_{f}(X)$ and any non-empty compact connected set $K$ contained in the convex closure ${\rm cov}\{\mu_{1},...,\mu_{k}\}$,
$\overline{\rm mdim}_{M}^{B}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Prior to the comparison of different types of specification-like property, we only focus on the systems with specification property and give a full answer to question (2). Actually, since the asymptotic behavior of orbits ${\rm orb}_{f}(x)=\{f^{n}(x):n\geq 0\}$ is a fundamental topic in ergodic theory, the following statements of main results involve more general (mixed) saturated sets. For simplicity, we write
$G_{K}^{\omega,U}:=G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},$
where $U\subset X$ is a nonempty open subset, $K\subset M_{f}(X)$ is a nonempty closed connected (or convex) subset. If $G_{K}^{\omega,U}\not=\emptyset$, then there exists a transitive point in the open set $U$ such that $V_{f}(x)=K$.
Saturated sets exhibit different dynamical behaviors in terms of different types of metric mean dimensions. There exist certain variational principles for Bowen and packing metric metric mean dimensions of saturated sets (Theorems 1.1 and 1.2), while the upper capacity metric mean dimension of saturated set has full metric mean dimension (Theorem 1.4).
Theorem 1.1.
Let $(X,d,f)$ be a dynamical system with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, and $U\subset X$ is a nonempty open set, then
$\overline{\rm mdim}_{M}^{P}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{P}\left(f,G_{K}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\},d\right)$
$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where $\overline{\rm mdim}_{M}^{P}(f,Z,d)$ is the packing upper metric mean dimension of $Z$, $h_{\mu}(f,\xi)$ is the measure-theoretic entropy of $\mu$ w.r.t. $\xi$, and the infimum is taken over all finite Borel measurable partitions of $X$.
Theorem 1.2.
Let $(X,d,f)$ be a dynamical system with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact connected set, and $U\subset X$ is a nonempty open set, then
$\overline{\rm mdim}_{M}^{B}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{B}\left(f,G_{K}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\},d\right)$
$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where $\overline{\rm mdim}_{M}^{B}(f,Z,d)$ is the Bowen upper metric mean dimension of $Z$.
Maybe, more satisfactory results are exchanging the order of $\limsup_{\epsilon\to 0}$ and $\sup_{\mu\in K}$ (or $\inf_{\mu\in K}$), but until now the authors do not know whether there are some counter-examples of infinite entropy systems with specification property such that the following equalities are strict:
$\sup_{\mu\in K}\left\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\right\}<\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d),$
and
$\inf_{\mu\in K}\left\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\right\}<\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d).$
Inspired by the work of [GS21, CPV24], we define two variants of mean Rényi information dimension of invariant measures. For any closed set $K$ of $M_{f}(X)$ and $\mu\in K$, by $M_{K}(\mu)$ we denote the space of sequences of Borel probability measures in $K$ which converge to $\mu$ in the weak-topology. We respectively define the (modified) lower and upper mean Rényi information dimensions of $\mu\in K$ as
$\underline{\rm mrid}_{\mu}(f,K,d):$ $=\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi),$
$\overline{\rm mrid}_{\mu}(f,K,d):$ $=\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
Then we have the following desired variational principles in terms of the modified lower and upper mean Rényi information dimensions.
Theorem 1.3.
Let $(X,d,f)$ be a dynamical system with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, and $U\subset X$ is a nonempty open set, then
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\},$
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
We remark that Theorem 1.3 does not require the connectedness of $K$ since any convex set in a topological vector space is automatically connected.
Theorem 1.4.
Let $(X,d,f)$ be a dynamical system with specification property. Let $K\subseteq M_{f}(X)$ be a nonempty compact connected set, and $U\subset X$ be a nonempty open set, then
$\overline{\rm mdim}_{M}^{UC}\left(f,G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},d\right)=\overline{\rm mdim}_{M}^{UC}\left(f,G_{K},d\right)=\overline{\rm mdim}_{M}(f,X,d),$
where $\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ is the upper capacity metric mean dimension of $Z$ defined by separated sets.
As a direct consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case is $K$ being exactly a single point. The Corollary 1.5 presents the metric mean dimensions of generic points of any invariant measures for systems with specification property, which extends the result given in [YCZ24, Theorem 1.1] to all non-ergodic measures.
Corollary 1.5.
Let $(X,d,f)$ be a dynamical system with specification property. Then for any $\mu\in M_{f}(X)$,
$\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}^{P}_{M}(f,G_{\mu},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi),$
$\overline{\rm mdim}_{M}^{UC}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
Consequently, if $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure such that $\overline{\rm mdim}_{M}(f,X,d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi)$, then we have
$\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}^{P}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}^{UC}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
The “Consequently” part can be realized for some dynamical systems. For example, the full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$ with the compatible product metric $d(x,y)=\sum_{n\in\mathbb{Z}}\frac{||x_{n}-y_{n}||_{\mathbb{R}^{m}}}{2^{|n|}}$, and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ is product measure of the Lebesgue measure $\mathcal{L}$ on $[0,1]^{m}$. The concept of maximal metric mean dimension measure was proposed in [YCZ23a], yet for what kind of infinite entropy systems can have maximal metric mean dimension measure is still a difficult problem. See [LR24] for the existence of maximal metric mean dimension measure for typical conservative homeomorphisms. By Gutman- Śpiewak variational principle (see Theorem 2.4) and Corollary 1.5 we have following (abstract) criterion for maximal metric mean dimension measure for systems with specification property.
Corollary 1.6.
Let $(X,d,f)$ be a dynamical system with specification property. An invariant measure $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure such that $\overline{\rm mdim}_{M}(f,X,d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi)$ if and only if $\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
Some comments are in order.
  • (1)
    There are plenty of infinite entropy systems with specification property, for instance, the usual full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$, and the examples presented in [BR23, Section IV] such that our main results can be applied to such systems.
  • (2)
    In Theorem 1.3, we do not know whether the equalities hold for Bowen upper metric mean dimension and packing lower metric mean dimension. The other aforementioned theorems are valid for the corresponding lower metric mean dimensions by considering changing the $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.
  • (3)
    We believe that it is trivial to extend the aforementioned results to metric mean dimensions with potential by considering some modifications of the present proofs. The novelty of our work not only presents a proper and full extension of the previous work in [PS07, ZCC12, HTW19] to the setting of infinite entropy systems of $\mathbb{Z}$-actions with specification property, but also may be helpful for the multifractal analysis of other types of fractal-like sets in infinite entropy systems. It remains unclear whether there are some proper substitutions of the uniform separation property so that all results are true for $g$-almost product property, or beyond $\mathbb{Z}$-actions.
The paper is organized as follows. In section 2, we recall the precise definitions of metric mean dimensions and the associated properties, and two types of measure-theoretic $\epsilon$-entropies, and the specification property. In section 3, we prove the main results. In section 4, we give two applications of the main results. In section 5, we show the metric mean dimensions of generic points of full shift over compact metric spaces.
2 Preliminaries
In this section, we recall the concepts of metric mean dimensions in both topological and measure-theoretic situations, and state the precise definition of specification property in dynamical systems.
2.1 Three types of metric mean dimensions
In this subsection, we recall the definitions of upper capacity metric mean dimension on subsets[LW00], Bowen metric mean dimension[LV21, W21] and packing metric mean dimension on subsets [YCZ22].
Let $(X,d,f)$ be a dynamical system. For $n\in\mathbb{N}$, we define the Bowen metric $d_{n}$ on $X$ as
$d_{n}(x,y):=\max_{0\leq i\leq n-1}\{d(f^{i}x,f^{i}y)\}$
for any $x,y\in X$. Then the Bowen open ball and closed ball of $x$ in the metric $d_{n}$ with radius $\epsilon$ are respectively given by
$B_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)<\epsilon\},$
$\overline{B}_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)\leq\epsilon\}.$
Let $Z$ be a nonempty subset of $X$. Fix $\epsilon>0$. A set $E\subset Z$ is said to be an $(n,\epsilon)$-separated set of $Z$ if $x,y\in E$ with $x\neq y$ implies $d_{n}(x,y)>\epsilon$. Let $s_{n}(Z,\epsilon)$ denote the largest cardinality of $(n,\epsilon)$-separated sets of $Z$. The $\epsilon$-upper capacity topological entropy of $Z$ is given by
$h_{top}^{UC}(f,Z,\epsilon)=\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(Z,\epsilon).$
Analogous to the definition of box dimension in fractal geometry, we can formulate the definition for upper capacity metric mean dimension of subsets using separated sets.
Definition 2.1.
We define the upper capacity metric mean dimension of $Z$ as
$\overline{\rm mdim}_{M}^{UC}(f,Z,d):=\limsup_{\epsilon\to 0}\frac{h_{top}^{UC}(f,Z,\epsilon)}{|\log\epsilon|}.$ (2.1)
When $Z=X$, we write $h_{top}(f,X,\epsilon)=h_{top}^{UC}(f,X,\epsilon)$, and $\overline{\rm mdim}_{M}(f,X,d)=\overline{\rm mdim}_{M}^{UC}(f,X,d)$, respectively. In this case, it is exactly reduced to Lindenstrauss and Weiss’s classical notion introduced in [LW00]. Such a formulation allows us to extend the metric mean dimension to any subsets of the phase space, which is not necessarily compact or $f$-invariant.
Recall that the upper capacity topological entropy of $Z$ is limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 00, that is,
$h_{top}^{UC}(f,Z)=\lim_{\epsilon\to 0}h_{top}^{UC}(f,Z,\epsilon)=\sup_{\epsilon>0}h_{top}^{UC}(f,Z,\epsilon).$
Especially, for any system admitting the infinite topological entropy, topological entropy does not provide the additional information of the dynamics except we have known the system has pretty complicated topological complexity. Thus, (upper capacity) metric mean dimension can be viewed as a quantity to detect how fact the $\epsilon$-upper capacity topological entropy diverges to the infinite (upper capacity) topological entropy as $\epsilon$ tends to 00. More precisely, for sufficiently $\epsilon>0$, we may think of
$h_{top}^{UC}(f,Z,\epsilon)\approx\overline{\rm mdim}_{M}^{UC}(f,Z,d)\cdot|\log\epsilon|.$
It is easy to see that any finite entropy system has zero metric mean dimension, and hence verifies that metric mean dimension is a good candidate to capture the dynamics of infinite entropy systems.
Similar to the definitions of Hausdorff dimension and packing dimension in fractal geometry, as an analogue of the classical Bowen topological entropy[Bow73] and packing topological entropy[FH12] in dimension theory, which is defined by Carathéodory-Pesin structures, the corresponding dimensional characterizations of metric mean dimensions via Carathéodory-Pesin structures are Bowen and packing metric mean dimensions, which has been introduced in [LV21, YCZ22, W21] and studied by several authors in [BR23, CLS21, LL24, FKO24, YCZ24].
Definition 2.2.
For $Z\subset X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
$M(Z,s,N,\epsilon)=\inf\left\{\sum_{i}\exp(-n_{i}s)\right\},$
where the infimum is taken over all finite or countable families $\{B_{n_{i}}(x_{i},\epsilon)\}$ such that $Z\subset\cup_{i}B_{n_{i}}(x_{i},\epsilon)$ with $x_{i}\in X,n_{i}\geq N$.
The quantity $M(Z,s,N,\epsilon)$ is non-decreasing as $N$ increases. So we set
$M(Z,s,\epsilon)=\lim_{N\to\infty}M(Z,s,N,\epsilon).$
There is a critical value of parameter $s$, which we call $\epsilon$-Bowen topological entropy, such that $M(Z,s,\epsilon)$ jumps from $\infty$ to 00, that is, the critical value is given by
$h_{top}^{B}(f,Z,\epsilon)$ $=\inf\{s>0:M(Z,s,\epsilon)=0\}$
$=\sup\{s>0:M(Z,s,\epsilon)=\infty\}.$
We define the Bowen upper metric mean dimension of $Z$ as
$\overline{\rm mdim}^{B}_{M}(f,Z,d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{B}(f,Z,\epsilon)}{|\log\epsilon|}.$
Definition 2.3.
For $Z\subseteq X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
$P(Z,s,N,\epsilon)=\sup\left\{\sum_{i}\exp(-sn_{i})\right\},$
where the supremum is taken over all finite or countable pairwise disjoint families $\{\overline{B}_{n_{i}}(x_{i},\epsilon)\}$ with $x_{i}\in Z$, $n_{i}\geq N$ for all $i$.
Since the quantity $P(Z,s,N,\epsilon)$ is non-increasing as $N$ increases, and thus the following limit exists:
$P(Z,s,\epsilon)=\lim_{N\to\infty}P(Z,s,N,\epsilon).$
Observe that the quantity $P(Z,s,\epsilon)$ does not satisfy the countable additive property with respect to $Z$. So some modifications are needed to obtain the property:
$\mathcal{P}(Z,s,\epsilon):=\inf\left\{\sum_{i=1}^{\infty}P(Z_{i},s,\epsilon):\bigcup_{i=1}^{\infty}Z_{i}\supseteq Z\right\}.$
There is a critical value of parameter $s$, which we call $\epsilon$-packing topological entropy, such that $\mathcal{P}(Z,s,\epsilon)$ jumps from $\infty$ to 00, that is, the critical value is given by
$h_{top}^{P}(f,Z,\epsilon)$ $=\inf\{s>0:\mathcal{P}(Z,s,\epsilon)=0\}$
$=\sup\{s>0:\mathcal{P}(Z,s,\epsilon)=\infty\}.$
We define the packing upper metric mean dimension of $Z$ as
$\overline{\rm mdim}_{M}^{P}(f,Z,d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{P}(f,Z,\epsilon)}{|\log\epsilon|}.$
One can replace $\limsup_{\epsilon\to 0}$ by $\liminf_{\epsilon\to 0}$ in above definitions to obtain the corresponding three types of lower metric mean dimensions on subsets, which are denoted by $\underline{\rm mdim}_{M}^{S}(f,Z,d)$ with $S\in\{UC,B,P\}$. If $\underline{\rm mdim}_{M}^{S}(f,Z,d)=\overline{\rm mdim}_{M}^{S}(f,Z,d)$, we call the common value ${\rm mdim}_{M}^{S}(f,Z,d)$ the (capacity/Bowen/packing) metric mean dimension of $Z$. We summarize some fundamental properties of metric mean dimensions derived in [YCZ22, Proposition 3.4] and [YCZ23b, Proposition 2.6].
Proposition 2.1.
$(1)$ If $Z_{1}\subset Z_{2}$, then
$\overline{\rm mdim}_{M}^{UC}(f,Z_{1},d)$ $\leq\overline{\rm mdim}_{M}^{UC}(f,Z_{2},d)$
$\overline{\rm mdim}^{B}_{M}(f,Z_{1},d)$ $\leq\overline{\rm mdim}^{B}_{M}(f,Z_{2},d)$
$\overline{\rm mdim}^{P}_{M}(f,Z_{1},d)$ $\leq\overline{\rm mdim}^{P}_{M}(f,Z_{2},d).$
$(2)$ If $Z=\bigcup_{i=1}^{\infty}Z_{i}$, then for every $\epsilon>0$,
$h_{top}^{B}(f,Z,\epsilon)$ $=\sup_{i\geq 1}h_{top}^{B}(f,Z_{i},\epsilon)$
$h_{top}^{P}(f,Z,\epsilon)$ $=\sup_{i\geq 1}h_{top}^{P}(f,Z_{i},\epsilon),$
and $h_{top}^{UC}(f,Z,\epsilon)=\max_{1\leq i\leq N}h_{top}^{UC}(f,Z_{i},\epsilon)$ if $Z=\bigcup_{i=1}^{N}Z_{i}$.
$(3)$ If $Z=\bigcup_{i=1}^{N}Z_{i}$, then
$\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}_{M}^{UC}(f,Z_{i},d)$
$\overline{\rm mdim}^{B}_{M}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}^{B}_{M}(f,Z_{i},d)$
$\overline{\rm mdim}^{P}_{M}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}^{P}_{M}(f,Z_{i},d).$
$(3)$ If $Z$ is a nonempty subset of $X$, then for every $10<\epsilon<$ 1, we have
$h_{top}^{B}(f,Z,3\epsilon)\leq h_{top}^{P}(f,Z,\epsilon)\leq h_{top}^{UC}(f,Z,\epsilon).$
Consequently, $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)\leq\overline{\rm mdim}_{M}^{UC}(f,Z,d)$; if $Z$ is $f$-invariant and compact, then the equalities hold.
The metric mean dimension of a given subset is in general not easy to calculate its precise value. However, the Entropy Distribution Principle allows us to get a lower bound for $\epsilon$-Bowen and $\epsilon$-packing topological entropies of some fractal-like sets in multifractal analysis of dynamical systems. Consequently, the powerful tool gives the lower bounds of entropy and metric mean dimension of the fractal-like sets.
Lemma 2.1.
[BR23, Lemma 13] Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Let $Z$ be a Borel set of $X$ and $s\geq 0$. If there exist a Borel probability measure $\mu\in M(X)$ and a constant $C>0$ satisfying:
$(1)$ for every $n$, $\mu(B_{n}(x,\epsilon))\leq Ce^{-ns}$ for each Bowen ball $B_{n}(x,\epsilon)$ such that $B_{n}(x,\epsilon)\cap Z\neq\emptyset$;
$(2)$ $\mu(Z)>0$,
then $h_{top}^{B}(f,Z,\epsilon)\geq s$.
$\epsilon$-packing topological entropy also admits an analogous Entropy Distribution Principle as follows:
Lemma 2.2.
Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Let $Z$ be a Borel set of $X$ and $s\geq 0$. If there exist a Borel probability measure $\mu\in M(X)$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ of positive integer that converges to $\infty$ satisfying:
$(1)$ $\mu(B_{n_{i}}(x,\epsilon))\leq Ce^{-n_{i}s}$ for each $i$ and for any $x\in Z$;
$(2)$ $\mu(Z)>0$,
then $h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s$.
Proof.
For sufficiently small $\delta>0$, by definition it is enough to prove that
$P\left(E,s-\delta,\frac{\epsilon}{6}\right)=\infty$
holds for any Borel subset $E\subset Z$ with $\mu(E)>0$. For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of $E$. Then by $5r$-covering lemma stated in [Mat95, Theorem 2.1], there is a countable pairwise disjoint subcollection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, and in particular the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint, such that
$E\subset\bigcup_{j}B_{n_{i}}(x_{j},\epsilon).$
Hence by condition $(1)$, one has
$P\left(E,s-\delta,n_{i},\frac{\epsilon}{6}\right)\geq\sum_{j}\exp[-n_{i}(s-\delta)]\geq\frac{e^{n_{i}\delta}}{C}\sum_{j\in J}\mu(B_{n_{i}}(x_{j},\epsilon))\geq\frac{e^{n_{i}\delta}}{C}\mu(E).$
Letting $i\to\infty$, we have $P(E,s-\delta,\frac{\epsilon}{6})=\infty$. ∎
2.2 Measure-theoretic $\epsilon$-entropy
In this subsection, we recall the definitions of Kolmogorov-Sinai $\epsilon$-entropy and Katok $\epsilon$-entropy of invariant measures.
Let $C(X)$ denote the space of continuous functions on $X$ equipped with the supermum norm. The Riesz representation theorem on compact metric space gives us an approach to endow with the weak-topology on $M(X)$, which is metrizable by the following metric
$D(\mu,\nu):=\sum_{k=1}^{\infty}\frac{|\int\phi_{k}\mathop{}\!\mathrm{d}\mu-\int\phi_{k}\mathop{}\!\mathrm{d}\nu|}{2^{k}},$
where $\{\phi_{k}\}$ is a family of countable dense subset of $C(X)$ taking values in $[0,1]$. Thus, the diameter of $M(X)$ in sense of metric $D$ is always less than $1$, and for fixed $\mu$ and $\nu$, both $D(\mu,\cdot)$ and $D(\cdot,\nu)$ are convex functions on $M(X)$. It follows from the weak-topology on $M(X)$ that the natural projection from $X$ to $M(X)$, given by $x\mapsto\delta_{x}$, is continuous. Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in sense of weak-topology.
Given an invariant measure $\mu\in M_{f}(X)$, a finite family $\xi=\left\{C_{1},\cdots,C_{k}\right\}$ of Borel subsets of $X$ is a measurable partition of $X$ if $\mu(C_{i}\bigcap C_{j})=0$ for all $i\neq j$ and $\mu(X\setminus\bigcup_{i=1}^{k}C_{i})=0$. The partition entropy of $\xi$ w.r.t. $\mu$ is defined by
$H_{\mu}(\xi)=-\sum_{i=1}^{k}\mu(C_{i})\log(\mu(C_{i})).$
Then we define measure-theoretic entropy of $f$ w.r.t. $\xi$ as
$h_{\mu}(f,\xi)=\lim_{n\to+\infty}\frac{1}{n}H_{\mu}(\xi^{n}),$
where $\xi^{n}:=\vee_{i=0}^{n-1}f^{-i}\xi$ is the $n$-th join of the $n$ partitions $\xi,f^{-1}\xi,\ldots,f^{-(n-1)}\xi$ on $X$. The measure-theoretic entropy of $f$ w.r.t. $\mu$ is defined as
$h_{\mu}(f)=\sup_{\xi}h_{\mu}(f,\xi),$
where the supremum is taken over all finite Borel measurable partitions $\xi$ of $X$.
For topological dynamical systems, the measure-theoretic counterpart of metric mean dimensions on subsets is the measure-theoretic metric mean dimensions of invariant measures. Unlike topological entropy, measure-theoretic entropy of ergodic measures admit several equivalent definitions by considering the limits of the measure-theoretic $\epsilon$-entropies: for instance, Brin-Katok $\epsilon$-entropy, Katok $\epsilon$-entropy, Kolmogorov-Sinai $\epsilon$-entropy, and the others. A systemic treatment about the measure-theoretic $\epsilon$-entropies is stated in [YCZ24], and the divergent rates of different types of measure-theoretic $\epsilon$-entropies coincide in terms of ergodic measures (cf. [YCZ24, Theorem 1.1]). Here, we only recall two types of them.
  • (a)
    The Kolmogorov-Sinai $\epsilon$-entropy (cf. [GS21]) of $\mu$ is given by
    $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
  • (b)
    For $\mu\in M_{f}(X)$, $\epsilon>0$, $\delta\in(0,1)$ and $n\in\mathbb{N}$, we put
    $N_{\mu}(n,\epsilon,\delta):=\min\{\#E:E\subset X~{}\emph{and}~{}\mu(\cup_{x\in E}B_{n}(x,\epsilon))>1-\delta\}.$
    The (lower) Katok $\epsilon$-entropy of $\mu$ is given by
    $\underline{h}_{\mu}^{K}(f,\epsilon,\delta)=\liminf_{n\to\infty}\frac{1}{n}\log N_{\mu}(n,\epsilon,\delta).$
The relation between Katok $\epsilon$-entropy and Kolmogorov-Sinai $\epsilon$-entropy is related by the following inequalities:
Lemma 2.3.
[S22, Lemma 4.1] Let $\mu\in M_{f}^{e}(X)$ and $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})<4\epsilon$ and ${\rm Leb}(\mathcal{U})\geq\epsilon$. Then for any $\delta\in(0,1)$, one has
$\underline{h}_{\mu}^{K}(f,4\epsilon,\delta)\leq\inf_{\xi\succ\mathcal{U}}h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta).$
Consequently, $\underline{h}_{\mu}^{K}(f,16\epsilon,\delta)\leq\inf_{{\rm diam}(\xi)<4\epsilon}\limits h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta).$
The next variational principles verify the relations between the metric mean dimension and the two types of measure-theoretic $\epsilon$-entropies.
Theorem 2.4.
(cf.[GS21, Theorem 3.1, Remark 3.6], [S22, Proposition 7.3]) Let $(X,f)$ be a t.d.s. with a metric $d$. Then for any $\delta\in(0,1)$,
$\overline{\rm mdim}_{M}(f,X,d)$ $=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(X)}F(\mu,\epsilon),$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}^{e}(X)}F(\mu,\epsilon),$
where $F(\mu,\epsilon)\in\{\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi),\underline{h}^{K}_{\mu}(f,\epsilon,\delta)\}$.
A nice survey involving the variational principles of metric mean dimension in terms of different types of measure-theoretic $\epsilon$-entropies [LW00, GS21, S22] is given in [YCZ24, Theorems 1.2 and 1.3].
2.3 Specification property
Sometimes, it is difficult to find a real orbit with desired dynamical behavior, while one can get an approximate orbit of a point that $\epsilon$-shadows any specified finite orbit-segments using specification-like property.
Definition 2.4 (Specification property).
One says that a t.d.s. $(X,f)$ has specification property if for any $\epsilon>0$, there is an integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists $x\in X$ such that
$d(f^{p+a_{j}}x,f^{p}x_{j})<\epsilon$
for any $p=0,\cdots,b_{j}-a_{j}$, $j=1,\cdots,k$.
In other words, that point $x$ $\epsilon$-shadows the $k$ pieces of orbits
$\{x_{j},f(x_{j}),\ldots,f^{b_{j}-a_{j}}(x_{j})\},\ j=1,\ldots,k.$
For instance, every topological mixing locally maximal hyperbolic set, and the full shift on $[0,1]^{\mathbb{Z}}$ have specification property. More examples of infinite entropy systems with specification property can be found in [BR23, Section IV].
A point $x\in X$ is almost periodic if for any open neighborhood $U$ of $x$, there exists $N\in\mathbb{N}$ such that for every $n\in\mathbb{N}$, it holds that $f^{k}(x)\in U$ for some $n\leq k\leq n+N$. Let $AP(X)$ denote the set of all almost periodic points of $X$.
The following result, which was proved for systems with $g$-almost product property [HTZ23, Proposition 2.11], gives a characterization of the structure of $AP(X)$.
Proposition 2.2.
If the dynamical system $(X,f)$ has specification property, then the almost periodic set AP(X) is dense in $X$.
3 Proofs of main results
This section aims to prove the main results of this paper. Essentially, the difficult part is the estimation of lower bound of $\overline{\rm mdim}_{M}^{S}(f,G_{K}^{\omega,U},d)$, where $S\in\{B,P,UC\}$. We shall construct a Moran fractal set containing in $G_{K}^{\omega,U}$, which is a limit set with geometric construction. Then such set is suitable for the purpose of the computation of the lower bound of metric mean dimensions of saturated sets.
3.1 Packing metric mean dimension of saturated sets
We first prove Theorem 1.1. To arrive this goal, we divide the proof of the Theorem 1.1 into two parts.
3.1.1 Upper bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K},d)$
We adapt a general method presented in [PS07, Theorem 4.1], [ZCC12, Theorem 1.1, $(2)$] and [JCZ22, Proposition 4] to obtain the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K})$.
Lemma 3.1.
[Y23, Lemma 3.1] Let $(X,f)$ be a dynamical system and $\epsilon>0$. Let $\{F_{n}\}$ be a sequence of non-empty subsets of $X$, and let $E_{n}$ be an $(n,\epsilon)$-separated subset of $F_{n}$ with the largest cardinality. We define
$v_{n}:=\frac{1}{\#E_{n}}\sum_{x\in E_{n}}\mathcal{E}_{n}(x).$
Then for any limit point $\mu$ of the sequence $\{\nu_{n}\}$, one has
$\limsup_{n\to\infty}\frac{1}{n}\log\#E_{n}\leq\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
This essentially follows the proof of the upper bound of the classical variational principle (cf.[Wal82, Theorem 9.10]). ∎
For $K\subset M_{f}(X)$, we set $G^{K}:=\{x\in X:V_{f}(x)\subset K\}$. Obviously, we have the fact $G_{K}\subset G^{K}$.
Lemma 3.2.
Let $(X,f)$ be a t.d.s..
$(1)$ Let $K\subset M_{f}(X)$ be a nonempty closed convex subset. Then for any $\epsilon>0$,
$h_{top}^{P}(f,G_{K},\epsilon)\leq h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
$(2)$ If $\mu\in M_{f}(X)$, then for any $\epsilon>0$,
$h_{top}^{P}(f,G_{\mu},\epsilon)\leq\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
It suffices to show $(1)$, since one can take $K=\{\mu\}$ as in $(1)$ and use the fact $G_{\mu}=G^{\{\mu\}}$ to get the desired $(2)$.
For $n\in\mathbb{N}$, $\delta>0$, we put
$\mathcal{R}_{K}(\delta,k):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)~{}\text{for infinitely many }n>k\right\},$
$\mathcal{Q}_{K}(\delta,n):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)\right\},$
and $\Theta(K,\epsilon):=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\log S(K,\delta,n,\epsilon),$ where $S(K,\delta,n,\epsilon)$ is the largest cardinality of $(n,\epsilon)$-separated sets of $\mathcal{Q}_{K}(\delta,n)$. Fix $x\in G^{K}$. Then, we can choose a subsequence $\{n_{i}\}$ such that $\mathcal{E}_{n_{i}}(x)\to\mu$ for some $\mu\in K$. This implies for every $\delta>0$, there exists a positive integer $k$ such that $\mathcal{E}_{n_{i}}(x)\in B(K,\delta)$ for sufficiently large $n_{i}$ greater than $k$. Then by the relation $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}_{K}(\delta,k)$ holds for every $\delta$, it follows from Proposition 2.1 that
$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{k}h_{top}^{P}(f,\mathcal{R}_{K}(\delta,k),\epsilon)\leq\sup_{k}h_{top}^{UC}(f,\mathcal{R}_{K}(\delta,k),\epsilon),$
which implies that $h_{top}^{P}(f,G^{K},\epsilon)\leq\Theta(K,\epsilon).$
Next we shall prove for any $\gamma>0$, there is $\mu\in K$ such that
$\Theta(K,\epsilon)-\gamma\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Take a strictly decreasing positive sequence $\{\delta_{k}\}$ with $\delta_{k}\to 0$ and a strictly increasing sequence $\{n_{k}\}$ with $n_{k}\to\infty$ such that
$S(K,\delta_{k},n_{k},\epsilon)>\exp[n_{k}(\Theta(K,\epsilon)-\gamma)].$
Let $F_{k}$ be a maximal $(n_{k},\epsilon)$-separated subset of $\mathcal{R}_{K}(\delta_{k},n_{k})$, and set
$\mu_{k}:=\frac{1}{\#F_{k}}\sum_{x\in F_{k}}\mathcal{E}_{n_{k}}(x)\in B(K,\delta_{k}).$
Choose a subsequence $k_{j}$ such that $\mu_{k_{j}}\to\mu$ as $j\to\infty$. Then $\mu\in K$ since $K$ is a closed convex set. By Lemma 3.1, we have
$\Theta(K,\epsilon)-\gamma\leq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log S(K,\delta_{k_{j}},n_{k_{j}},\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
This completes the proof. ∎
Thus, by the Lemma 3.2 we obtain an upper bound of the packing metric mean dimension of $G_{K}$:
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{P}(f,G_{K},\epsilon)}{|\log\epsilon|}\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.1)
3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
For any small enough $\epsilon^{\ast}>0$, we shall prove that
$\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$ (3.2)
One may assume $H:=\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)>0$. Let $\zeta>0$. Then take $\mu^{\ast}\in K$ such that
$\inf_{\mathrm{diam}(\xi)<20\epsilon^{\ast}}h_{\mu^{\ast}}(f,\xi)>H-\zeta.$ (3.3)
Since $K$ is connected and compact, we choose a sequence $\{\mu_{(k,i)}:k\in\mathbb{N},1\leq i\leq Q_{k}\}\subseteq K$ such that
[…]
$D(\mu_{(k,Q_{k})},\mu_{(k+1,1)})$ $\leq\frac{1}{k},\ \mu_{(k,Q_{k})}=\mu^{\ast}.$
Now the set $\mathcal{A}=\{(0,0)\}\cup\{(k,j):k\geq 1,0\leq j\leq Q_{k}\}$ with lexicographic order. Denote by $\iota(k,j)$ the position of $(k,j)\in\mathcal{A}$. For $(k,j)\in\mathcal{A}$, we shall use $(k,j)^{\ast}$ and $(k,j)_{\ast}$ to represent the one before and the one after $(k,j)$, respectively. Let $\{\delta_{(k,j)}\}$ be a strictly decreasing sequence indexed by $\mathcal{A}$ with $\delta_{(0,0)}<\zeta$, $\delta_{(k,j)}\to 0$ as $(k,j)\to\infty$. We shall write $\delta_{k}:=\delta_{(k,Q_{k})}$. For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$, where $m$ is a positive integer given in the definition of specification property.
For any nonempty open set $U$, there exist $\tilde{\epsilon}>0$ and a point $a_{0}\in X$ such that $\overline{B(a_{0},\tilde{\epsilon})}\subseteq U$. One may assume that $\epsilon^{\ast}<\tilde{\epsilon}$. Set
$\mathcal{C}_{(0,0)}=\{a_{0}\},\ n_{(0,0)}=1,\ N_{(0,0)}=1.$
For any $k\in\mathbb{N}$, there exists $\Delta_{k}=\{a_{1}^{k},\ldots,a_{q_{k}}^{k}\}\subset AP(X)$ which is $\epsilon_{(k,0)}$-dense in $X$ by Proposition 2.2. Then according to specification property, there exists $a_{k}\in X$ such that
$d\left(f^{(i-1)m_{(k,0)}+i-1}(a_{k}),a_{i}^{k}\right)<\epsilon_{(k,0)},\ i=1,2,\ldots,q_{k}.$
Write $\mathcal{C}_{(k,0)}=\{a_{k}\}$, $n_{(k,0)}=q_{k}+(q_{k}-1)m_{(k,0)}$, $N_{(k,0)}=1$.
Let $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$ and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$. The following measure approximation of $\mu^{*}$ derived from a slight modification of [LL24, Lemma 2.1], so we omit the proof.
Lemma 3.3.
For each $k\in\mathbb{N}$, there is a measure $\nu_{k}\in M_{f}(X)$ satisfying
$(1)$ $v_{k}=\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\nu_{i}^{k}$, where $\{\lambda_{i}^{k}:1\leq i\leq j_{k}\}$ are rational positive numbers, $\forall 1\leq i\leq j_{k}$ with $\sum_{i=1}^{j_{k}}\lambda_{i}^{k}=1$ and $\nu_{i}^{k}\in M_{f}^{e}(X)$;
$(2)$ one has
$\inf_{\xi\succ\mathcal{U}}h_{\mu^{\ast}}(f,\xi)\leq\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)+\zeta.$ (3.4)
$(3)$ $D(\mu^{\ast},\nu_{k})<\delta_{k}$.
By the ergodicity of $\nu_{i}^{k}$, there exists $N\in\mathbb{N}$ such that the set
$Y_{i}(N)=\left\{x\in X:D(\mathcal{E}_{n}(x),\nu_{i}^{k})<\delta_{k},\ \forall n>N\right\}$
has $\nu_{i}^{k}$-measure greater than $1-\zeta$ for all $1\leq i\leq j_{k}$. Applying Lemma 2.3 one has
$\underline{h}_{\nu_{i}^{k}}^{K}(f,20\epsilon^{\ast},\zeta)\leq\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)\leq\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)<\infty.$
Now take an integer $n_{k}^{\prime}\in\mathbb{N}$ such that
$(1)$ $\lambda_{i}^{k}n_{k}^{\prime}$ is an integer holds for all $1\leq i\leq j_{k}$, and
$\frac{n_{k}^{\prime}}{n_{k}}>\max\left\{1-\zeta,1-\frac{1}{k}\right\},$ (3.5)
where $n_{k}=n_{(k,Q_{k})}=n_{k}^{\prime}+(j_{k}-1)m_{(k,Q_{k})}$.
$(2)$ there exists a subset $E_{k,i}$ which is a $(\lambda_{i}^{k}n_{k}^{\prime},5\epsilon^{\ast})$-separated set of $Y_{i}(N)$ satisfying
#$\#E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right].$
$(3)$ for every $x\in E_{k,i}$,
$D(\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x),\nu_{i}^{k})<\delta_{k}.$
By the specification property, for $k\geq 1$ and each $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})\in E_{k,1}\times E_{k,2}\times\cdots\times E_{k,j_{k}}$, there exists $y=y(\underline{x}_{k})\in X$ so that
[…]
where $T_{1}=0$ and $T_{i}=\lambda_{1}^{k}n_{k}^{\prime}+\cdots+\lambda_{i-1}^{k}n_{k}^{\prime}+(i-1)m_{k}$ for $2\leq i\leq j_{k}$. One may assume that $d(x,y)<\frac{\epsilon^{\ast}}{2^{k+2}}$ implies $D(\delta_{x},\delta_{y})<\frac{1}{k}$ since the map $x\mapsto\delta_{x}$ is uniform continuous.
For any distinct $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})$, $\underline{x}_{k}^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{j_{k}}^{\prime})$ in $E_{k,1}\times\cdots\times E_{k,j_{k}}$, one has $y(\underline{x}_{k})\neq y^{\prime}(\underline{x}_{k}^{\prime})$. Indeed, if $x_{i}\neq x_{i}^{\prime}$ for some $1\leq i\leq j_{k}$,
$5\epsilon^{\ast}<d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(x_{i},x_{i}^{\prime}\right)\leq d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(x_{i},f^{T_{i}}y\right)+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y,f^{T_{i}}y^{\prime}\right)+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y^{\prime},x_{i}^{\prime}\right)$
$<\frac{\epsilon^{\ast}}{2^{k+1}}+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y,f^{T_{i}}y^{\prime}\right),$
which implies $d_{n_{k}}(y,y^{\prime})>4\epsilon^{\ast}$. Let
$\mathcal{C}_{(k,Q_{k})}=\left\{y(x_{1},x_{2},\ldots,x_{j_{k}}):\ (x_{1},x_{2},\ldots,x_{j_{k}})\in E_{k,1}\times E_{k,2}\times\cdots\times E_{k,j_{k}}\right\}.$
Then $\mathcal{C}_{(k,Q_{k})}$ is a $(n_{k},4\epsilon^{\ast})$-separated set and
[…]
The last inequality holds since ${\rm diam}(\xi)<20\epsilon^{\ast}$ and formulas (3.3), (3.5).
For $n\in\mathbb{N}$, $\varphi\in C(X)$, write $S_{n}\varphi=\sum_{l=0}^{n-1}\varphi\circ f^{i}$. Then for $y\in\mathcal{C}_{(k,Q_{k})}$ and $\varphi\in C(X,[0,1])$ taking values in $[0,1]$,
$\left|\int\varphi\mathop{}\!\mathrm{d}\mathcal{E}_{n_{k}}(y)-\int\varphi\mathop{}\!\mathrm{d}\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y))\right)\right|$
$\leq$ $\left|\frac{1}{n_{k}}\sum_{i=1}^{j_{k}}S_{\lambda_{i}^{k}n_{k}^{\prime}}\varphi(f^{T_{i}}(y))-\frac{1}{n_{k}^{\prime}}\sum_{i=1}^{j_{k}}S_{\lambda_{i}^{k}n_{k}^{\prime}}\varphi(f^{T_{i}}(y))\right|+\frac{n_{k}-n_{k}^{\prime}}{n_{k}}\|\varphi\|$
$\leq$ $\left[\left|\frac{1}{n_{k}^{\prime}}-\frac{1}{n_{k}}\right|\cdot n_{k}^{\prime}+\frac{n_{k}-n_{k}^{\prime}}{n_{k}}\right]\cdot\|\varphi\|<\frac{2}{k}\|\varphi\|,$
which shows that $D(\mathcal{E}_{n_{k}}(y),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)))<\frac{2}{k}$. Recall we have
$D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i}),\nu_{k}\right)<\delta_{k},\ \ D\left(\nu_{k},\mu^{\ast}\right)<\delta_{k}.$
By the construction, one has
$D\left(\mathcal{E}_{n_{k}}(y),\mu^{\ast}\right)\leq\frac{2}{k}+D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i})\right)+2\delta_{k}<\frac{3}{k}+2\delta_{k}.$ (3.7)
Until now, using the specification property we have obtained two sets $\mathcal{C}_{(k,0)}$ and $\mathcal{C}_{(k,Q_{k})}$ for each $k$. We proceed to consider the remaining cases for $1\leq j<Q_{k}$.
Since the system has specification property, $G_{\mu}\neq\emptyset$ holds for any $\mu\in M_{f}(X)$. Applying this fact for any $k\geq 1$ and $1\leq j<Q_{k}$, then there exist a point $x_{k,j}$ and $n_{(k,j)}\in\mathbb{N}$ such that
$D(\mathcal{E}_{n_{(k,j)}}(x_{k,j}),\mu_{(k,j)})<\delta_{(k,j)},$
and
$\frac{n_{(k,j)}}{n_{(k,j)}+m_{(k,j)}}>1-\frac{1}{k}.$
Now set $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$.
Step $1$. Construction of intermediate sets $\mathcal{S}_{(k,j)}$.
Recall that we have defined $N_{(k,0)}=1$ for any $k\geq 0$. Now choose a strictly increasing sequence of integers $\{N_{(k,j)}\}_{(k,j)\in\mathcal{A},~{}j\neq 0}$ such that $N_{(1,1)}=1$ and
$\begin{split}\sum_{0\leq k^{\prime}\leq k+1}(n_{(k^{\prime},0)}+m_{(k^{\prime},0)})+n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}&\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})],\\ \sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})]&\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k},\\ n_{k}N_{(k,Q_{k})}&\geq(1-\zeta)\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}].\end{split}$ (3.8)
For $(k,j)\in\mathcal{A}$ with $j=0$, set $\mathcal{S}_{(k,j)}=\mathcal{C}_{(k,j)}$ and $c_{k,j}=n_{(k,j)}$. We divide the following two cases to define the rest $\mathcal{S}_{(k,j)}$ indexed by $j\neq 0$.
Case $1$. $j\neq Q_{k}$.
Recall that $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$. By specification property, there exists $y_{k,j}\in X$ so that the pieces of orbits $\{x_{k,j},\cdots,f^{n_{(k,j)-1}}(x_{(k,j)})\}$ repeating $N_{(k,j)}$ times, can be $\epsilon_{(k,j)}$-traced by $y_{(k,j)}$ with gap $m_{(k,j)}$. Now set $\mathcal{S}_{(k,j)}=\{y_{k,j}\}$, and $c_{k,j}=n_{(k,j)}N_{(k,j)}+(N_{(k,j)}-1)m_{(k,j)}$.
Case $2$. $j=Q_{k}$.
Enumerate the points in $\mathcal{C}_{(k,Q_{k})}$ as
$\mathcal{C}_{(k,Q_{k})}=\{y_{i}^{k}:i=1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}.$
For any $(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}$, let $y(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in X$ be given by the specification property, such that the orbit $\epsilon_{(k,Q_{k})}$-shadows, with gap $m_{(k,Q_{k})}$, the pieces of orbits $\{y_{i_{l}}^{k},f(y_{i_{l}}^{k}),\ldots,f^{n_{k}-1}(y_{i_{l}}^{k})\}$, $l=1,2,\ldots,N_{(k,Q_{k})}$. Then define
$\mathcal{S}_{(k,Q_{k})}=\left\{y=y(i_{1},\ldots,i_{N_{(k,Q_{k})}}):(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}\right\}.$
Write $c_{k,Q_{k}}=n_{k}\cdot N_{(k,Q_{k})}+(N_{(k,Q_{k})}-1)m_{(k,Q_{k})}$, then it follows from the facts that $\mathcal{C}_{(k,Q_{k})}$ is a $(n_{k},4\epsilon^{\ast})$-separated set, and $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set. In particular, by (3.6) and (3.8), one has
#$\#\mathcal{S}_{(k,Q_{k})}=\#\mathcal{C}_{(k,Q_{k})}^{N_{(k,Q_{k})}}\geq\exp\left[\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}]\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$ (3.9)
Similar to the argument in (3.7) one has
$D(\mathcal{E}_{c_{k,j}}(y_{k,j}),\mu_{(k,j)})<\frac{2}{k}+\delta_{(k,j)},\ y_{k,j}\in\mathcal{S}_{(k,j)}\ \text{for }k\geq 1\ \text{and }1\leq j<Q_{k},$
and
$D(\mathcal{E}_{c_{k,Q_{k}}}(y),\mu^{\ast})<\frac{5}{k}+2\delta_{k},\ \forall y\in\mathcal{S}_{(k,Q_{k})}.$
Hence, in an unified way we always have
$D(\mathcal{E}_{c_{k,j}}(y),\mu_{(k,j)})<\frac{5}{k}+2\delta_{(k,j)},\ \forall y\in\mathcal{S}_{(k,j)}\ \text{with }k\geq 1,\ 1\leq j\leq Q_{k}.$ (3.10)
Step $2$. Constructions of $\mathbb{F}_{(k,j)}$ and Moran set $\mathbb{F}$.
Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$. Now suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$, and then we will describe how to construct $\mathcal{I}_{(k,j)}$. Let
$M_{(k,j)}$ $=M_{(k,j)^{\ast}}+m_{(k,j)}+c_{k,j};$
[…]
Note that $M_{(k,j)}^{N_{(k,j)}}=M_{(k,j)}$ and the number of $\mathcal{S}_{(k,Q_{k})}$ given in formula (3.9) has the estimation:
#$\#\mathcal{S}_{(k,Q_{k})}\geq\exp\left[(M_{(k,Q_{k})}-M_{(k-1,Q_{k-1})})(1-\zeta)^{2}(H-3\zeta)\right].$
For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, by specification property let $z=z(x,y)\in X$ be some point such that
$d_{M_{(k,j)^{\ast}}}(x,z)<\epsilon_{(k,j)},\ d_{c_{k,j}}(y,f^{M_{(k,j)^{\ast}}+m_{(k,j)}}(z))<\epsilon_{(k,j)}.$
We set $\mathcal{I}_{(k,j)}=\{z(x,y):x\in\mathcal{I}_{(k,j)^{\ast}},y\in\mathcal{S}_{(k,j)}\}$. For $x\in\mathcal{I}_{(k,Q_{k})^{\ast}}$ and distinct $y,y^{\prime}\in\mathcal{S}_{(k,Q_{k})}$, since $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set, one has
$d_{M_{k,Q_{k}}}(z(x,y),z(x,y^{\prime}))>2\epsilon^{\ast},$
which leads to
#$\#\mathcal{I}_{(k,j)}=\#\mathcal{I}_{(k,j)^{\ast}}\cdot\#\mathcal{S}_{(k,j)}=\Pi_{(k^{\prime},j^{\prime})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},j^{\prime})}=\Pi_{(k^{\prime},Q_{k^{\prime}})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}.$
Moreover, according to (3.10), for $z\in\mathcal{I}_{(k,j)}$ with $j\neq 0$, we have
$D(\mathcal{E}_{c_{k,j}}(f^{M_{(k,j)^{\ast}}}(z)),\mu_{(k,j)})<\frac{6}{k}+2\delta_{(k,j)}.$ (3.11)
Now set
$\mathbb{F}_{(k,j)}=\bigcup_{x\in\mathcal{I}_{(k,j)}}\overline{B}_{M_{(k,j)}}(x,2\epsilon_{(k,j)}).$
It is easy to see that for each $(k,j)\in\mathcal{A}$ the following is satisfied:
$(1)$ $\overline{B}_{M_{(k,j)}}(x,2\epsilon_{(k,j)})\cap\overline{B}_{M_{(k,j)}}(x^{\prime},2\epsilon_{(k,j)})=\emptyset$ holds for any distinct $x,x^{\prime}\in\mathcal{I}_{(k,j)}$;
$(2)$ if $z\in\mathcal{I}_{(k,j)}$ descends from $x\in\mathcal{I}_{(k,j)^{\ast}}$, one has $\overline{B}_{M_{(k,j)}}(z,2\epsilon_{(k,j)})\subset\overline{B}_{M_{(k,j)^{\ast}}}(x,2\epsilon_{(k,j)^{\ast}})$. That is $\{\mathbb{F}_{(k,j)}:(k,j)\in\mathcal{A}\}$ is a sequence of decreasing nonempty closed subsets of $X$.
Let
$\mathbb{F}=\bigcap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)}=\bigcap_{k\in\mathbb{N}}\mathbb{F}_{(k,Q_{k})}.$
Lemma 3.4.
One has the fractal set $\mathbb{F}$ is contained in $\mathbb{F}\subset G_{K}^{\omega,U}$, that is,
$\mathbb{F}\subset G_{K}^{\omega,U}=G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\}.$
Proof.
Firstly, we know $\mathbb{F}\subset\mathbb{F}_{(0,0)}=\overline{B}_{n(0,0)}(a_{0},\frac{\epsilon^{\ast}}{8})\subset\overline{B(a_{0},\tilde{\epsilon})}\subset U$. For $p\in\mathbb{F}$ and any $k\geq 1$, we have $p\in\mathbb{F}_{(k,0)}$. Hence for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there is a $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<3\epsilon_{(k,0)}$, which implies that the orbit of $p$ is $4\epsilon_{(k,0)}$-dense in $X$. By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have $\mathbb{F}\subset\{x\in X:\omega_{f}(x)=X\}$.
For $p\in\mathbb{F}$ and sufficiently large $n\in\mathbb{N}$, in order to estimate $D(\mathcal{E}_{n}(p),K)$, we split into four cases to discuss.
Case $(1)$. There exists $k>1$ such that $M_{(k+1,0)}\leq n<M_{(k+1,1)}^{1}$.
Then there is some $z\in\mathcal{I}_{(k,Q_{k})}$ such that $p\in\overline{B}_{M_{(k,Q_{k})}}(z,\frac{\epsilon_{(k,Q_{k})}}{2})$, and
$D(\mathcal{E}_{n}(p),\mu^{\ast})\leq$ $\frac{M_{(k,Q_{k})^{\ast}}}{n}D(\mathcal{E}_{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})+\frac{c_{k,Q_{k}}}{n}D(\mathcal{E}_{c_{k,Q_{k}}}(f^{M_{(k,Q_{k})^{\ast}}}(p)),\mu^{\ast})$
$+\frac{n-M_{(k,Q_{k})}}{n}D(\mathcal{E}_{n-M_{(k,Q_{k})}}(f^{M_{(k,Q_{k})}}(p)),\mu^{\ast}).$
Noticing that both $\frac{M_{(k,Q_{k})^{\ast}}}{n}$ and $\frac{n-M_{(k,Q_{k})}}{n}$ are smaller than $\frac{1}{k}$, and using formula (3.11), we have $D(\mathcal{E}_{c_{k,Q_{k}}}f^{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})<\frac{7}{k}+2\delta_{(k,Q_{k})}$. Thus,
$D(\mathcal{E}_{n}(p),K)\leq\frac{9}{k}+2\delta_{(k,Q_{k})}.$
Case $(2)$. There exists $(k,j)\in\mathcal{A}$ with $j\neq 0$, such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
Then there is some $z\in\mathcal{I}_{(k,j)}$ such that $p\in\overline{B}_{M_{(k,j)}}(z,\frac{\epsilon_{(k,j)}}{2})$, and
$D(\mathcal{E}_{n}(p),\mu_{(k,j)})\leq$ $\frac{M_{(k,j)^{\ast}}}{n}D(\mathcal{E}_{M_{(k,j)^{\ast}}}(p),\mu_{(k,j)})+\frac{c_{k,j}}{n}D(\mathcal{E}_{c_{k,j}}(f^{M_{(k,j)^{\ast}}}(p)),\mu_{(k,j)})$
$+\frac{n-M_{(k,j)}}{n}D(\mathcal{E}_{n-M_{(k,j)}}(f^{M_{(k,j)}}(p)),\mu_{(k,j)}).$
The similar argument as in Case $(1)$ yields that
$D(\mathcal{E}_{n}(p),K)\leq\frac{9}{k}+2\delta_{(k,j)}.$
Case $(3)$. There exist $(k,1)\in\mathcal{A}$ and $i>1$, such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
One has
$D($ $\mathcal{E}_{n}(p),\mu_{(k,1)})\leq\frac{M_{(k-1,Q_{k-1})^{\ast}}}{n}D(\mathcal{E}_{M_{(k-1,Q_{k-1})^{\ast}}}(p),\mu_{(k,1)})$
$+\frac{c_{k-1,Q_{k-1}}}{n}D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})+\frac{c_{k,0}}{n}D(\mathcal{E}_{c_{k,0}}(f^{M_{(k-1,Q_{k-1})}}(p)),\mu_{(k,1)})$
$+\frac{n_{(k,1)}+m_{(k,1)}}{n}\sum_{l=0}^{i-1}D(\mathcal{E}_{n_{(k,1)}+m_{(k,1)}}(f^{M_{(k,1)}^{l}}(p)),\mu_{(k,1)})$
$+\frac{n-M_{(k,1)}^{i}}{n}D(\mathcal{E}_{n-M_{(k,1)}^{i}}(f^{M_{(k,1)}^{i}}(p)),\mu_{(k,1)}).$
Note that
$\frac{M_{(k-1,Q_{k-1})^{\ast}}}{n}<\frac{1}{k-1},\ \frac{c_{k,0}}{n}<\frac{1}{k},\ \frac{n-M_{(k,1)}^{i}}{n}<\frac{1}{k},$
$D(\mathcal{E}_{n_{(k,1)}+m_{(k,1)}}(f^{M_{(k,1)}^{l}}(p)),\mu_{(k,1)})\leq\frac{5}{k}+2\delta_{(k,1)},$
and
$D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})$ $\leq D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu^{\ast})+D(\mu^{\ast},\mu_{(k,1)})$
$\leq\frac{8}{k-1}+2\delta_{(k-1,Q_{k-1})}.$
This gives us
$D(\mathcal{E}_{n}(p),K)\leq\frac{16}{k-1}+4\delta_{(k-1,0)}.$
Case $(4)$. There exists $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$, such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
In this case, using the expression
$\mathcal{E}_{n}(z)=$ $\frac{M_{(k,j-1)^{\ast}}}{n}\mathcal{E}_{M_{(k,j-1)^{\ast}}}(p)+\frac{c_{k,j-1}}{n}\mathcal{E}_{c_{k,j-1}}(f^{M_{(k,j-1)^{\ast}}}(p))$
$+$ $\frac{n_{(k,j)}+m_{(k,j)}}{n}\sum_{l=0}^{i-1}\mathcal{E}_{n_{(k,j)}+m_{(k,j)}}(f^{M_{(k,j)}^{l}}(p))+\frac{n-M_{(k,1)}^{i}}{n}\mathcal{E}_{n-M_{(k,1)}^{i}}(f^{M_{(k,1)}^{i}}(p)),$
and repeating the similar argument as above, one can obtain
$D(\mathcal{E}_{n}(p),K)\leq\frac{15}{k}+4\delta_{(k,j)}.$
To sum up, whatever case we consider, by letting $n\to\infty$ we get $V_{f}(p)\subset K$. Based on above arguments, one can also show sequences $\{\mathcal{E}_{n}(p)\}_{n\in\mathbb{N}}$ and $\{\mu_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ have the same limit points. Notice that $K=\overline{\{\mu_{(k^{\prime},j^{\prime})}:(k^{\prime},j^{\prime})>(k,j)\}}$ holds for any fixed $(k,j)\in\mathcal{A}$. Consequently, we have $V_{f}(p)=K$, and hence $\mathbb{F}\subset G_{K}$. ∎
Step $3$. Construction of measure.
Now we define a probability measure $\rho$ on $\mathbb{F}$ satisfying that the $\rho$-measure of some appropriate dynamical balls decay exponentially fast along sequence $\{M_{k}\}$, where $M_{k}=M_{k,Q_{k}}$. Then we can use the Entropy Distribution Principle developed in Lemma 2.2 to estimate the lower bound of $\epsilon$-packing topological entropy of $\mathbb{F}$.
Define
$\rho_{k}=\frac{1}{\#\mathcal{I}_{(k,Q_{k})}}\sum_{x\in\mathcal{I}_{(k,Q_{k})}}\delta_{x}.$
For simplicity, we write $\mathcal{I}_{k}$, $\mathbb{F}_{k}$ instead of $\mathcal{I}_{(k,Q_{k})}$, $\mathbb{F}_{(k,Q_{k})}$, respectively. Since $M(X)$ is compact, we may assume that $\rho_{k}\to\rho\in M(X)$ as $k\to\infty$. It is clear that $\rho_{k}(\mathbb{F}_{k})=1$ for every $k$, and for any $k>k_{0}$, one has $\rho_{k}(\mathbb{F}_{k_{0}})=1$ since $\mathbb{F}_{k}\subset\mathbb{F}_{k_{0}}$. Thus $\rho(\mathbb{F}_{k_{0}})\geq\limsup_{k\to\infty}\rho_{k}(\mathbb{F}_{k_{0}})=1$, which implies that $\rho(\mathbb{F})=1$. By (3.9), we have
#$\#\mathcal{I}_{k}\geq\exp\left[M_{k}\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
Lemma 3.5.
Let $k_{0}\geq 1$, and $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ be any Bowen ball which intersects $\mathbb{F}$. Then for any $k>k_{0}$, one has
$\rho_{k}\left(B_{M_{k_{0}}}\left(z,\frac{\epsilon^{\ast}}{2}\right)\right)\leq\frac{1}{\#\mathcal{I}_{k_{0}}}.$
Proof.
If $k=k_{0}+1$, it is easy to check that
#$\#(\mathbb{F}_{k_{0}+1}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))\leq\#\mathcal{S}_{(k_{0}+1,Q_{k_{0}+1})},$
and hence
$\rho_{k_{0}+1}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{\#\mathcal{S}_{(k_{0}+1,Q_{k_{0}+1})}}{\#\mathcal{I}_{k_{0}+1}}=\frac{1}{\#\mathcal{I}_{k_{0}}}.$
In general, for any $k>k_{0}$,
$\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{\Pi_{k_{0}+1\leq k^{\prime}\leq k}\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}}{\#\mathcal{I}_{k}}=\frac{1}{\#\mathcal{I}_{k_{0}}}.$
For large enough $k_{0}$ and any Bowen ball $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ with $z\in\mathbb{F}$, since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ is open, one has
$\rho\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)$ $\leq\liminf_{k\to\infty}\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{1}{\#\mathcal{I}_{k_{0}}}$
$\leq\exp\left[-M_{k_{0}}(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
Using Lemma 2.2, one has
$(1-\zeta)^{2}\cdot(H-3\zeta)\leq h_{top}^{P}(f,\mathbb{F},\frac{\epsilon^{\ast}}{12})\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$
The arbitrariness of $\zeta$ implies that
$\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$
Now dividing $|\log\epsilon^{\ast}|$ in both sides and then taking $\limsup_{\epsilon^{\ast}\to 0}$, we have
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d).$ (3.12)
Finally, by (3.1) and (3.12) we get
$\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
3.2 Bowen metric mean dimension of saturated sets
We prove Theorems 1.2 and 1.3 in this subsection.
3.2.1 Upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$
The following lemma gives an upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$, which was already given in [Y23, Lemma 3.3,(3)]
Lemma 3.6.
Let $(X,f)$ be a t.d.s.. Let $K\subset M_{f}(X)$ be a nonempty closed connected subset. Then for any $\epsilon>0$, we have
$h_{top}^{B}(f,G_{K},\epsilon)\leq\inf_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Thus, by the Lemma 3.6 we obtain an upper bound of the Bowen metric mean dimension of saturated set:
$\overline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.13)
3.2.2 Lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
The proof of the lower bound is similar to that of in subsection 3.1.2. So we only give a sketch for the estimation of lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$.
We shall show that for any small enough $\epsilon^{\ast}>0$, it holds that
$\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2}).$ (3.14)
Fix $\zeta>0$ and pick a sequence $\{\mu_{(k,i)}:k\geq 1,1\leq i\leq Q_{k}\}\subset K$ of measures in $K$ such that
$K\subset\cup_{i=1}^{Q_{k}}B(\mu_{(k,i)},\frac{1}{k}),\ D(\mu_{(k,i)},\mu_{(k,i+1)})<\frac{1}{k},\ D(\mu_{(k,Q_{k})},\mu_{(k+1,1)})<\frac{1}{k}.$
Put $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$ and endow
$\mathcal{A}=\{(k,0):k\geq 0\}\cup\mathcal{A}^{\prime}$
with the lexicographic order again. Using the fixed notation in subsection 3.1.2, we have the data $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ and $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$.
Assume that $H:=\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)>0$. Then one has
$\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu_{(k,i)}}(f,\xi)>H-\zeta,\ \forall(k,i)\in\mathcal{A}^{\prime}.$
Let $\mathcal{U}$ be an open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$, and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$. For $(k,i)\in\mathcal{A}^{\prime}$, by Lemma 3.3 let $\nu_{(k,i)}=\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\nu_{l}^{k,i}$ be the finite convex combination of ergodic $f$-invariant probability measures with rational coefficients such that
$D(\mu_{(k,i)},\nu_{(k,i)})<\delta_{(k,i)},\ \ \inf_{\xi\succ\mathcal{U}}h_{\mu_{(k,i)}}(f,\xi)\leq\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)+\zeta.$
Applying Katok’s $\epsilon$-entropy formula and repeating the previous argument as in subsection 3.1.2, we choose an integer $n_{(k,i)}^{\prime}$ such that the following conditions are satisfied:
$(1)$ for any $1\leq l\leq p_{k,i}$, $\lambda_{k,i,l}n_{(k,i)}^{\prime}$ is an integer and
$\frac{n_{(k,i)}^{\prime}}{n_{(k,i)}+m_{(k,i)}}>\max\{1-\zeta,1-\frac{1}{k}\},$
where $n_{(k,i)}:=n_{(k,i)}^{\prime}+(p_{k,i}-1)m_{(k,i)}$.
$(2)$ there exist subsets $E_{k,i,l}$, $1\leq l\leq p_{k,i}$ which are $(\lambda_{k,i,l}n_{(k,i)}^{\prime},5\epsilon^{\ast})$-separated and
#$\#E_{k,i,l}\geq\exp\left[\lambda_{k,i,l}n_{(k,i)}^{\prime}\cdot\left(\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)-\zeta\right)\right];$
$(3)$ for any $x\in E_{k,i,l}$,
$D(\mathcal{E}_{\lambda_{k,i,l}n_{(k,i)}^{\prime}}(x),\nu_{l}^{k,i})<\delta_{(k,i)}.$
Let
$\mathcal{C}_{(k,i)}=\left\{y=y(x_{1},\ldots,x_{p_{k,i}}):(x_{1},\ldots,x_{p_{k,i}})\in E_{k,i,1}\times\cdots\times E_{k,i,p_{k,i}}\right\},$
where $y(x_{1},\ldots,x_{p_{k,i}})$ is the point that $\epsilon_{(k,i)}$-shadows the pieces of orbits
[…]
Then $\mathcal{C}_{(k,i)}$ is a $(n_{(k,i)},4\epsilon^{\ast})$-separated set with
#$\#\mathcal{C}_{(k,i)}\geq\exp\left[n_{(k,i)}(1-\zeta)\cdot(H-2\zeta)\right].$
For $k\geq 0$, let $\mathcal{C}_{(k,0)}$, $n_{(k,0)}$, $N_{(k,0)}$ be defined as in subsection 3.1.2. Choose a strictly increasing sequence $\{N_{(k,i)}\}_{(k,i)\in\mathcal{A}^{\prime}}$ such that $N_{(1,1)}=1$ and
$\begin{split}\sum_{0\leq k^{\prime}\leq k+1}(n_{(k^{\prime},0)}+m_{(k^{\prime},0)})+n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}&\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})],\\ \sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})]&\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k}.\end{split}$
Set $\mathcal{S}_{(k,0)}=\mathcal{C}_{(k,0)}$, $c_{k,0}=n_{(k,0)}$ for $k\geq 0$. For $(k,i)\in\mathcal{A}^{\prime}$, and any tuple of points $(y_{k,i,1},y_{(k,i,2)},\ldots,y_{k,i,N_{(k,i)}})\in\mathcal{C}_{(k,i)}^{N_{(k,i)}}$, there exists a point $z=z(y_{k,i,1},y_{(k,i,2)},\ldots,y_{k,i,N_{(k,i)}})$ that $\epsilon_{(k,i)}$-shadows the pieces of orbits
$\{y_{k,i,l},f(y_{k,i,l}),\ldots,f^{n_{(k,i)}-1}(y_{k,i,l})\},\ l=1,\cdots,N_{(k,i)},$
with the same gap $m_{(k,i)}$. Denote by $\mathcal{S}_{(k,i)}$ the set of such shadowing points, which is $(c_{k,i},3\epsilon^{\ast})$-separated set, where $c_{k,i}=n_{(k,i)}N_{(k,i)}+(N_{(k,i)}-1)m_{(k,i)}$.
Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$. Suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$. Let $M_{(k,j)}=M_{(k,j)^{\ast}}+m_{(k,j)}+c_{k,j}$. For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, let $z=z(x,y)$ be the shadowing point such that $d_{M_{(k,j)^{\ast}}}(z,x)<\epsilon_{(k,j)}$ and
$d_{c_{k,j}}(y,f^{M_{(k,j)^{\ast}+m_{(k,j)}}}(z))<\epsilon_{(k,j)}.$
We set $\mathcal{I}_{(k,j)}=\{z=z(x,y):x\in\mathcal{I}_{(k,j)^{\ast}},y\in\mathcal{S}_{(k,j)}\}$, and then define
$\mathbb{F}=\cap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)},$
where $\mathbb{F}_{(k,j)}=\cup_{z\in\mathcal{I}_{(k,j)}}\overline{B}_{M_{(k,j)}}(z,2\epsilon_{(k,j)})$. One can verify that $\mathbb{F}\subset G_{K}^{\omega,U}$.
We proceed to define a sequence of measures $\{\rho_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ by
$\rho_{(k,j)}:=\frac{1}{\#\mathcal{I}_{(k,j)}}\sum_{x\in\mathcal{I}_{(k,j)}}\delta_{x}.$
Without loss of generality, we can assume that $\rho_{(k,j)}\to\rho$. Furthermore, one can check that $\rho(\mathbb{F})=1$, and for every $n$ large enough and any ball $B_{n}(z,\frac{\epsilon^{\ast}}{2})$ intersecting with $\mathbb{F}$, it holds that
$\rho(B_{n}(z,\frac{\epsilon^{\ast}}{2}))\leq\exp\left[-n(1-\zeta)^{2}(H-2\zeta)\right].$
Then by the Lemma 2.1, we have
$(1-\zeta)^{2}(H-2\zeta)\leq h_{top}^{B}(f,\mathbb{F},\frac{\epsilon^{\ast}}{2})\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2}).$
Letting $\zeta\to 0$ gives us
$\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2})$
Therefore, by (3.13) and (3.14) we get
$\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{B}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{\mathrm{diam}\xi<\epsilon}h_{\mu}(f,\xi).$ (3.15)
3.2.3 Proof of Theorem 1.3
Firstly, we show
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\}.$ (3.16)
Fix $\mu\in K$ and $(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)$. Then, by Lemma 3.6, for every $\epsilon>0$,
$h_{top}^{B}(f,G_{K},\epsilon)\leq\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
This shows
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)$
and hence $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\underline{\rm mrid}_{\mu}(f,K,d)$ for every $\mu\in K$.
On the other hand, we may assume that $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)$ is finite; otherwise there is nothing left to prove. Let $\gamma>0$. Then according to Theorem 1.2 we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence of $\{\mu_{k}\}$ in $K$ such that for any $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi)<\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
Assume that $\mu_{k}\rightarrow\mu$. Then $\mu\in M_{f}(X)$ since $K$ is closed. This shows
$\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
Letting $\gamma\to 0$, we have
$\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\}\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d).$
This completes the proof of (3.16).
Using formula (3.2), the inequality
$\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}\leq\overline{\rm mdim}_{M}^{P}(f,G_{K},d).$
is clear. On the other hand, let $\gamma>0$ be sufficiently small and put $M_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{P}(f,G_{K},d)-\gamma\}$.Then we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence of $\{\mu_{k}\}$ in $K$ such that for sufficiently large $k$,
$M_{\gamma}<\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi).$
Again, we assume that $\mu_{k}\rightarrow\mu\in K$. This shows
$M_{\gamma}\leq\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
Letting $\gamma\to 0$, we have
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)\leq\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
3.3 Upper capacity metric mean dimension of saturated sets
We prove the Theorem 1.4 in this subsection.
The inequality
$\overline{\rm mdim}_{M}^{UC}\left(f,G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},d\right)\leq\overline{\rm mdim}_{M}^{UC}\left(f,G_{K},d\right)\leq\overline{\rm mdim}_{M}(f,X,d)$
is clear. Thus, it suffices to show
$\overline{\rm mdim}_{M}(f,X,d)\leq\overline{\rm mdim}_{M}^{UC}\left(f,G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},d\right).$
Once we prove that for $\epsilon^{\ast}>0$, $\zeta>0$ small enough and $\delta\in(0,1)$, it holds that
$\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta\leq h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast}),$
then by Theorem 2.4, the proof of Theorem 1.4 will be done. Since the proof is similar to the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$ in subsection 3.1.2, we give a sketch for it.
We write $H=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)>0$. Take $\mu^{\ast}\in M_{f}^{e}(X)$ such that
$\underline{h}_{\mu^{\ast}}^{K}(f,3\epsilon^{\ast},\delta)>H-\zeta.$
Then for large enough $\mathcal{N}\in\mathbb{N}$, there exists a $(\mathcal{N},3\epsilon^{\ast})$-separated subset $S_{\mathcal{N}}$ with
#$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)].$
Now fix such $\mathcal{N}$, and write $\mathcal{C}_{(0,0)}=S_{\mathcal{N}}$, $c_{(0,0)}=\mathcal{N}$, $N_{(0,0)}=1$.
Since $K$ is a connected closed subset, a standard argument shows that there exists a family $\{\mu_{k,i}:k\geq 1,1\leq i\leq Q_{k}\}\subset K$ of measures in $K$ such that for all $k\geq 1$, $1\leq i<Q_{k}$,
$K\subset\bigcup_{i=1}^{Q_{k}}B(\mu_{k,i},\frac{1}{k}),\ D(\mu_{k,i},\mu_{k,i+1})<\frac{1}{k},\ \text{and }D(\mu_{k,Q_{k}},\mu_{k+1,1})<\frac{1}{k}.$
Consider the sets $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$, and
$\mathcal{A}=\{(k,0):k\in\mathbb{Z}\ \text{and }k\geq-1\}\cup\mathcal{A}^{\prime}$
with the lexicographic order. The sequences $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ are constructed similarly as in subsection 3.1.2.
For the nonempty open set $U$, there exist $a_{0}\in X$ and $\epsilon_{0}\in(0,\epsilon^{\ast})$ with $\overline{B(a_{0},\epsilon_{0})}\subset U$. Set $\mathcal{C}_{(-1,0)}=\{a_{0}\}$, $n_{(-1,0)}=c_{(-1,0)}=N_{(-1,0)}=1$. For $k\geq 1$, let $\mathcal{C}_{(k,0)}$, $n_{(k,0)}$, $c_{(k,0)}$ and $N_{(k,0)}$ be the same as in subsection 3.1.2. For $(k,i)\in\mathcal{A}^{\prime}$, since $G_{\mu_{k,i}}\neq\emptyset$, one can pick $x_{k,i}\in X$ and $n_{(k,i)}\in\mathbb{N}$ such that
$D(\mathcal{E}_{n_{(k,i)}}(x),\mu_{k,i})<\delta_{(k,i)},\ \frac{n_{(k,i)}}{n_{(k,i)}+m_{(k,i)}}>\max\left\{1-\zeta,1-\frac{1}{k}\right\}.$
Set $\mathcal{C}_{(k,i)}=\{x_{k,i}\}$. Now choose a strictly increasing sequence of integers $\{N_{(i,j)}\}_{(k,j)\in\mathcal{A}\ \text{with }j\neq 0}$ with $N_{(1,1)}=1$ and
$\begin{split}&n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}\left[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})\right],\\ &\sum_{(k^{\prime},j^{\prime})\leq(k,j)}\left[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})\right]\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k}.\end{split}$
Based on the data $\mathcal{C}_{(k,j)}$, $N_{(k,j)}$ for $(k,j)\in\mathcal{A}$, similar to the previous subsections, one can construct the intermediate sets $\mathcal{S}_{(k,j)}$, $\mathcal{I}_{(k,j)}$, $\mathbb{F}_{(k,j)}$ to obtain the fractal set
$\mathbb{F}=\bigcap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)}\subset G_{K}^{\omega,U}.$
It is worth mentioning that for any $(k,j)>(0,0)$, #$\#\mathcal{I}_{(k,j)}=\#S_{\mathcal{N}}$. Moreover, for any distinct $x$, $x^{\prime}\in\mathcal{I}_{(0,0)}$, one has
$d_{M_{(0,0)}}(x,x^{\prime})>2\epsilon^{\ast}.$
Recall that $M_{(0,0)}=1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}$. Hence for $y$, $y^{\prime}\in\mathbb{F}$ with $y\in\overline{B}_{M_{(0,0)}}(x,2\epsilon_{(0,0)})$, and $y^{\prime}\in\overline{B}_{M_{(0,0)}}(x^{\prime},2\epsilon_{(0,0)})$ one has
$d_{M_{(0,0)}}(y,y^{\prime})>d_{M_{(0,0)}}(x,x^{\prime})-d_{M_{(0,0)}}(x,y)-d_{M_{(0,0)}}(x^{\prime},y^{\prime})>\epsilon^{\ast}.$
Hence for any large enough $\mathcal{N}$, there is a $(1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N},\epsilon^{\ast})$-separated subset in $\mathbb{F}$ with cardinality #$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)]$. Thus
$h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast})\geq\limsup_{\mathcal{N}\to\infty}\frac{\mathcal{N}(H-2\zeta)}{1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}}=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta.$
4 Applications
In this section, we apply the main results to multifractal analysis and mean Li-Yorke chaos. The main results are Theorems 4.1 and 4.3.
4.1 Variational principle of level sets
In this subsection, we establish variational principles for level sets of continuous functions on $M_{f}(X)$ in the sense of an abstract framework.
Let $(X,f)$ be a dynamical system, and let $\varphi:M_{f}(X)\to\mathbb{R}$ be a continuous function. Let
$L_{\varphi}=\left[\inf_{\mu\in M_{f}(X)}\varphi(\mu),\ \sup_{\mu\in M_{f}(X)}\varphi(\mu)\right],\$
For $\alpha\in L$, let $M_{f}(\varphi,\alpha):=\left\{\mu\in M_{f}(X):\varphi(\mu)=\alpha\right\}$, and let
$K_{\alpha}:=\left\{x\in X:\inf_{\mu\in V_{f}(x)}\varphi(\mu)=\sup_{\mu\in V_{f}(x)}\varphi(\mu)=\alpha\right\}$
denote the level set of $\alpha$ w.r.t. $\varphi$ consisting of the set of points of $X$ such that $\varphi$ is taking constant value $\alpha$ on $V_{f}(x)$. Then the continuity of $\varphi$ implies that $M_{f}(\varphi,\alpha)$ is a nonempty closed subset of $M_{f}(X)$.
Condition A: Given a continuous function $\varphi$ on $M_{f}(X)$, for every $\alpha\in L_{\varphi}$ we always assume that $M_{f}(\varphi,\alpha)$ is a convex subset.
Some examples of $\varphi$ satisfying Condition A are given as follows:
  • (a)
    If $\varphi=\alpha\in\mathbb{R}$ is a constant, then $L_{\varphi}=\{\alpha\}$, $K_{\alpha}=X$, $M_{f}(\varphi,\alpha)=M_{f}(X)$;
  • (b)
    Let $\psi$, $\phi\in C(X)$ such that $\int\phi\mathop{}\!\mathrm{d}\mu\neq 0$ holds for all $\mu\in M_{f}(X)$ (e.g. $\phi>0$). We define
    $\varphi(\mu):=\frac{\int\psi\mathop{}\!\mathrm{d}\mu}{\int\phi\mathop{}\!\mathrm{d}\mu}.$
    Then for any $\alpha\in L_{\varphi}$,
    $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\int\phi\mathop{}\!\mathrm{d}\mu\},$
    $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\int\phi\mathop{}\!\mathrm{d}\mu\};$
  • (c)
    Let $\Phi:=\{\phi_{n}\}$ be an asymptotically additive potential, i.e. for any $\delta>0$, there exists $\psi_{\delta}\in C(X)$ such that $\limsup_{n\to\infty}\frac{1}{n}||\phi_{n}-S_{n}\psi_{\delta}||<\delta$. We define
    $\varphi:M_{f}(X)\to\mathbb{R},~{}\mu\mapsto\lim_{n\to\infty}\int\frac{\phi_{n}(x)}{n}\mathop{}\!\mathrm{d}\mu(x).$
    Then $\varphi$ is continuous (cf. [FH10, Lemma A.4]), and $M_{f}(\varphi,\alpha)$ is a convex subset for any $\Phi$ and $\alpha\in L_{\Phi}$.
    Specially, when the sequence $\{\phi_{n}\}$ is an additive function $\{S_{n}\psi\}$ with $\psi\in C(X)$, in this case we have
    $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\ \int\psi\mathop{}\!\mathrm{d}\mu=\alpha\},$
    $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
Given a nonempty open subset $U\subset X$, we put
$K_{\alpha}^{\omega,U}=K_{\alpha}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}.$
Theorem 4.1.
Let $(X,d,f)$ be a dynamical system with specification property and $U\subset X$ be a nonempty open subset. Suppose that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function satisfying Condition A. Then for any $\alpha\in L$ and $S\in\{B,P\}$, we have
$\overline{\rm mdim}_{M}^{S}(f,K_{\alpha}^{\omega,U},d)=$ $\overline{\rm mdim}_{M}^{S}(f,K_{\alpha},d)$
$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
[…]
Proof.
Put $G_{\mu}^{\omega,U}:=G_{\mu}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}$. By (3.14), there exists a constant $c>0$, which is not dependent on $\mu$, such that
$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu}^{\omega,U},\epsilon)$ (4.1)
holds for all $\mu\in M_{f}(X)$ and any small enough $\epsilon>0$. Together with the fact $G_{\mu}^{\omega,U}\subset K_{\alpha}^{\omega,U}$ holds for any $\mu\in M_{f}(\varphi,\alpha)$, we have
$\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,K_{\alpha}^{\omega,U},\epsilon)$
for all sufficiently small $\epsilon>0$, which implies that
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\overline{\rm mdim}_{M}^{B}(f,K_{\alpha}^{\omega,U},d).$ (4.2)
Recall $G^{M_{f}(\varphi,\alpha)}=\{x\in X:V_{f}(x)\subset M_{f}(\varphi,\alpha)\}$. Since the statement $\varphi(\mu)=\alpha$ for all $\mu\in V_{f}(x)$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\varphi,\alpha)$, then $K_{\alpha}=G^{M_{f}(\varphi,\alpha)}$. By Lemma 3.2 $(1)$, we get
$\overline{\rm mdim}_{M}^{P}(f,K_{\alpha},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (4.3)
By (4.2), (4.3), and the fact stated in Proposition 2.1 $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)$ for any $Z\subset X$, we complete the proof. ∎
A direct consequence of Theorem 4.1 is the variational principles for Bowen metric mean dimension of level sets obtained by Backes and Rodrigues [BR23, Theorem A], but we can strength the result which states that the variational principle is still valid for packing metric mean dimension of level sets.
Corollary 4.2.
Let $(X,d,f)$ be a dynamical system with specification property and $\psi$ be a continuous function on $X$. Let
$E_{\alpha}:=\{x\in X:\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha\}$
denote the $\alpha$-level set of $\psi$. If $\alpha\in\mathbb{R}$ is a constant such that $E_{\alpha}\not=\emptyset$, then
$\overline{\rm mdim}_{M}^{B}(f,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(f,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
[…]
Proof.
[…]
$K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
Fix $x\in X$. Since the statement $\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\psi,\alpha)$, this shows $K_{\alpha}=E_{\alpha}$. We get the desired result by Theorem 4.1. ∎
Readers can turn to [O03, Tho09, ZC13] and references therein for more examples about the multifractal analysis of level sets using topological pressure. Then, by Theorem 4.1, one can get analogous variational principles for the metric mean dimensions of level sets of some special functions, which are determined by the dynamical behaviors of the system.
4.2 Linking metric mean dimension and mean Li-Yorke chaotic
In this subsection, we link the concepts of metric mean dimension and mean Li-Yorke chaos in the sense of infinite entropy. We give a quantitative characterization for the chaotic phenomenon of infinite entropy systems. Namely, we show the packing metric mean dimension of the set of mean Li-Yorke pairs is bounded from blow by the metric mean dimension of the phase space (Theorem 4.3).
Let $(X,d,f)$ be a dynamical system. This naturally gives a product system $(X\times X,d\times d,f\times f)$, where $X\times X$ is compact endowed with the product metric $d\times d$ defined by $d\times d((x_{1},y_{1}),(x_{2},y_{2}))=\max\{d(x_{1},x_{2}),d(y_{1},y_{2})\}$, and the map
$f\times f:X\times X\rightarrow X\times X$
given by $(f\times f)(x,y):=(fx,fy)$ is continuous.
It is well-known that if two systems admit the specification property, then the product system also has the specification property(cf. [DGS76, Proposition 21.4]). A pair $(x,y)\in X\times X$ is called a mean Li-Yorke pair if
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)=0,\ \limsup_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)>0.$
The set of mean Li-Yorke pairs is denoted by ${\rm MLY}(f)$. A subset $S$ of $X$ is called a mean Li-Yorke set if any two distinct points in $S$ form a mean Li-Yorke pair. We say that the system $(X,f)$ is mean Li-Yorke chaotic if there is an uncountable mean Li-Yorke set.
Theorem 4.3.
Let $(X,d,f)$ be a dynamical system with specification property. Then
$\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d),$
$\overline{\rm mdim}_{M}^{UC}(f\times f,{\rm MLY}(f),d\times d)=\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d).$
Proof.
Claim: there exists $\mu_{0}\in M_{f\times f}(X\times X)$ with $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}(x,y)>0$.
Proof of the claim.
Specification property implies that the system has positive entropy. Then by the well-known fact that positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]), we can choose two distinct points of $x,y\in X$, a sequence $\{n_{i}\}$ of positive integer that converges to $\infty$ as $i\to\infty$, and $\theta>0$ (depending on $x,y$) such that $\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta$ for all sufficiently large $i$. We put $\mu_{n_{i}}=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$. Then any an accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfies the desired claim. ∎
Now take $\nu\in M_{f}(X)$ and put $\mu_{1}=\phi_{\ast}\nu\in M_{f\times f}(X\times X)$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$. Then one has $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$. Assume that $\overline{\rm mdim}_{M}(f,X,d)$ is finite. For any $\gamma>0$, by Theorem 2.4 we pick $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and $\rho_{k}\in M_{f}(X)$ such that for each $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)>\overline{\rm mdim}_{M}(f,X,d)-\gamma.$ (4.4)
Let $\rho_{k}^{\ast}=\phi_{\ast}\rho_{k}\in M_{f\times f}(X\times X)$. Then for every $k$,
$\inf_{{\rm diam}(\xi,d)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)=\inf_{{\rm diam}(\eta,d\times d)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta).$ (4.5)
Indeed, let $\xi$ be a finite Borel partition of $X$, then $\eta=\{A\times B:A,B\in\xi\}$ is a partition of $X\times X$ so that ${\rm diam}(\eta,d\times d)={\rm diam}(\xi,d)$. For any $n\geq 1$,
$H_{\rho_{k}^{\ast}}(\eta^{n})$ $=-\sum_{\tilde{C}\in\eta^{n}}\rho_{k}^{\ast}(\tilde{C})\log\rho_{k}^{\ast}(\tilde{C})=-\sum_{C\in\xi^{n}}\rho_{k}^{\ast}(C\times C)\log\rho_{k}^{\ast}(C\times C)$
$=-\sum_{C\in\xi^{n}}\rho_{k}(C)\log\rho_{k}(C)=H_{\rho_{k}}(\xi^{n}),$
which implies that $\inf_{{\rm diam}(\eta,d\times d)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)\leq\inf_{{\rm diam}(\xi,d)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)$. Similarly, if $\eta$ is a finite Borel partition of $X\times X$, then $\xi=\{\phi^{-1}(D\cap\Delta):D\in\eta\}$ is a finite partition of $\Delta=\{(x,x):x\in X\}$, such that ${\rm diam}(\xi)\leq{\rm diam}(\eta)$. Moreover, one has
$H_{\rho_{k}}(\xi^{n})=-\sum_{C\in\xi^{n}}\rho_{k}(C)\log\rho_{k}(C)=-\sum_{\tilde{C}\in\eta^{n}}\rho_{k}^{\ast}(\tilde{C})\log\rho_{k}^{\ast}(\tilde{C})=H_{\rho_{k}^{\ast}}(\eta^{n}),$
which implies that $\inf_{{\rm diam}(\eta,d\times d)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)\geq\inf_{{\rm diam}(\xi,d)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)$. This shows (4.5).
Let $K=\overline{{\rm Conv}\{\mu_{0},\mu_{1},\rho_{k}^{\ast},k\geq 1\}}$, i.e. the convex closure of set $\{\mu_{0},\mu_{1}\}\cup\{\rho_{k}^{\ast}:k\geq 1\}$. Then $K$ is a closed convex subset of $M_{f\times f}(X\times X)$. For any $(x_{0},y_{0})\in G_{K}$, there are two sequences of integers $\{t_{n}^{0}\}$ and $\{t_{n}^{1}\}$ with $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$. Hence
$\lim_{n\to\infty}\frac{1}{t_{n}^{0}}\sum_{i=0}^{t_{n}^{0}-1}d(f^{i}x_{0},f^{i}y_{0})=\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}>0,$
and
$\lim_{n\to\infty}\frac{1}{t_{n}^{1}}\sum_{i=0}^{t_{n}^{1}-1}d(f^{i}x_{0},f^{i}y_{0})=\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0.$
This shows $G_{K}\subset{\rm MLY}(f)$. Therefore, by Theorem 1.1, one has
$\overline{\rm mdim}_{M}^{P}(f\times f,{\rm MLY}(f),d\times d)\geq$ $\overline{\rm mdim}_{M}^{P}(f\times f,G_{K},d\times d)$
$=$ $\limsup_{\epsilon\to 0}\frac{\sup_{\nu\in K}\inf_{{\rm diam}(\eta)<\epsilon}h_{\nu}(f\times f,\eta)}{|\log\epsilon|}$
$\geq$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\eta)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)}{|\log\epsilon_{k}|}$
$=$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)}{|\log\epsilon_{k}|}\geq\overline{\rm mdim}_{M}(f,X,d)-\gamma.$
Since $\gamma$ is arbitrary, by Proposition 2.1 one has
$\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d).$
The case of $\overline{\rm mdim}_{M}(f,X,d)=\infty$ is still valid by considering a slight modification.
For sufficiently small $\gamma>0$, let $c_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d)-\gamma\}$. By Theorem 2.4, we pick $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and $m_{k}\in M_{f\times f}(X\times X)$ such that for each $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f,\xi)>c_{\gamma}.$ (4.6)
Put $\hat{K}=\overline{{\rm Conv}\{\mu_{0},\mu_{1},m_{k},k\geq 1\}}$. Then the similar argument gives us
$\overline{\rm mdim}_{M}^{P}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d),$
which implies the desired result by Proposition 2.1.
Finally, we give an example to verify the Corollary 4.2 and Theorem 4.3.
Example 4.4.
Endow the product space $[0,1]^{\mathbb{Z}}$ with the product metric
$d((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}})=\sum_{n\in\mathbb{Z}}\frac{|x_{n}-y_{n}|}{2^{|n|}},$
and let $\sigma:[0,1]^{\mathbb{Z}}\to[0,1]^{\mathbb{Z}}$ be the left shift map defined by $\sigma((x_{n})_{n\in\mathbb{Z}})\mapsto(x_{n+1})_{n\in\mathbb{Z}}$. It is shown that the system $([0,1]^{\mathbb{Z}},d,\sigma)$ has the specification property (cf. [DGS76, Proposition 21.2]), and ${\rm{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d)=1$ [LT18].
  • (1)
    Thus, by the product formula of metric mean dimension one has ${\rm mdim}_{M}(\sigma\times\sigma,[0,1]^{\mathbb{Z}}\times[0,1]^{\mathbb{Z}},d\times d)=2$. Consequently, by Theorem 4.3, we have
    $\overline{\rm mdim}_{M}^{UC}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=2.$
  • (2)
    Let $\varphi:[0,1]^{\mathbb{Z}}\to\mathbb{R}$ be map defined by $\varphi((x_{n})_{n\in\mathbb{Z}})=x_{0}$. Then for any $\alpha\in[0,1]$, we have
    $\overline{\rm mdim}_{M}^{B}(\sigma,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(\sigma,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{\sigma}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi),$
[…]
5 Examples
Applying the main results, we give the specific formulae of the metric mean dimensions of generic points of full shift over compact metric space.
Example 5.1.
Consider the Example 4.4 again. Let $\mathcal{L}$ be the Lebesgue measure on $[0,1]$ and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ be the product measure, which is a $\sigma$-ergodic measure on $[0,1]^{\mathbb{Z}}$. Then the ergodicity of $\mu$ enables us to define the upper $\epsilon$-Brin-Katok local entropy of $\mu$ as
$h_{\mu}^{BK}(\epsilon):=\limsup_{n\to\infty}-\frac{1}{n}\log\mu(B_{n}(x,\epsilon)),\ \mu\text{-a.e. }x\in[0,1]^{\mathbb{Z}}.$
Shi proved that $\lim_{\epsilon\to 0}\frac{h_{\mu}^{BK}(\epsilon)}{|\log\epsilon|}=1$ [S22, Example 12]. Together with [YCZ24, Theorem 1.1], we have
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi)=1={\rm\overline{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d).$
Then, by Corollary 1.5, we have
${\rm\overline{mdim}}_{M}^{B}(\sigma,G_{\mu},d)={\rm\overline{mdim}}_{M}^{P}(\sigma,G_{\mu},d)={\rm\overline{mdim}}_{M}^{UC}(\sigma,G_{\mu},d)=1.$
We extend the Example 5.1 to more general full shift.
Example 5.2.
Let $K$ be a compact metric space with a metric $d$. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with following product metric
$d^{\mathbb{Z}}\left((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}}\right)=\sum_{n\in\mathbb{Z}}\frac{d(x_{n},y_{n})}{2^{|n|}}.$
For $n\geq 0$, let $\pi_{n}$ be the projection from $K^{\mathbb{Z}}$ to its first $n$ coordinates.
It was proved in [S22, Theorem 15]
$\overline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K),$
where $\overline{\rm dim}_{B}(K)$ is the upper box dimension of $K$ under the metric $d$.
Theorem 5.3.
Let $K$ be a compact metric space with a metric $d$. Then there exists $\mu\in M_{\sigma}(K^{\mathbb{Z}})$ such that
$\overline{\rm mdim}_{M}^{B}(\sigma,G_{\mu},d^{\mathbb{Z}})$ $=\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K)$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi).$
Proof.
It is known that the system $(K^{\mathbb{Z}},\sigma)$ has specification property. Let $S_{\epsilon}$ be a maximal $\epsilon$-separated subset of $K$ with respect to metric $d$. For $n\geq 1$, we put
${\mathcal{S}_{n}}=\{x\in K^{n}:x_{k}\in S_{\epsilon},\forall~{}0\leq k\leq n-1\},$
and consider the set $\hat{\mathcal{S}}_{n}\subset K^{\mathbb{Z}}$ consisting of one preimage of each element in ${\mathcal{S}_{n}}$ under $\pi_{n}$. Then $\hat{\mathcal{S}}_{n}$ is an $(n,\epsilon)$-separated set of $K^{\mathbb{Z}}$ and #$\#{\hat{\mathcal{S}}_{n}}=(\#S_{\epsilon})^{n}$. Define
$\mu_{n}:=\frac{1}{\#\hat{\mathcal{S}}_{n}}\sum_{x\in\hat{\mathcal{S}}_{n}}\mathcal{E}_{n}(x).$
We assume that $\mu_{n}\rightarrow\mu_{\epsilon}\in M_{\sigma}(K^{\mathbb{Z}})$. Then, by Lemma 3.1, for each $\epsilon>0$ we have
$\log\#S_{\epsilon}=\limsup_{n\to\infty}\frac{1}{n}\log\#\hat{\mathcal{S}}_{n}\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(\sigma,\xi).$
Assume that $\mu_{\epsilon}\rightarrow\mu\in M_{\sigma}(K^{\mathbb{Z}})$ as $\epsilon\to 0$. Combining with the Theorem 1.3, one has
$\overline{\rm dim}_{B}(K)=\limsup_{\epsilon\to 0}\frac{\log\#S_{\epsilon}}{|\log\epsilon|}\leq\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})\leq\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})\leq\overline{\rm dim}_{B}(K).$
Finally, by the Corollary 1.5 we complete the proof.
Availability of data and materials: Not applicable.
Competing interests: The authors declare that they have no competing interests.
Acknowledgments: The first and second authors were supported by NNSF of China (12101340). The third author was supported by the China Postdoctoral Science Foundation (Grant Number: 2024M763856); R.Yang would like to thank Yunping Wang for many helpful communication and the warm hospitality when he was visiting the Ningbo University. The authors also thank Prof. Ercai Chen for many useful comments.