### v3

# Yong Ji1, Junye Li1 and Rui Yang∗2,3

Abstract. Saturated sets and its reduced case, the set of generic points, constitute two significant types of fractal-like sets in multifractal analysis of dynamical systems. In the context of infinite entropy systems, we aim to provide some qualitative aspects of saturated sets and the set of generic points from both topological and measure-theoretic perspectives.
For systems with the specification property, we establish certain variational principles for saturated sets in terms of Bowen and packing metric mean dimensions, and show the upper capacity metric mean dimension of saturated sets has full metric mean dimension. As applications, we further present some qualitative aspects of the metric mean dimensions of level sets and the set of mean Li-Yorke pairs in infinite entropy systems.
Keywords and phrases: Metric mean dimension; Saturated set; Generic point; Variational principle; Multifractal analysis
1 Introduction
Let $(X,d,f)$ be a topological dynamical system (t.d.s. for short), where $(X,d)$ is a compact metric space with a metric $d$, and $f:X\to X$ is a homeomorphism. Let $M(X)$, $M_{f}(X)$, $M_{f}^{e}(X)$ denote the sets of Borel probability measures endowed with weak-topology, $f$-invariant Borel probability measures, and ergodic $f$-invariant Borel probability measures on $X$, respectively.
Quantitative analysis using some entropy-like quantities is of great significance in understanding the dynamics of abstract dynamical systems. Measure-theoretic entropy and topological entropy are two crucial invariants that helps us understand the topological complexity of dynamical systems, which are related by the classical variational principle:
$h_{top}(f)=\sup_{\mu\in M_{f}(X)}{h_{\mu}(f)},$
where $h_{top}(f)$ is the topological entropy of $X$, $h_{\mu}(f)$ is the measure-theoretic entropy of $\mu$, and the supremum ranges over all $f$-invariant measures of $X$. In 1973, resembling the definition of Hausdorff dimension Bowen [Bow73] introduced Bowen topological entropy for any subset of the phase space $X$. The pioneering work has directly contributed to the development of dimension theory in dynamical systems, especially serving as a powerful tool for studying the “size” of fractal-like sets in multifractal analysis. The multifractal analysis, with its origin dating back to Besicovitch, takes into account the decomposition of the whole phase space into smaller subsets consisting of points with the similar dynamical behaviors(or known as multifractal spectrum), and then, from a geometrical or topological perspective, describes the “size” of those subsets by using fractal-like dimensions and dimensional-like entropies. Such a procedure allows us to get more information about the dynamics and even recovers the partial dynamical behaviors of the system, a phenomenon known as multifractal rigidity.
We aim to study the multifractal analysis of saturated sets in infinite entropy systems. Let us mention some relevant results about the dimensional entropies of saturated sets to clarify the motivation. Given $x\in X$, by
$\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}\in M(X),$
where $\delta_{x}$ is the Dirac mass at $x$, we denote the $n$-th empirical measure of $x$. Then, in the sense of weak-topology, the limit-point set $V_{f}(x)$ of the sequence $\{\mathcal{E}_{n}(x)\}_{n\geq 1}$ is a nonempty compact connected subset of $M_{f}(X)$ (cf. [DGS76, Proposition 3.8]). Specially, when $V_{f}(x)=\{\mu\}$ for some invariant measure $\mu$, then $x$ is called a generic point of $\mu$. Denote by $G_{\mu}$ the set of all generic points of $\mu$. The Birkhoff’s ergodic theorem implies that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic. In 1973, Bowen [Bow73] proved that for any ergodic measure $\mu$, the Bowen topological entropy of $G_{\mu}$ equals $h_{\mu}(f)$, i.e.,
$h_{top}^{B}(f,G_{\mu})=h_{\mu}(f).$
For the case of non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$, and there even exist the examples showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$. Therefore, Bowen’s result fails for non-ergodic measures. It is natural to ask whether there are some proper conditions such that the Bowen’s equality still holds for non-ergodic measures.
A positive answer to the above question is due to Pfister and Sullivan [PS07], who provided it within a more general framework. A subset $K$ of $X$ is called a saturated set if $x\in K$ and $V_{f}(x)=V_{f}(y)$, then $y\in K$. More generally, given a nonempty closed connected subset $K\subset M_{f}(X)$, we have a saturated set by defining $G_{K}=\{x\in X:V_{f}(x)=K\}$ associated with $K$. For any topological dynamical system (t.d.s) possessing the specification property, Sigmund [Sig72, Sig74] firstly proved the existence of the saturated sets, and then was later extended to non-uniformly hyperbolic and non-uniformly expanding systems by several authors [LST13, TV17]. In 2007, Pfister and Sullivan [PS07] introduced the $g$-almost product property that is strictly weaker than specification property, which solely requires a point partially $\epsilon$-shadowing the given finite pieces of orbit-segments and the uniform separation property (e.g. expansive and more generally asymptotically $h$-expansive systems), and established the variational principle of Bowen topological entropy of nonempty compact connected set $K\subset M_{f}(X)$:
$h_{top}^{B}(T,G_{K})=\inf\{h_{\mu}(f):\mu\in K\}.$
Furthermore, under only the assumption of $g$-almost product property, they [PS07] showed that for all $\mu\in M_{f}(X)$,
$h_{top}^{B}(f,G_{\mu})=h_{\mu}(f),$
which extends Bowen’s remarkable result to non-ergodic measures. Consequently, the work of [Bow73, PS07] leads to an intensive study toward the estimation of different types of dimensional-like entropies of saturated sets. Under the specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] also obtained the variational principle of packing topological entropy of nonempty compact connected set $K\subset M_{f}(X)$,
$h_{top}^{P}(f,G_{K})=\sup\{h_{\mu}(f):\mu\in K\}.$
By weakening the specification property and considering systems with the $g$-almost product property, Hou, Tian and Zhang [HTZ23] not only demonstrated that the variational principle $h_{top}^{P}(f,G_{K}^{C})=\sup\{h_{\mu}(f):\mu\in K\}$ holds for (mixed) saturated sets with more complicated dynamics, but also showed that the upper capacity topological entropy of the saturated sets has full topological entropy, i.e.,
$h_{top}^{UC}(f,G_{K}^{C})=h_{top}(f),$
where $G_{K}^{C}:=G_{K}\cap\{x\in X:C_{f}(X)\subset\omega_{f}(x)\},$ $\omega_{f}(x):=\bigcap_{n=0}^{\infty}\overline{\bigcup_{k=n}^{\infty}\{f^{k}x\}}$ is the $\omega$-limit set of $x$, and $C_{f}(X)$ is the measure center given by $C_{f}(X):=\overline{\cup_{\mu\in M_{f}(X)}{\rm supp}{(\mu)}}$, ${\rm supp}(\mu)$ is the supported set of $\mu$. Additionally, Huang, Tian and Wang [HTW19] showed that Pfister and Sullivan’s variational principle is still valid for transitive-saturated set $G_{K}^{T}:=G_{K}\cap{\rm Tran}$, where ${\rm Tran}$ means the set of transitive points. The generalizations of the aforementioned results can be found in [JCZ22, LLS17] for the scenarios of nonuniformly hyperbolic systems, and in [ZC18, DZZ23, Z18, RTZ23, ZRZ23] for the cases of the actions of amenable groups
Based on the possible values of topological entropy, dynamical systems can be can be classified into three types: zero entropy systems, finite positive entropy systems and infinite entropy systems. Infinite entropy systems typically imply that these systems may exhibit rather complicated topological complexity. As a result, the concept of topological entropy fails to offer more information about the dynamics of the system. In 2000, Lindenstrauss and Weiss [LW00] introduced a refined notion for topological entropy, called metric mean dimension, which is a dynamical analogue of box dimension to measure how fast the $\epsilon$-topological entropy diverges to $\infty$ as $\epsilon$ goes to 00. It has been shown that every system with finite entropy has a metric mean dimension of zero. See some applications of metric mean dimension in the analog compression of information theory [GS20], and in estimating the upper bound of mean dimension of some complex systems [Tsu18a, Tsu18b]. The notion of measure-theoretic counterpart of metric mean dimension, called upper and lower mean Rényi information dimensions, is due to Gutman and Śpiewak[GS21]. More precisely, the lower and upper mean Rényi information dimensions of $\mu\in M_{f}(X)$ are respectively defined by
$\underline{\rm MRID}_{\mu}(f,X,d):$ $=\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
$\overline{\rm MRID}_{\mu}(f,X,d):$ $=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where the infimum ranges over the finite Borel measurable partition with diameter at most $\epsilon$. This measures how fast the Kolmogorov-Sinai $\epsilon$-entropy $\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ diverges to $\infty$ as $\epsilon\to 0$. Several variational principles regarding the metric mean dimension of $\mathbb{Z}$-actions can be found in the works [LW00, GS21, S22, YCZ23a, YCZ25]. Then, in the context of the infinite entropy systems, the aforementioned work suggests the following questions for saturated sets:
Questions:
  • (1)
    For ergodic measures, does the metric mean dimension of the set of generic points coincide with the mean Rényi information dimension?
  • (2)
    For non-ergodic measures, under the assumption of specification-like property, are there certain variational principles for metric mean dimensions of saturated sets? If not, can metric mean dimensions of saturated sets have the full metric mean dimension?
For (1), the authors in [YCZ25] showed that the packing metric mean dimension of the set of generic points is equal to the mean Rényi information dimension. However, this equality holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula within the context of infinite entropy systems (cf.[YCZ25, §5, Question,(1)]); for (2), since the uniform separation property implies the system has finite topological entropy, extending the work of [PS07] to the framework of infinite entropy systems encounters some nontrivial challenges. Under the assumption of $g$-almost product property, for any non-empty compact connected set $K$ contained in the convex closure of finite invariant measures, Yuan [Y23] investigated the Bowen metric mean dimension of $G_{K}$.
Prior to the comparison of the different types of specification-like property, we concentrate solely on the systems possessing the specification property and provide a comprehensive answer to question (2). Actually, since the asymptotic behavior of orbits ${\rm orb}_{f}(x)=\{f^{n}(x):n\geq 0\}$ is a fundamental topic in ergodic theory, the subsequent statements of main results involve more general (mixed) saturated sets. For simplicity, we write
$G_{K}^{\omega,U}:=G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},$
where $U\subset X$ is a nonempty open subset, $K\subset M_{f}(X)$ is a nonempty closed connected (or convex) subset. If $G_{K}^{\omega,U}\not=\emptyset$, then there exists a transitive point in the open set $U$ such that $V_{f}(x)=K$.
Saturated sets exhibit different dynamical behaviors in terms of different types of metric mean dimensions. There are indeed certain variational principles for the Bowen and packing metric mean dimensions of saturated sets, as presented in Theorems 1.1, 1.2 and 1.3. Although the $\mu$-measure of $G_{\mu}$ is zero for every $\mu\in M_{f}(X)$, its topological complexity can be sufficiently large, that is, the upper capacity metric mean dimension of saturated sets carries the full metric mean dimension, as stated in Theorem 1.4.
Theorem 1.1.
Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, and $U\subset X$ is a nonempty open set, then
$\overline{\rm mdim}_{M}^{P}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where $\overline{\rm mdim}_{M}^{P}(f,Z,d)$ is the packing upper metric mean dimension of $Z$, $h_{\mu}(f,\xi)$ is the measure-theoretic entropy of $\mu$ w.r.t. $\xi$.
Theorem 1.2.
Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact connected set, and $U\subset X$ is a nonempty open set, then
$\overline{\rm mdim}_{M}^{B}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where $\overline{\rm mdim}_{M}^{B}(f,Z,d)$ is the Bowen upper metric mean dimension of $Z$.
Perhaps, more satisfactory results would be to exchange the order of $\limsup_{\epsilon\to 0}$ and $\sup_{\mu\in K}$ (or $\inf_{\mu\in K}$). However, up to now, the authors remain unaware of whether there exist any counterexamples among infinite entropy systems with the specification property for which the following equalities hold strictly
$\sup_{\mu\in K}\left\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\right\}<\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d),$
and
$\inf_{\mu\in K}\left\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\right\}<\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d).$
Inspired by the work of [GS21, CPV24], we define two variants of mean Rényi information dimension of invariant measures. Let $K$ be a closed set of $M_{f}(X)$. For any $\mu\in K$, we use $M_{K}(\mu)$ to denote the space of sequences of invariant measures within $K$ which converge to $\mu$ in the weak-topology. We define the (modified) lower and upper mean Rényi information dimensions of $\mu$ as
$\underline{\rm mrid}_{\mu}(f,K,d):$ $=\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\},$
$\overline{\rm mrid}_{\mu}(f,K,d):$ $=\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}$
respectively. Then this allows us to obtain the the variational principles in terms of the (modified) lower and upper mean Rényi information dimensions.
Theorem 1.3.
Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, then
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\},$
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
We remark that Theorem 1.3 does not require the connectedness of $K$ since any convex set in a topological vector space is automatically connected.
Theorem 1.4.
Let $(X,d,f)$ be a t.d.s. with specification property. Let $K\subseteq M_{f}(X)$ be a nonempty compact connected set, and $U\subset X$ be a nonempty open set. Then
$\overline{\rm mdim}_{M}^{UC}\left(f,G_{K},d\right)=\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}(f,X,d),$
where $\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ is the upper capacity metric mean dimension of $Z$.
As an immediate consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case occurs when $K$ consists of exactly a single point. The Corollary 1.5 extends the result given in [YCZ25, Theorem 1.1] to all non-ergodic measures for systems with specification property.
Corollary 1.5.
Let $(X,d,f)$ be a dynamical system with specification property. Then for every $\mu\in M_{f}(X)$,
$\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}^{P}_{M}(f,G_{\mu},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi),$
$\overline{\rm mdim}_{M}^{UC}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
Consequently, if $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure in the sense of $\overline{\rm mdim}_{M}(f,X,d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$, then we have
$\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}^{P}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}^{UC}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
The “Consequently” part can be realized for some dynamical systems. For example, the full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$ with the compatible product metric $d(x,y)=\sum_{n\in\mathbb{Z}}\frac{||x_{n}-y_{n}||_{\mathbb{R}^{m}}}{2^{|n|}}$, and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ is product measure of the Lebesgue measure $\mathcal{L}$ on $[0,1]^{m}$.
The concept of maximal metric mean dimension measure was proposed in [YCZ23a]. However, determining what types of infinite entropy systems can possess a maximal metric mean dimension measure remains a challenging task. See [LR24] for the existence of maximal metric mean dimension measure for typical conservative homeomorphisms. By Gutman- Śpiewak variational principle (as presented in Theorem 2.4) and Corollary 1.5, we have an abstract criterion for maximal metric mean dimension measure for identifying maximal metric mean dimension measures in systems with the specification property.
Corollary 1.6.
Let $(X,d,f)$ be a t.d.s. with specification property. An invariant measure $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure if and only if $\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
Some comments are in order.
  • (1)
    There are plenty of infinite entropy systems with specification property, for instance, the usual full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$, and the examples presented in [BR23, Section IV], and our main results can be applied to these systems.
  • (2)
    In Theorem 1.3, we do not know whether the equalities hold for Bowen upper metric mean dimension and packing lower metric mean dimension. The other theorems mentioned above are valid for the corresponding lower metric mean dimensions by considering changing the $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.
  • (3)
    We believe that it is trivial to extend the aforementioned results to metric mean dimensions with potential by making some modifications to the current proofs. The novelty of our work not only gives a proper and full extension of the previous work in [PS07, ZCC12, ZC13, HTW19] to the setting of infinite entropy systems of $\mathbb{Z}$-actions with specification property, but also may be helpful for studying the multifractal analysis of other types of fractal-like sets in infinite entropy systems. It remains unclear whether there are some possible substitutions of the uniform separation property so that all the above results are true for $g$-almost product property, or even for cases beyond $\mathbb{Z}$.
The paper is organized as follows. In section 2, we recall the precise definitions of metric mean dimensions and the associated properties, and two types of measure-theoretic $\epsilon$-entropies, and the specification property. In section 3, we prove the main results. In section 4, we consider some applications of the main results, and apply it to investigate the level sets of Birkhoff average of a continuous potential, Li-Yorke chaos and the metric mean dimensions of generic points of the full shift over compact metric spaces.
2 Preliminaries
In this section, we recall the concepts of metric mean dimensions in both topological and measure-theoretic situations, and state the precise definition of specification property in dynamical systems.
2.1 Three types of metric mean dimensions
In this subsection, we review the definitions of upper capacity metric mean dimension of subsets[LW00], Bowen metric mean dimension[LV21, W21] and packing metric mean dimension [YCZ22] of subsets.
Let $(X,d,f)$ be a t.d.s.. For any natural number $n\in\mathbb{N}$, we define the $n$-th Bowen metric $d_{n}$ on the space $X$ as
$d_{n}(x,y):=\max_{0\leq i\leq n-1}\{d(f^{i}x,f^{i}y)\},$
where $x,y\in X$. Then the Bowen open ball and closed ball centered at $x$ with radius $\epsilon$ in the metric $d_{n}$ are given by
$B_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)<\epsilon\},$
$\overline{B}_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)\leq\epsilon\}$
respectively.
Let $Z$ be a nonempty subset of $X$. Fix $\epsilon>0$. A set $E\subset Z$ is said to be an $(n,\epsilon)$-separated set of $Z$ if any distinct $x,y\in E$ implies $d_{n}(x,y)>\epsilon$. Denote by $s_{n}(Z,\epsilon)$ the largest cardinality of $(n,\epsilon)$-separated sets of $Z$. The $\epsilon$-upper capacity topological entropy of $Z$ is given by
$h_{top}^{UC}(f,Z,\epsilon)=\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(Z,\epsilon).$
Recall that the upper capacity topological entropy of $Z$ is the limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 00, i.e.,
$h_{top}^{UC}(f,Z)=\lim_{\epsilon\to 0}h_{top}^{UC}(f,Z,\epsilon)=\sup_{\epsilon>0}h_{top}^{UC}(f,Z,\epsilon).$
Specifically, topological entropy fails to distinguish different levels of topological complexity within the infinite entropy systems
Analogous to the definition of box dimension in fractal geometry, we formulate the refined entropy-like quantity for infinite entropy systems using $\epsilon$-upper capacity topological entropy.
Definition 2.1.
The upper capacity metric mean dimension of $Z$ is defined by
$\overline{\rm mdim}_{M}^{UC}(f,Z,d):=\limsup_{\epsilon\to 0}\frac{h_{top}^{UC}(f,Z,\epsilon)}{|\log\epsilon|}.$
In the case where $Z$ is equal to $X$, we use the notation $h_{top}(f,X,\epsilon)$ instead of $h_{top}^{UC}(f,X,\epsilon)$, and $\overline{\rm mdim}_{M}(f,X,d)$ instead of $\overline{\rm mdim}_{M}^{UC}(f,X,d)$ respectively. Then the upper capacity metric mean dimension of $X$ is reduced to Lindenstrauss-Weiss’s definition introduced in [LW00]. Such a formulation extends the concept of the metric mean dimension to any subsets of the phase space, which is not necessarily compact or $f$-invariant.
Notice that the (upper capacity) metric mean dimension is defined by detecting how fast the $\epsilon$-upper capacity topological entropy diverges to $\infty$ as $\epsilon$ tends to 00. That is, for sufficiently small $\epsilon>0$, we may think of
$h_{top}^{UC}(f,Z,\epsilon)\approx\overline{\rm mdim}_{M}^{UC}(f,Z,d)\cdot|\log\epsilon|.$
It is straightforward to observe that any system with finite topological entropy has zero metric mean dimension. This, in turn, shows that the metric mean dimension is a suitable “measure” for capturing the dynamics of the infinite entropy systems.
The classical Bowen topological entropy[Bow73] and packing topological entropy[FH12] in dimension theory, which are defined by Carathéodory-Pesin structures, are the analogue of the Hausdorff dimension and packing dimension in fractal geometry. The corresponding dimensional characterizations of metric mean dimensions, which have been introduced in [LV21, YCZ22, W21] through Carathéodory-Pesin structures, are the Bowen and packing metric mean dimensions. One may refer to [BR23, CLS21, LL24, FKO24, YCZ25] for the results involving Bowen and packing metric mean dimensions.
For $Z\subset X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
$M(Z,s,N,\epsilon)=\inf\left\{\sum_{i\in I}\exp(-n_{i}s)\right\},$
where the infimum is taken over all finite or countable families $\{B_{n_{i}}(x_{i},\epsilon)\}_{i\in I}$ such that $Z\subset\cup_{i\in I}B_{n_{i}}(x_{i},\epsilon)$ with $x_{i}\in X,n_{i}\geq N$.
The quantity $M(Z,s,N,\epsilon)$ is non-decreasing as $N$ increases, this ensures that the limit
$M(Z,s,\epsilon)=\lim_{N\to\infty}M(Z,s,N,\epsilon)$
exists. There is a critical value of the parameter $s$, which we call $\epsilon$-Bowen topological entropy, such that $M(Z,s,\epsilon)$ jumps from $\infty$ to 00. That is, the critical value is given by
$h_{top}^{B}(f,Z,\epsilon)=\inf\{s>0:M(Z,s,\epsilon)=0\}=\sup\{s>0:M(Z,s,\epsilon)=\infty\}.$
Definition 2.2.
The Bowen upper metric mean dimension of $Z$ is defined by
$\overline{\rm mdim}^{B}_{M}(f,Z,d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{B}(f,Z,\epsilon)}{|\log\epsilon|}.$
For $Z\subseteq X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
$P(Z,s,N,\epsilon)=\sup\left\{\sum_{i\in I}\exp(-sn_{i})\right\},$
where the supremum is taken over all finite or countable pairwise disjoint families $\{\overline{B}_{n_{i}}(x_{i},\epsilon)\}_{i\in I}$ with $x_{i}\in Z$, $n_{i}\geq N$ for all $i\in I$.
Since the quantity $P(Z,s,N,\epsilon)$ is non-increasing as $N$ increases, and thus the following limit exists:
$P(Z,s,\epsilon)=\lim_{N\to\infty}P(Z,s,N,\epsilon).$
Observe that the quantity $P(Z,s,\epsilon)$ does not satisfy the countable additive property with respect to $Z$. So some modifications are needed to obtain the property:
$\mathcal{P}(Z,s,\epsilon):=\inf\left\{\sum_{i=1}^{\infty}P(Z_{i},s,\epsilon):\bigcup_{i=1}^{\infty}Z_{i}\supseteq Z\right\}.$
There is a critical value of the parameter $s$, which we call $\epsilon$-packing topological entropy, such that $\mathcal{P}(Z,s,\epsilon)$ jumps from $\infty$ to 00. That is, the critical value is given by
$h_{top}^{P}(f,Z,\epsilon)$ $=\inf\{s>0:\mathcal{P}(Z,s,\epsilon)=0\}$
$=\sup\{s>0:\mathcal{P}(Z,s,\epsilon)=\infty\}.$
Definition 2.3.
The packing upper metric mean dimension of $Z$ is defined by
$\overline{\rm mdim}_{M}^{P}(f,Z,d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{P}(f,Z,\epsilon)}{|\log\epsilon|}.$
One can replace $\limsup_{\epsilon\to 0}$ by $\liminf_{\epsilon\to 0}$ in above definitions to obtain the corresponding three types of lower metric mean dimensions on subsets, which are denoted by $\underline{\rm mdim}_{M}^{S}(f,Z,d)$ with $S\in\{UC,B,P\}$. If $\underline{\rm mdim}_{M}^{S}(f,Z,d)=\overline{\rm mdim}_{M}^{S}(f,Z,d)$, we call the common value ${\rm mdim}_{M}^{S}(f,Z,d)$ the (capacity/Bowen/packing) metric mean dimension of $Z$. We summarize some fundamental properties of metric mean dimensions derived in [YCZ22, Proposition 3.4] and [YCZ23b, Proposition 2.6].
Proposition 2.1.
$(1)$ If $Z_{1}\subset Z_{2}\subset X$, then for each $S\in\{UC,B,P\}$,
$\overline{\rm mdim}_{M}^{S}(f,Z_{1},d)\leq\overline{\rm mdim}_{M}^{S}(f,Z_{2},d).$
$(2)$ If $Z=\bigcup_{i=1}^{\infty}Z_{i}$, then for every $\epsilon>0$,
$h_{top}^{B}(f,Z,\epsilon)$ $=\sup_{i\geq 1}h_{top}^{B}(f,Z_{i},\epsilon)$
$h_{top}^{P}(f,Z,\epsilon)$ $=\sup_{i\geq 1}h_{top}^{P}(f,Z_{i},\epsilon),$
and $h_{top}^{UC}(f,Z,\epsilon)=\max_{1\leq i\leq N}h_{top}^{UC}(f,Z_{i},\epsilon)$ if $Z=\bigcup_{i=1}^{N}Z_{i}$.
$(3)$ If $Z=\bigcup_{i=1}^{N}Z_{i}$, then for each $S\in\{UC,B,P\}$,
$\overline{\rm mdim}_{M}^{S}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}_{M}^{S}(f,Z_{i},d).$
$(3)$ If $Z$ is a nonempty subset of $X$, then for every $10<\epsilon<$ 1, we have
$h_{top}^{B}(f,Z,3\epsilon)\leq h_{top}^{P}(f,Z,\epsilon)\leq h_{top}^{UC}(f,Z,\epsilon).$
Consequently, $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)\leq\overline{\rm mdim}_{M}^{UC}(f,Z,d)$. Furthermore, if $Z$ is $f$-invariant and compact, then the three types of metric mean dimensions are equal.
Given a subset of phase space, it is not easy to calculate its precise Bowen and packing metric mean dimensions. However, in the multifractal analysis of dynamical systems the Entropy Distribution Principles allow us to get a lower bound for $\epsilon$-Bowen and $\epsilon$-packing topological entropies of certain fractal-like sets. Consequently, the powerful tool yields the lower bounds of metric mean dimension of the fractal-like sets. $\epsilon$-Bowen and packing topological entropies of Entropy Distribution Principles are stated as follows:
Lemma 2.1.
[BR23, Lemma 13] Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Suppose that $Z$ is a Borel set of $X$ and $s\geq 0$. If there exist a $\mu\in M(X)$ and a constant $C>0$ such that $\mu(Z)>0$ and for every $n$ and for each Bowen ball $B_{n}(x,\epsilon)$ with $B_{n}(x,\epsilon)\cap Z\neq\emptyset$, one has $\mu(B_{n}(x,\epsilon))\leq Ce^{-ns}$, then $h_{top}^{B}(f,Z,\epsilon)\geq s$.
Lemma 2.2.
Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Suppose that $Z$ is a Borel set of $X$ and $s\geq 0$. If there exist a $\mu\in M(X)$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ that converges to $\infty$ satisfying $\mu(Z)>0$ and $\mu(B_{n_{i}}(x,\epsilon))\leq Ce^{-n_{i}s}$ for each $i$ and for any $x\in Z$, then $h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s$.
Proof.
By definition, it is enough to prove that for sufficiently small $\delta>0$,
$P\left(E,s-\delta,\frac{\epsilon}{6}\right)=\infty$
holds for any Borel subset $E\subset Z$ with $\mu(E)>0$.
For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of the set $E$. Then, by $5r$-covering lemma stated in [Mat95, Theorem 2.1], there is a countable pairwise disjoint subcollection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, where $J$ is a index set, and in particular the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint such that $E\subset\bigcup_{j\in J}B_{n_{i}}(x_{j},\epsilon).$ Hence, one has
$P\left(E,s-\delta,n_{i},\frac{\epsilon}{6}\right)\geq\sum_{j\in J}\exp[-n_{i}(s-\delta)]\geq\frac{e^{n_{i}\delta}}{C}\sum_{j\in J}\mu(B_{n_{i}}(x_{j},\epsilon))\geq\frac{e^{n_{i}\delta}}{C}\mu(E).$
Letting $i\to\infty$, we have $P(E,s-\delta,\frac{\epsilon}{6})=\infty$. This shows that $\mathcal{P}(Z,s-\delta,\frac{\epsilon}{6})=\infty$, and hence $h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s$ by letting $\delta\to 0$. ∎
2.2 Measure-theoretic $\epsilon$-entropy of invariant measures
In this subsection, we recall the definitions of Kolmogorov-Sinai $\epsilon$-entropy and Katok $\epsilon$-entropy of invariant measures.
Let $C(X)$ denote the space consisting of real-valued continuous functions defined on $X$, equipped with the supremum norm. By virtue of the Riesz representation theorem for compact metric spaces, we are able to endow $M(X)$ with the weak-topology, which is metrizable by the following metric:
$D(\mu,\nu):=\sum_{k=1}^{\infty}\frac{|\int\phi_{k}\mathop{}\!\mathrm{d}\mu-\int\phi_{k}\mathop{}\!\mathrm{d}\nu|}{2^{k}},$ (2.1)
where $\{\phi_{k}\}$ is a countable family of dense subsets of $C(X)$ taking values in $[0,1]$. Thus, the diameter of $M(X)$ in the sense of the metric $D$ is always less than $1$, and for fixed $\mu,\nu\in M(X)$, both $D(\mu,\cdot)$ and $D(\cdot,\nu)$ are convex functions on $M(X)$. It follows from the weak-topology on $M(X)$ that the natural embedding from $X$ to $M(X)$, given by $x\mapsto\delta_{x}$, is continuous. Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in the sense of the weak-topology.
Given an invariant measure $\mu\in M_{f}(X)$, a finite family $\xi=\left\{C_{1},\cdots,C_{k}\right\}$ of Borel subsets of $X$ is a measurable partition of $X$ if $\mu(C_{i}\cap C_{j})=0$ for all $i\neq j$ and $\mu(X\setminus\cup_{i=1}^{k}C_{i})=0$. Then the partition entropy of $\xi$ w.r.t. $\mu$ is defined by
$H_{\mu}(\xi)=-\sum_{i=1}^{k}\mu(C_{i})\log(\mu(C_{i})).$
We define the measurable partition of $\mu\in M_{f}(X)$ w.r.t. $\xi$ as
$h_{\mu}(f,\xi)=\lim_{n\to+\infty}\frac{1}{n}H_{\mu}(\xi^{n}),$
where $\xi^{n}:=\vee_{i=0}^{n-1}f^{-i}\xi$ is the $n$-th join of the $n$ partitions $\xi,f^{-1}\xi,\ldots,f^{-(n-1)}\xi$ on $X$. The measure-theoretic entropy of $\mu$ is defined as
$h_{\mu}(f)=\sup_{\xi}h_{\mu}(f,\xi),$
where the supremum is taken over all finite Borel measurable partitions of $X$.
For topological dynamical systems, the measure-theoretic counterpart of metric mean dimensions is the measure-theoretic metric mean dimensions of invariant measures by measuring the divergent rates of the limits of the measure-theoretic $\epsilon$-entropies such as Brin-Katok $\epsilon$-entropy, Katok $\epsilon$-entropy, Kolmogorov-Sinai $\epsilon$-entropy, and the others. A systemic treatment about the measure-theoretic $\epsilon$-entropies is stated in [YCZ25], and the divergent rates of different types of measure-theoretic $\epsilon$-entropies coincide in terms of ergodic measures (cf. [YCZ25, Theorem 1.1]). Here, we only recall two types of the measure-theoretic $\epsilon$-entropies.
  • (a)
    The Kolmogorov-Sinai $\epsilon$-entropy (cf. [GS21]) of $\mu\in M_{f}(X)$ is given by
    $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
  • (b)
    For $\mu\in M_{f}(X)$, $\epsilon>0$, $\delta\in(0,1)$ and $n\in\mathbb{N}$, we put
    $N_{\mu}(n,\epsilon,\delta):=\min\{\#E:E\subset X~{}\emph{and}~{}\mu(\cup_{x\in E}B_{n}(x,\epsilon))>1-\delta\}.$
    The lower Katok $\epsilon$-entropy of $\mu$ is given by
    $\underline{h}_{\mu}^{K}(f,\epsilon,\delta)=\liminf_{n\to\infty}\frac{1}{n}\log N_{\mu}(n,\epsilon,\delta).$
The Katok $\epsilon$-entropy and Kolmogorov-Sinai $\epsilon$-entropy are related by the following inequalities:
Lemma 2.3.
[S22, Lemma 4.1] Let $\mu\in M_{f}^{e}(X)$ and $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 4\epsilon$ and ${\rm Leb}(\mathcal{U})\geq\epsilon$. Then for any $\delta\in(0,1)$, one has
$\underline{h}_{\mu}^{K}(f,4\epsilon,\delta)\leq\inf_{\xi\succ\mathcal{U}}h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta).$
Consequently, $\underline{h}_{\mu}^{K}(f,16\epsilon,\delta)\leq\inf_{{\rm diam}(\xi)<4\epsilon}\limits h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta).$
The following variational principles verify the relations between the metric mean dimension and the two types of measure-theoretic $\epsilon$-entropies.
Theorem 2.4.
(cf.[GS21, Theorem 3.1, Remark 3.6], [S22, Proposition 7.3]) Let $(X,f)$ be a t.d.s. with a metric $d$. Then for any $\delta\in(0,1)$,
$\overline{\rm mdim}_{M}(f,X,d)$ $=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(X)}F(\mu,\epsilon),$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}^{e}(X)}F(\mu,\epsilon),$
where $F(\mu,\epsilon)\in\{\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi),\underline{h}^{K}_{\mu}(f,\epsilon,\delta)\}$.
The results are also valid for lower metric mean dimension.
A nice survey involving the variational principles of metric mean dimension in terms of different types of measure-theoretic $\epsilon$-entropies [LW00, GS21, S22] is summarized in [YCZ25, Theorems 1.2 and 1.3].
2.3 Specification property
Sometimes, it is difficult to find a real orbit with desired dynamical behavior, while one can get an approximate orbit of a point that $\epsilon$-shadows any specified finite orbit-segments, which is known as specification property.
Definition 2.4.
A t.d.s. $(X,f)$ has the specification property if for any $\epsilon>0$, there is a positive integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists a point $x\in X$ such that
$d(f^{p}x_{j},f^{p+a_{j}}x)<\epsilon$
for any $p=0,\cdots,b_{j}-a_{j}$, $j=1,\cdots,k$.
In other words, the point $x$ $\epsilon$-shadows the $k$ pieces of finite orbits of $x_{j}$:
$\{x_{j},f(x_{j}),\ldots,f^{b_{j}-a_{j}}(x_{j})\},\ j=1,\ldots,k.$
For instance, every topological mixing locally maximal hyperbolic set, and the full shift on $[0,1]^{\mathbb{Z}}$ have specification property. More examples of infinite entropy systems with specification property can be found in [BR23, Section IV].
A point $x\in X$ is called almost periodic if for any open neighborhood $U$ of $x$, there exists $N\in\mathbb{N}$ such that for every $n\in\mathbb{N}$, there exists an integer $k$ with $n\leq k\leq n+N$ for which $f^{k}(x)\in U$. Let $AP(X)$ denote the set of all almost periodic points of $X$. The following statement, proved in [HTZ23, Proposition 2.11] for systems with $g$-almost product property, gives a characterization of the structure of $AP(X)$.
Proposition 2.2.
If the dynamical system $(X,f)$ has specification property, then the almost periodic set AP(X) is dense in $X$.
3 Proofs of main results
This section proves the main results of this paper. The challenging aspect lies in the estimation of the lower bound of $\overline{\rm mdim}_{M}^{S}(f,G_{K}^{\omega,U},d)$, where $S\in\{B,P,UC\}$. We will construct a Moran fractal set, which is geometrically constructed and is contained in $G_{K}^{\omega,U}$. Consequently, this set is well-suited for computing the lower bound of the metric mean dimensions of saturated sets.
3.1 Packing metric mean dimension of saturated sets
We first prove Theorem 1.1. To arrive at this goal, we divide its proof into two parts.
3.1.1 Upper bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K},d)$
We employ a general method presented in [PS07, Theorem 4.1], [ZCC12, Theorem 1.1, $(2)$] and [JCZ22, Proposition 4] to derive the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K},d)$.
Lemma 3.1.
[Y23, Lemma 3.1] Let $(X,f)$ be a t.d.s. and $\epsilon>0$. Let $\{F_{n}\}$ be a sequence of non-empty subsets of $X$, and let $E_{n}$ be an $(n,\epsilon)$-separated subset of $F_{n}$ with the largest cardinality. We define
$v_{n}:=\frac{1}{\#E_{n}}\sum_{x\in E_{n}}\mathcal{E}_{n}(x).$
Then for any limit point $\mu$ of the sequence $\{\nu_{n}\}$, one has
$\limsup_{n\to\infty}\frac{1}{n}\log\#E_{n}\leq\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
This essentially adheres to the proof of the upper bound of the classical variational principle (cf.[Wal82, Theorem 8.6]). ∎
Recall $G_{K}=\{x\in X:V_{f}(x)=K\}$. For $K\subset M_{f}(X)$, we set
$G^{K}:=\{x\in X:V_{f}(x)\subset K\}.$
Obviously, we have $G_{K}\subset G^{K}$.
Lemma 3.2.
Let $(X,f)$ be a t.d.s..
$(1)$ Let $K\subset M_{f}(X)$ be a nonempty closed convex subset. Then for any $\epsilon>0$,
$h_{top}^{P}(f,G_{K},\epsilon)\leq h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
$(2)$ If $\mu\in M_{f}(X)$, then for any $\epsilon>0$,
$h_{top}^{P}(f,G_{\mu},\epsilon)\leq\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
It suffices to show $(1)$ since one can take $K=\{\mu\}$ as in $(1)$ and use $G_{\mu}=G^{\{\mu\}}$ to get the desired $(2)$.
For $n\in\mathbb{N}$ and $\delta>0$, we put
$\mathcal{R}_{K}(\delta,k):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta),~{}\forall n\geq k\right\},$
$\mathcal{Q}_{K}(\delta,n):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)\right\},$
and
$\Theta(K,\epsilon):=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta,n),\epsilon),$
where $s_{n}(\mathcal{Q}_{K}(\delta,n),\epsilon)$ is the largest cardinality of $(n,\epsilon)$-separated sets of $\mathcal{Q}_{K}(\delta,n)$.
Fix $x\in G^{K}$. Since we have $V_{f}(x)\subset B(K,\delta)$ for every $\delta>0$, then the index $n$ with $\mathcal{E}_{n}(x)\in M_{f}(X)\backslash B(K,\delta)$ must be finite. This shows that $x\in\mathcal{R}_{K}(\delta,k)$ for some $k$. Then, by $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}_{K}(\delta,k)$ for every $\delta>0$, it follows from Proposition 2.1 that
$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{k}h_{top}^{P}(f,\mathcal{R}_{K}(\delta,k),\epsilon)\leq\sup_{k}h_{top}^{UC}(f,\mathcal{R}_{K}(\delta,k),\epsilon).$ (3.1)
For any $\tau>0$, there exists $\delta_{0}>0$ such that
$\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta_{0},n),\epsilon)<\Theta(K,\epsilon)+\tau.$
For any $k\geq 1$, by definitions, it holds that $\mathcal{R}_{K}(\delta_{0},k)\subset\mathcal{Q}_{K}(\delta_{0},n)$ for any $n\geq k$. Thus, we have
$h_{top}^{UC}(f,\mathcal{R}_{K}(\delta_{0},k),\epsilon)\leq\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta_{0},n),\epsilon)<\Theta(K,\epsilon)+\tau.$
Combining (3.1) and letting $\tau\to 0$, we conclude that $h_{top}^{P}(f,G^{K},\epsilon)\leq\Theta(K,\epsilon)$.
Next we shall prove for any $\gamma>0$, there is $\mu\in K$ such that
$\Theta(K,\epsilon)-\gamma\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Take a strictly decreasing sequence $\{\delta_{k}\}$ of positive real numbers with $\delta_{k}\to 0$ and a strictly increasing sequence $\{n_{k}\}$ with $n_{k}\to\infty$ such that
$s_{n_{k}}(\mathcal{Q}_{K}(\delta_{k},n_{k}),\epsilon)>\exp[n_{k}(\Theta(K,\epsilon)-\gamma)].$
Let $F_{k}$ be a maximal $(n_{k},\epsilon)$-separated subset of $\mathcal{R}_{K}(\delta_{k},n_{k})$. We set
$\mu_{k}:=\frac{1}{\#F_{k}}\sum_{x\in F_{k}}\mathcal{E}_{n_{k}}(x).$
Then $\mu_{k}\in B(K,\delta_{k})$ since $K$ is convex. In the weak-topology, choose a subsequence $k_{j}$ such that $\mu_{k_{j}}\to\mu$ as $j\to\infty$. Then $\mu\in K$ since $K$ is a closed. By Lemma 3.1, we have
$\Theta(K,\epsilon)-\gamma\leq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log s_{n_{k}}(\mathcal{Q}_{K}(\delta_{k},n_{k}),\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
This implies that $h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$. ∎
Thus, by the Lemma 3.2 the upper bound of the packing metric mean dimension of $G_{K}$ is given by
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.2)
3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
For any small enough $\epsilon^{\ast}>0$, we prove that
$\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$ (3.3)
Then the above inequality yields the desired lower bound.
Without loss of generality, one may assume $H:=\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)>0$. Let $\zeta>0$ and then take $\mu^{\ast}\in K$ such that
$\inf_{\mathrm{diam}(\xi)<20\epsilon^{\ast}}h_{\mu^{\ast}}(f,\xi)>H-\zeta.$ (3.4)
Since $K$ is a connected and compact set, we choose a sequence $\{\mu_{(k,i)}:k\in\mathbb{N},1\leq i\leq Q_{k}\}\subseteq K$ such that
[…]
$D(\mu_{(k,Q_{k})},\mu_{(k+1,1)})$ $\leq\frac{1}{k},\ \mu_{(k,Q_{k})}=\mu^{\ast}.$
Now endow the set $\mathcal{A}=\{(0,0)\}\cup\{(k,j):k\geq 1,0\leq j\leq Q_{k}\}$ with the lexicographic order. Denote by $\iota(k,j)$ the position of $(k,j)\in\mathcal{A}$. For $(k,j)\in\mathcal{A}$, we shall use $(k,j)^{\ast}$ and $(k,j)_{\ast}$ to represent the one before and the one after $(k,j)$, respectively.
Let $\{\delta_{(k,j)}\}$ be a strictly decreasing sequence indexed by $\mathcal{A}$ with $\delta_{(0,0)}<\zeta$, $\delta_{(k,j)}\to 0$ as $(k,j)\to\infty$. We shall simply write $\delta_{k}=\delta_{(k,Q_{k})}$. For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$, where $m$ is chosen as in the definition of specification property.
For any nonempty open set $U$, there exist $\tilde{\epsilon}>0$ and a point $a_{0}\in X$ such that $\overline{B(a_{0},\tilde{\epsilon})}\subseteq U$. One may assume that $\epsilon^{\ast}<\tilde{\epsilon}$. Set
$\mathcal{C}_{(0,0)}=\{a_{0}\},\ n_{(0,0)}=1,\ N_{(0,0)}=1.$
For any $k\in\mathbb{N}$, there exists $\Delta_{k}=\{a_{1}^{k},\ldots,a_{q_{k}}^{k}\}\subset AP(X)$, which is $\epsilon_{(k,0)}$-dense in $X$ by Proposition 2.2. By the specification property, there exists $a_{k}\in X$ such that
$d\left(f^{(i-1)m_{(k,0)}+i-1}(a_{k}),a_{i}^{k}\right)<\epsilon_{(k,0)},\ i=1,2,\ldots,q_{k}.$
Write $\mathcal{C}_{(k,0)}=\{a_{k}\}$, $n_{(k,0)}=q_{k}+(q_{k}-1)m_{(k,0)}$ and $N_{(k,0)}=1$.
Let $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$ and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$. The following measure approximation of $\mu^{*}$ is a slight modification of [LL24, Lemma 2.1], so we omit the proof.
Lemma 3.3.
For each $k\in\mathbb{N}$, there exists $\nu_{k}\in M_{f}(X)$ satisfying
$(1)$ $v_{k}=\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\nu_{i}^{k}$, where $\{\lambda_{i}^{k}:1\leq i\leq j_{k}\}$ are rational positive numbers, $\forall 1\leq i\leq j_{k}$ with $\sum_{i=1}^{j_{k}}\lambda_{i}^{k}=1$ and $\nu_{i}^{k}\in M_{f}^{e}(X)$;
$(2)$ one has
$\inf_{\xi\succ\mathcal{U}}h_{\mu^{\ast}}(f,\xi)\leq\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)+\zeta;$ (3.5)
$(3)$ $D(\mu^{\ast},\nu_{k})<\delta_{k}$, where $D$ is the compatible metric on $M(X)$ given in (2.1).
Since $\nu_{i}^{k}$ is ergodic, by Birkhoff ergodic theorem there exists $N\in\mathbb{N}$ such that the set
$Y_{i}(N)=\left\{x\in X:D(\mathcal{E}_{n}(x),\nu_{i}^{k})<\delta_{k},\ \forall n>N\right\}$
has $\nu_{i}^{k}$-measure greater than $1-\zeta$ for all $1\leq i\leq j_{k}$. Applying Lemma 2.3, one has
$\underline{h}_{\nu_{i}^{k}}^{K}(f,20\epsilon^{\ast},\zeta)\leq\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)\leq\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)<\infty.$
Now take an integer $n_{k}^{\prime}\in\mathbb{N}$ such that
$(1)$ $\lambda_{i}^{k}n_{k}^{\prime}$ is an integer for all $1\leq i\leq j_{k}$, and
$\frac{n_{k}^{\prime}}{n_{k}}>\max\left\{1-\zeta,1-\frac{1}{k}\right\},$ (3.6)
where $n_{k}=n_{(k,Q_{k})}=n_{k}^{\prime}+(j_{k}-1)m_{(k,Q_{k})}$.
$(2)$ there exists a subset $E_{k,i}$ that is a $(\lambda_{i}^{k}n_{k}^{\prime},5\epsilon^{\ast})$-separated set of $Y_{i}(N)$ satisfying
#$\#E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right],$
where the second inequality is derived from Lemma 2.3.
$(3)$ for every $x\in E_{k,i}$,
$D(\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x),\nu_{i}^{k})<\delta_{k}.$
By the specification property, for $k\geq 1$ and each $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})\in E_{k,1}\times E_{k,2}\times\cdots\times E_{k,j_{k}}$, there exists $y=y(\underline{x}_{k})\in X$ so that
[…]
where $T_{1}=0$ and $T_{i}=\lambda_{1}^{k}n_{k}^{\prime}+\cdots+\lambda_{i-1}^{k}n_{k}^{\prime}+(i-1)m_{k}$ for $2\leq i\leq j_{k}$. One may assume that $d(x,y)<\frac{\epsilon^{\ast}}{2^{k+2}}$ implies $D(\delta_{x},\delta_{y})<\frac{1}{k}$ since the map $x\mapsto\delta_{x}$ is uniformly continuous.
For any two distinct $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})$, $\underline{x}_{k}^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{j_{k}}^{\prime})$ in $E_{k,1}\times\cdots\times E_{k,j_{k}}$, one has $y(\underline{x}_{k})\neq y^{\prime}(\underline{x}_{k}^{\prime})$. Indeed, if $x_{i}\neq x_{i}^{\prime}$ for some $1\leq i\leq j_{k}$, then
$5\epsilon^{\ast}<d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(x_{i},x_{i}^{\prime}\right)\leq d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(x_{i},f^{T_{i}}y\right)+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y,f^{T_{i}}y^{\prime}\right)+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y^{\prime},x_{i}^{\prime}\right)$
$<\frac{\epsilon^{\ast}}{2^{k+1}}+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y,f^{T_{i}}y^{\prime}\right),$
which implies $d_{n_{k}}(y,y^{\prime})>4\epsilon^{\ast}$. Let
$\mathcal{C}_{(k,Q_{k})}=\left\{y(x_{1},x_{2},\ldots,x_{j_{k}}):\ (x_{1},x_{2},\ldots,x_{j_{k}})\in E_{k,1}\times E_{k,2}\times\cdots\times E_{k,j_{k}}\right\}.$
Then $\mathcal{C}_{(k,Q_{k})}$ is an $(n_{k},4\epsilon^{\ast})$-separated set and
#$\#\mathcal{C}_{(k,Q_{k})}$ $=\Pi_{i=1}^{j_{k}}\#E_{k,i}\geq\Pi_{i=1}^{j_{k}}\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right]$
[…]
${\geq}\exp\left[n_{k}(1-\zeta)\left(H-3\zeta\right)\right],$ (3.7)
where the last inequality holds since ${\rm diam}(\xi)<20\epsilon^{\ast}$ and (3.4), (3.6).
For $n\in\mathbb{N}$ and $\varphi\in C(X)$, we write $S_{n}\varphi=\sum_{l=0}^{n-1}\varphi\circ f^{i}$ as the $n$-th Birkhoff sum of $\varphi$. Then for $y\in\mathcal{C}_{(k,Q_{k})}$ and $\varphi\in C(X)$ taking values in $[0,1]$,
$\left|\int\varphi\mathop{}\!\mathrm{d}\mathcal{E}_{n_{k}}(y)-\int\varphi\mathop{}\!\mathrm{d}\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y))\right)\right|$
$\leq$ $\left|\frac{1}{n_{k}}\sum_{i=1}^{j_{k}}S_{\lambda_{i}^{k}n_{k}^{\prime}}\varphi(f^{T_{i}}(y))-\frac{1}{n_{k}^{\prime}}\sum_{i=1}^{j_{k}}S_{\lambda_{i}^{k}n_{k}^{\prime}}\varphi(f^{T_{i}}(y))\right|+\frac{n_{k}-n_{k}^{\prime}}{n_{k}}\|\varphi\|$
$\leq$ $\left[\left|\frac{1}{n_{k}^{\prime}}-\frac{1}{n_{k}}\right|\cdot n_{k}^{\prime}+\frac{n_{k}-n_{k}^{\prime}}{n_{k}}\right]\cdot\|\varphi\|<\frac{2}{k}\|\varphi\|,$
which shows that $D(\mathcal{E}_{n_{k}}(y),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)))<\frac{2}{k}$. Since
$D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i}),\nu_{k}\right)<\delta_{k},\ \ D\left(\nu_{k},\mu^{\ast}\right)<\delta_{k},$
one has that for any $y\in\mathcal{C}_{(k,Q_{k})}$,
$D\left(\mathcal{E}_{n_{k}}(y),\mu^{\ast}\right)\leq\frac{2}{k}+D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i})\right)+2\delta_{k}<\frac{3}{k}+2\delta_{k}.$ (3.8)
Until now, by virtue of the specification property, we have obtained two sets $\mathcal{C}_{(k,0)}$ and $\mathcal{C}_{(k,Q_{k})}$ for each $k$. We proceed to consider the cases for $1\leq j<Q_{k}$.
Since the system has specification property, the generic point $G_{\mu}\neq\emptyset$ holds for any $\mu\in M_{f}(X)$ [PS07]. Applying this fact for any $k\geq 1$ and $1\leq j<Q_{k}$, then there exist $x_{k,j}\in G_{\mu_{(k,j)}}$ and $n_{(k,j)}\in\mathbb{N}$ such that
$D(\mathcal{E}_{n_{(k,j)}}(x_{k,j}),\mu_{(k,j)})<\delta_{(k,j)},$
and
$\frac{n_{(k,j)}}{n_{(k,j)}+m_{(k,j)}}>1-\frac{1}{k}.$
Now set $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$.
With these data $\{\mathcal{C}_{(k,j)}\}$ and $\{n_{(k,j)}\}$, we are ready for the construction:
Step $1$. Construct the intermediate sets $\mathcal{S}_{(k,j)}$.
Recall that we have defined $N_{(k,0)}=1$ for any $k\geq 0$. Now choose a strictly increasing sequence of integers $\{N_{(k,j)}\}_{(k,j)\in\mathcal{A},~{}j\neq 0}$ such that $N_{(1,1)}=1$ and
$\begin{split}\sum_{0\leq k^{\prime}\leq k+1}(n_{(k^{\prime},0)}+m_{(k^{\prime},0)})+n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}&\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})],\\ \sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})]&\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k},\\ n_{k}N_{(k,Q_{k})}&\geq(1-\zeta)\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}].\end{split}$ (3.9)
For $(k,j)\in\mathcal{A}$ with $j=0$, we set $\mathcal{S}_{(k,j)}=\mathcal{C}_{(k,j)}$ and $c_{k,j}=n_{(k,j)}$. We divide the following two cases to define the rest $\mathcal{S}_{(k,j)}$ indexed by $j\neq 0$.
Case $1$. $j\neq Q_{k}$.
Recall that $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$ is defined above. By specification property again, there exists $y_{k,j}\in X$ so that the pieces of orbits $\{x_{k,j},\cdots,f^{n_{(k,j)-1}}(x_{(k,j)})\}$ repeating $N_{(k,j)}$ times, can be $\epsilon_{(k,j)}$-traced by $y_{(k,j)}$ with the gap $m_{(k,j)}$. Now we set $\mathcal{S}_{(k,j)}=\{y_{k,j}\}$, and $c_{k,j}=n_{(k,j)}N_{(k,j)}+(N_{(k,j)}-1)m_{(k,j)}$.
Case $2$. $j=Q_{k}$.
Enumerate the points in $\mathcal{C}_{(k,Q_{k})}$ as
$\mathcal{C}_{(k,Q_{k})}=\{y_{i}^{k}:i=1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}.$
For any $(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}$, let $y(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in X$ be given by the specification property, such that the orbit $\epsilon_{(k,Q_{k})}$-shadows, with the gap $m_{(k,Q_{k})}$, the pieces of orbits $\{y_{i_{l}}^{k},f(y_{i_{l}}^{k}),\ldots,f^{n_{k}-1}(y_{i_{l}}^{k})\}$, $l=1,2,\ldots,N_{(k,Q_{k})}$. We write
$\mathcal{S}_{(k,Q_{k})}=\left\{y=y(i_{1},\ldots,i_{N_{(k,Q_{k})}}):(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}\right\}$
and $c_{k,Q_{k}}=n_{k}\cdot N_{(k,Q_{k})}+(N_{(k,Q_{k})}-1)m_{(k,Q_{k})}$. Since $\mathcal{C}_{(k,Q_{k})}$ is an $(n_{k},4\epsilon^{\ast})$-separated set, this implies that $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set. Then, by (3.1.2) and (3.9), one has the estimation for the cardinality of $\mathcal{S}_{(k,Q_{k})}$:
#$\#\mathcal{S}_{(k,Q_{k})}=\#\mathcal{C}_{(k,Q_{k})}^{N_{(k,Q_{k})}}\geq\exp\left[\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}]\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$ (3.10)
Similar to the arguments in (3.8) one has
$D(\mathcal{E}_{c_{k,j}}(y_{k,j}),\mu_{(k,j)})<\frac{2}{k}+\delta_{(k,j)},\ \forall\ y_{k,j}\in\mathcal{S}_{(k,j)}~{}\text{with}~{}k\geq 1\ \text{and }1\leq j<Q_{k},$
and
$D(\mathcal{E}_{c_{k,Q_{k}}}(y),\mu^{\ast})<\frac{5}{k}+2\delta_{k},\ \forall\ y\in\mathcal{S}_{(k,Q_{k})}.$
Hence, noting that $\mu_{(k,Q_{k})}=\mu^{*}$, in an unified manner we always have
$D(\mathcal{E}_{c_{k,j}}(y),\mu_{(k,j)})<\frac{5}{k}+2\delta_{(k,j)},\ \forall\ y\in\mathcal{S}_{(k,j)}\ \text{with }k\geq 1,\ 1\leq j\leq Q_{k}.$ (3.11)
Step $2$. Construct the $\mathbb{F}_{(k,j)}$ and the Moran set $\mathbb{F}$.
One may regard this process as an analogue of the construction of the middle three Cantor set of the unit interval $[0,1]$.
Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$. Suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$, and then we will describe how to construct $\mathcal{I}_{(k,j)}$. Let
$M_{(k,j)}$ $=M_{(k,j)^{\ast}}+m_{(k,j)}+c_{k,j};$
[…]
Note that $M_{(k,j)}^{N_{(k,j)}}=M_{(k,j)}$ and the number of $\mathcal{S}_{(k,Q_{k})}$ given in formula (3.10) has the estimation:
#$\#\mathcal{S}_{(k,Q_{k})}\geq\exp\left[(M_{(k,Q_{k})}-M_{(k-1,Q_{k-1})})(1-\zeta)^{2}(H-3\zeta)\right].$ (3.12)
For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, by the specification property let $z=z(x,y)\in X$ be some point such that
$d_{M_{(k,j)^{\ast}}}(x,z)<\epsilon_{(k,j)},\ d_{c_{k,j}}(y,f^{M_{(k,j)^{\ast}}+m_{(k,j)}}(z))<\epsilon_{(k,j)}.$
We set $\mathcal{I}_{(k,j)}=\{z(x,y):x\in\mathcal{I}_{(k,j)^{\ast}},y\in\mathcal{S}_{(k,j)}\}$. For $x\in\mathcal{I}_{(k,Q_{k})^{\ast}}$ and two distinct $y,y^{\prime}\in\mathcal{S}_{(k,Q_{k})}$, since $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set, one has
$d_{M_{(k,Q_{k})}}(z(x,y),z(x,y^{\prime}))>2\epsilon^{\ast},$
which leads to
#$\#\mathcal{I}_{(k,j)}=\#\mathcal{I}_{(k,j)^{\ast}}\cdot\#\mathcal{S}_{(k,j)}=\Pi_{(k^{\prime},j^{\prime})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},j^{\prime})}=\Pi_{(k^{\prime},Q_{k^{\prime}})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}.$ (3.13)
Moreover, according to (3.11), for $z\in\mathcal{I}_{(k,j)}$ with $j\neq 0$, we have
$D(\mathcal{E}_{c_{k,j}}(f^{M_{(k,j)^{\ast}}}(z)),\mu_{(k,j)})<\frac{6}{k}+2\delta_{(k,j)}.$ (3.14)
Now set
$\mathbb{F}_{(k,j)}=\bigcup_{x\in\mathcal{I}_{(k,j)}}\overline{B}_{M_{(k,j)}}(x,2\epsilon_{(k,j)}).$
Then for each $(k,j)\in\mathcal{A}$, the following statements are satisfied:
$(1)$ $\overline{B}_{M_{(k,j)}}(x,2\epsilon_{(k,j)})\cap\overline{B}_{M_{(k,j)}}(x^{\prime},2\epsilon_{(k,j)})=\emptyset$ holds for any distinct $x,x^{\prime}\in\mathcal{I}_{(k,j)}$;
$(2)$ if $z\in\mathcal{I}_{(k,j)}$ descends from $x\in\mathcal{I}_{(k,j)^{\ast}}$, one has
$\overline{B}_{M_{(k,j)}}(z,2\epsilon_{(k,j)})\subset\overline{B}_{M_{(k,j)^{\ast}}}(x,2\epsilon_{(k,j)^{\ast}}).$
This implies that $\{\mathbb{F}_{(k,j)}:(k,j)\in\mathcal{A}\}$ is a sequence of decreasing nonempty closed subsets of $X$. Let
$\mathbb{F}=\bigcap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)}=\bigcap_{k\in\mathbb{N}}\mathbb{F}_{(k,Q_{k})}.$
Lemma 3.4.
The fractal set $\mathbb{F}$ is contained in $G_{K}^{\omega,U}$, i.e.,
$\mathbb{F}\subset G_{K}^{\omega,U}=G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\}.$
Proof of Lemma 3.4.
By the construction of $\mathbb{F}$, we know
$\mathbb{F}\subset\mathbb{F}_{(0,0)}=\overline{B}_{n(0,0)}(a_{0},\frac{\epsilon^{\ast}}{8})\subset\overline{B(a_{0},\tilde{\epsilon})}\subset U.$
For $p\in\mathbb{F}$ and any $k\geq 1$, we have $p\in\mathbb{F}_{(k,0)}$. Hence, for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there exists $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<3\epsilon_{(k,0)}$. This implies that the orbit of $p$ is $4\epsilon_{(k,0)}$-dense in $X$. By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have
$\mathbb{F}\subset\{x\in X:\omega_{f}(x)=X\}.$
For $p\in\mathbb{F}$ and sufficiently large $n\in\mathbb{N}$, to estimate $D(\mathcal{E}_{n}(p),K)$ we split into four cases to discuss.
Case $(1)$. There exists $k>1$ such that $M_{(k+1,0)}\leq n<M_{(k+1,1)}^{1}$.
Then there is some $z\in\mathcal{I}_{(k,Q_{k})}$ such that $p\in\overline{B}_{M_{(k,Q_{k})}}(z,\frac{\epsilon_{(k,Q_{k})}}{2})$, and
$D(\mathcal{E}_{n}(p),\mu^{\ast})\leq$ $\frac{M_{(k,Q_{k})^{\ast}}}{n}D(\mathcal{E}_{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})+\frac{c_{k,Q_{k}}}{n}D(\mathcal{E}_{c_{k,Q_{k}}}(f^{M_{(k,Q_{k})^{\ast}}}(p)),\mu^{\ast})$
$+\frac{n-M_{(k,Q_{k})}}{n}D(\mathcal{E}_{n-M_{(k,Q_{k})}}(f^{M_{(k,Q_{k})}}(p)),\mu^{\ast}).$
Noticing that both $\frac{M_{(k,Q_{k})^{\ast}}}{n}$ and $\frac{n-M_{(k,Q_{k})}}{n}$ are smaller than $\frac{1}{k}$, and using formula (3.14), we have $D(\mathcal{E}_{c_{k,Q_{k}}}f^{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})<\frac{7}{k}+2\delta_{(k,Q_{k})}$. Thus,
$D(\mathcal{E}_{n}(p),K)\leq\frac{9}{k}+2\delta_{(k,Q_{k})}.$
Case $(2)$. There exists $(k,j)\in\mathcal{A}$ with $j\neq 0$ such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
Then there is some $z\in\mathcal{I}_{(k,j)}$ such that $p\in\overline{B}_{M_{(k,j)}}(z,\frac{\epsilon_{(k,j)}}{2})$, and
$D(\mathcal{E}_{n}(p),\mu_{(k,j)})\leq$ $\frac{M_{(k,j)^{\ast}}}{n}D(\mathcal{E}_{M_{(k,j)^{\ast}}}(p),\mu_{(k,j)})+\frac{c_{k,j}}{n}D(\mathcal{E}_{c_{k,j}}(f^{M_{(k,j)^{\ast}}}(p)),\mu_{(k,j)})$
$+\frac{n-M_{(k,j)}}{n}D(\mathcal{E}_{n-M_{(k,j)}}(f^{M_{(k,j)}}(p)),\mu_{(k,j)}).$
The similar argument as in Case $(1)$ yields that
$D(\mathcal{E}_{n}(p),K)\leq\frac{9}{k}+2\delta_{(k,j)}.$
Case $(3)$. There exist $(k,1)\in\mathcal{A}$ and $i>1$ such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
One has
$D($ $\mathcal{E}_{n}(p),\mu_{(k,1)})\leq\frac{M_{(k-1,Q_{k-1})^{\ast}}}{n}D(\mathcal{E}_{M_{(k-1,Q_{k-1})^{\ast}}}(p),\mu_{(k,1)})$
$+\frac{c_{k-1,Q_{k-1}}}{n}D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})+\frac{c_{k,0}}{n}D(\mathcal{E}_{c_{k,0}}(f^{M_{(k-1,Q_{k-1})}}(p)),\mu_{(k,1)})$
$+\frac{n_{(k,1)}+m_{(k,1)}}{n}\sum_{l=0}^{i-1}D(\mathcal{E}_{n_{(k,1)}+m_{(k,1)}}(f^{M_{(k,1)}^{l}}(p)),\mu_{(k,1)})$
$+\frac{n-M_{(k,1)}^{i}}{n}D(\mathcal{E}_{n-M_{(k,1)}^{i}}(f^{M_{(k,1)}^{i}}(p)),\mu_{(k,1)}).$
Note that
$\frac{M_{(k-1,Q_{k-1})^{\ast}}}{n}<\frac{1}{k-1},\ \frac{c_{k,0}}{n}<\frac{1}{k},\ \frac{n-M_{(k,1)}^{i}}{n}<\frac{1}{k},$
$D(\mathcal{E}_{n_{(k,1)}+m_{(k,1)}}(f^{M_{(k,1)}^{l}}(p)),\mu_{(k,1)})\leq\frac{5}{k}+2\delta_{(k,1)},$
and
$D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})$ $\leq D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu^{\ast})+D(\mu^{\ast},\mu_{(k,1)})$
$\leq\frac{8}{k-1}+2\delta_{(k-1,Q_{k-1})}.$
This gives us
$D(\mathcal{E}_{n}(p),K)\leq\frac{16}{k-1}+4\delta_{(k-1,0)}.$
Case $(4)$. There exist $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$ such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
In this case, using the expression
$\mathcal{E}_{n}(z)=$ $\frac{M_{(k,j-1)^{\ast}}}{n}\mathcal{E}_{M_{(k,j-1)^{\ast}}}(p)+\frac{c_{k,j-1}}{n}\mathcal{E}_{c_{k,j-1}}(f^{M_{(k,j-1)^{\ast}}}(p))$
$+$ $\frac{n_{(k,j)}+m_{(k,j)}}{n}\sum_{l=0}^{i-1}\mathcal{E}_{n_{(k,j)}+m_{(k,j)}}(f^{M_{(k,j)}^{l}}(p))+\frac{n-M_{(k,1)}^{i}}{n}\mathcal{E}_{n-M_{(k,1)}^{i}}(f^{M_{(k,1)}^{i}}(p)),$
and repeating the similar argument as above, one can obtain
$D(\mathcal{E}_{n}(p),K)\leq\frac{15}{k}+4\delta_{(k,j)}.$
To sum up, whatever case we consider, by letting $n\to\infty$ we get $V_{f}(p)\subset K$. Based on the above arguments, one can also show that the sequences $\{\mathcal{E}_{n}(p)\}_{n\in\mathbb{N}}$ and $\{\mu_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ have the same limit points. Notice that $K=\overline{\{\mu_{(k^{\prime},j^{\prime})}:(k^{\prime},j^{\prime})>(k,j)\}}$ holds for any fixed $(k,j)\in\mathcal{A}$. Consequently, we have $V_{f}(p)=K$, and hence $\mathbb{F}\subset G_{K}$. ∎
Step $3$. Construct the measure $\rho$ supported on $\mathbb{F}$.
Now we define a probability measure $\rho$ on $\mathbb{F}$ satisfying the $\rho$-measure of some appropriate dynamical balls decays exponentially fast along the sequence $\{M_{k}\}$, where $M_{k}:=M_{(k,Q_{k})}$. Then we can invoke the Entropy Distribution Principle developed in Lemma 2.2 to estimate the lower bound of $\epsilon$-packing topological entropy of $\mathbb{F}$.
Define
$\rho_{k}=\frac{1}{\#\mathcal{I}_{(k,Q_{k})}}\sum_{x\in\mathcal{I}_{(k,Q_{k})}}\delta_{x}.$
For simplicity, we write $\mathcal{I}_{k}$, $\mathbb{F}_{k}$ instead of $\mathcal{I}_{(k,Q_{k})}$, $\mathbb{F}_{(k,Q_{k})}$ respectively. Since $M(X)$ is compact, we may assume that $\rho_{k}\to\rho\in M(X)$ as $k\to\infty$. It is clear that $\rho_{k}(\mathbb{F}_{k})=1$ for every $k$, and for any $k>k_{0}$, one has $\rho_{k}(\mathbb{F}_{k_{0}})=1$ since $\mathbb{F}_{k}\subset\mathbb{F}_{k_{0}}$. Thus one has $\rho(\mathbb{F}_{k_{0}})\geq\limsup_{k\to\infty}\rho_{k}(\mathbb{F}_{k_{0}})=1$. This implies that $\rho(\mathbb{F})=1$. Recall that $M_{(0,0)}=n_{(0,0)}=1$, by (3.12) and (3.13), we have
#$\#\mathcal{I}_{k}\geq\exp\left[(M_{k}-1)\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
Lemma 3.5.
Let $k_{0}\geq 1$, $z\in X$ and $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ be any Bowen ball that intersects $\mathbb{F}$. Then for any $k>k_{0}$, one has
$\rho_{k}\left(B_{M_{k_{0}}}\left(z,\frac{\epsilon^{\ast}}{2}\right)\right)\leq\frac{1}{\#\mathcal{I}_{k_{0}}}.$
Proof of Lemma 3.5.
Since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}\not=\emptyset$, the set $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}_{k_{0}}$ is also nonempty by the construction of $\mathbb{F}$ and #$\#(\mathbb{F}_{k_{0}}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))=1$. Notice that $\mathbb{F}_{k_{0}+1}$ descends from $\mathbb{F}_{k_{0}}$. Then one has
#$\#(\mathbb{F}_{k_{0}+1}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))\leq\#\mathcal{S}_{(k_{0}+1,Q_{k_{0}+1})},$
and hence
$\rho_{k_{0}+1}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{\#\mathcal{S}_{(k_{0}+1,Q_{k_{0}+1})}}{\#\mathcal{I}_{k_{0}+1}}=\frac{1}{\#\mathcal{I}_{k_{0}}}.$
In general, for any $k>k_{0}$,
$\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{\Pi_{k_{0}+1\leq k^{\prime}\leq k}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}}{\#\mathcal{I}_{k}}=\frac{1}{\#\mathcal{I}_{k_{0}}}.$
For sufficiently large $k_{0}$ and any Bowen ball $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ with $z\in\mathbb{F}$, since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ is open, one has
$\rho\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)$ $\leq\liminf_{k\to\infty}\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{1}{\#\mathcal{I}_{k_{0}}}$
$\leq\exp\left[-(M_{k_{0}}-1)(1-\zeta)^{2}\cdot(H-3\zeta)\right]$
$\leq\exp\left[-M_{k_{0}}(1-\zeta)^{3}\cdot(H-3\zeta)\right]$
Using Lemma 2.2, one has
$(1-\zeta)^{3}\cdot(H-3\zeta)\leq h_{top}^{P}(f,\mathbb{F},\frac{\epsilon^{\ast}}{12})\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$
The arbitrariness of $\zeta$ implies that
$\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$
Dividing in both sides of the above inequality by $|\log\epsilon^{\ast}|$, and then taking $\limsup_{\epsilon^{\ast}\to 0}$, we have
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d).$ (3.15)
Finally, by (3.2) and (3.15) we get
$\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
which proves Theorem 1.1.
3.2 Bowen metric mean dimension of saturated sets
In this subsection, we prove Theorems 1.2 and 1.3.
3.2.1 Upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$
The following lemma gives an upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$, which was already obtained in [Y23, Lemma 3.3,(3)].
Lemma 3.6.
Let $(X,f)$ be a t.d.s., and let $K\subset M_{f}(X)$ be a nonempty closed subset.
  • (1)
    For any $\epsilon>0$, we have
    $h_{top}^{B}(f,\{x\in X:V_{f}(x)\cap K\not=\emptyset\},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
  • (2)
    For any $\epsilon>0$, we have
    $h_{top}^{B}(f,G_{K},\epsilon)\leq\inf_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Thus, by the Lemma 3.6 the upper bound of the Bowen metric mean dimension of saturated set is given by:
$\overline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.16)
3.2.2 Lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
The proof for the lower bound is similar to that in subsection 3.1.2. So we only give a sketch for the estimation of the lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$.
We show that for any small enough $\epsilon^{\ast}>0$, it holds that
$\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2}).$ (3.17)
Fix $\zeta>0$ and pick a sequence $\{\mu_{(k,i)}:k\geq 1,1\leq i\leq Q_{k}\}\subset K$ of measures in $K$ such that
$K\subset\cup_{i=1}^{Q_{k}}B(\mu_{(k,i)},\frac{1}{k}),\ D(\mu_{(k,i)},\mu_{(k,i+1)})<\frac{1}{k},\ D(\mu_{(k,Q_{k})},\mu_{(k+1,1)})<\frac{1}{k}.$
Put $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$, $\mathcal{A}=\{(k,0):k\geq 0\}\cup\mathcal{A}^{\prime}$, and endow these sets with the lexicographic order again. Using the fixed notation in subsection 3.1.2, we have the data $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ and $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$.
Assume that $H:=\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)>0$. Then one has
$\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu_{(k,i)}}(f,\xi)>H-\zeta,\ \forall(k,i)\in\mathcal{A}^{\prime}.$
Let $\mathcal{U}$ be an open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$, and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$. For $(k,i)\in\mathcal{A}^{\prime}$, by Lemma 3.3 let $\nu_{(k,i)}=\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\nu_{l}^{k,i}$ be the finite convex combination of ergodic $f$-invariant probability measures with rational coefficients such that
$D(\mu_{(k,i)},\nu_{(k,i)})<\delta_{(k,i)},\ \ \inf_{\xi\succ\mathcal{U}}h_{\mu_{(k,i)}}(f,\xi)\leq\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)+\zeta.$
Applying Katok’s $\epsilon$-entropy formula and repeating the previous argument as in subsection 3.1.2, we choose an integer $n_{(k,i)}^{\prime}$ such that the following conditions are satisfied:
$(1)$ for any $1\leq l\leq p_{k,i}$, $\lambda_{k,i,l}n_{(k,i)}^{\prime}$ is an integer and
$\frac{n_{(k,i)}^{\prime}}{n_{(k,i)}+m_{(k,i)}}>\max\{1-\zeta,1-\frac{1}{k}\},$
where $n_{(k,i)}:=n_{(k,i)}^{\prime}+(p_{k,i}-1)m_{(k,i)}$;
$(2)$ there exist subsets $E_{k,i,l}$, $1\leq l\leq p_{k,i}$, which are $(\lambda_{k,i,l}n_{(k,i)}^{\prime},5\epsilon^{\ast})$-separated and
#$\#E_{k,i,l}\geq\exp\left[\lambda_{k,i,l}n_{(k,i)}^{\prime}\cdot\left(\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)-\zeta\right)\right];$
$(3)$ for any $x\in E_{k,i,l}$,
$D(\mathcal{E}_{\lambda_{k,i,l}n_{(k,i)}^{\prime}}(x),\nu_{l}^{k,i})<\delta_{(k,i)}.$
Let
$\mathcal{C}_{(k,i)}=\left\{y=y(x_{1},\ldots,x_{p_{k,i}}):(x_{1},\ldots,x_{p_{k,i}})\in E_{k,i,1}\times\cdots\times E_{k,i,p_{k,i}}\right\},$
where $y(x_{1},\ldots,x_{p_{k,i}})$ is the point that $\epsilon_{(k,i)}$-shadows the pieces of orbits
[…]
with gap $m_{(k,i)}$. Then $\mathcal{C}_{(k,i)}$ is a $(n_{(k,i)},4\epsilon^{\ast})$-separated set with
#$\#\mathcal{C}_{(k,i)}\geq\exp\left[n_{(k,i)}(1-\zeta)\cdot(H-2\zeta)\right].$
For $k\geq 0$, let $\mathcal{C}_{(k,0)}$, $n_{(k,0)}$, $N_{(k,0)}$ be defined as in subsection 3.1.2. Choose a strictly increasing sequence $\{N_{(k,i)}\}_{(k,i)\in\mathcal{A}^{\prime}}$ such that $N_{(1,1)}=1$ and
$\begin{split}\sum_{0\leq k^{\prime}\leq k+1}(n_{(k^{\prime},0)}+m_{(k^{\prime},0)})+n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}&\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})],\\ \sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})]&\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k}.\end{split}$
Set $\mathcal{S}_{(k,0)}=\mathcal{C}_{(k,0)}$, $c_{k,0}=n_{(k,0)}$ for $k\geq 0$. For $(k,i)\in\mathcal{A}^{\prime}$, and any tuple of points $(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})\in\mathcal{C}_{(k,i)}^{N_{(k,i)}}$, there exists a point $z=z(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})$ that $\epsilon_{(k,i)}$-shadows the pieces of orbits
$\{y_{k,i,l},f(y_{k,i,l}),\ldots,f^{n_{(k,i)}-1}(y_{k,i,l})\},\ l=1,\cdots,N_{(k,i)},$
with the same gap $m_{(k,i)}$. Denote by $\mathcal{S}_{(k,i)}$ the set of such shadowing points. It is a $(c_{k,i},3\epsilon^{\ast})$-separated set, where $c_{k,i}=n_{(k,i)}N_{(k,i)}+(N_{(k,i)}-1)m_{(k,i)}$.
Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$. Suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$. Let $M_{(k,j)}=M_{(k,j)^{\ast}}+m_{(k,j)}+c_{k,j}$. For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, let $z=z(x,y)$ be the shadowing point such that $d_{M_{(k,j)^{\ast}}}(z,x)<\epsilon_{(k,j)}$ and
$d_{c_{k,j}}(y,f^{M_{(k,j)^{\ast}}+m_{(k,j)}}(z))<\epsilon_{(k,j)}.$
We set $\mathcal{I}_{(k,j)}=\{z=z(x,y):x\in\mathcal{I}_{(k,j)^{\ast}},y\in\mathcal{S}_{(k,j)}\}$, and then define
$\mathbb{F}=\cap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)},$
where $\mathbb{F}_{(k,j)}=\cup_{z\in\mathcal{I}_{(k,j)}}\overline{B}_{M_{(k,j)}}(z,2\epsilon_{(k,j)})$. One can verify that $\mathbb{F}\subset G_{K}^{\omega,U}$.
We proceed to define a sequence of measures $\{\rho_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ by
$\rho_{(k,j)}:=\frac{1}{\#\mathcal{I}_{(k,j)}}\sum_{x\in\mathcal{I}_{(k,j)}}\delta_{x}.$
Without loss of generality, we can assume that $\rho_{(k,j)}\to\rho$. Furthermore, one can check that $\rho(\mathbb{F})=1$, and for every $n$ large enough and any ball $B_{n}(z,\frac{\epsilon^{\ast}}{2})$ intersecting with the fractal set $\mathbb{F}$, it holds that
$\rho(B_{n}(z,\frac{\epsilon^{\ast}}{2}))\leq\exp\left[-n(1-\zeta)^{2}(H-2\zeta)\right].$
Then, by the Lemma 2.1, we have
$(1-\zeta)^{2}(H-2\zeta)\leq h_{top}^{B}(f,\mathbb{F},\frac{\epsilon^{\ast}}{2})\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2}).$
Letting $\zeta\to 0$ gives us
$\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2})$
Therefore, by (3.16) and (3.17) we get
$\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{B}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{\mathrm{diam}\xi<\epsilon}h_{\mu}(f,\xi).$
3.2.3 Proof of Theorem 1.3
With the help of Theorems 1.1 and 1.2, we prove Theorem 1.3.
Firstly, we show
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\}.$ (3.18)
Fix $\mu\in K$ and $(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)$. Then, by Lemma 3.6, for every $\epsilon>0$,
$h_{top}^{B}(f,G_{K},\epsilon)\leq\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
This shows
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)$
and hence $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\underline{\rm mrid}_{\mu}(f,K,d)$ for every $\mu\in K$.
On the other hand, we may assume that $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)$ is finite; otherwise there is nothing left to prove. Let $\gamma>0$. Then, by Theorem 1.2, we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for any $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi)<\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
Assume that $\mu_{k}\rightarrow\mu$ as $k\to\infty$. Then $\mu\in K$ since $K$ is closed. This shows that
$\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
Letting $\gamma\to 0$, we have
$\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\}\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d).$
This completes the proof of (3.18).
Next we show the second equality:
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
Using formula (3.3), the inequality $\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}\leq\overline{\rm mdim}_{M}^{P}(f,G_{K},d)$ is clear. On the other hand, let $\gamma>0$ be sufficiently small and put
$M_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{P}(f,G_{K},d)-\gamma\}.$
Then we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for sufficiently large $k$,
$M_{\gamma}<\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi).$
Again, we assume that $\mu_{k}\rightarrow\mu\in K$. This shows that
$M_{\gamma}\leq\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}.$
Letting $\gamma\to 0$, we have
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)\leq\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
3.3 Upper capacity metric mean dimension of saturated sets
In this subsection, we prove the Theorem 1.4.
By Proposition 2.1, the inequality
$\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d)\leq\overline{\rm mdim}_{M}^{UC}(f,G_{K},d)\leq\overline{\rm mdim}_{M}(f,X,d)$
is clear. Thus, it suffices to show
$\overline{\rm mdim}_{M}(f,X,d)\leq\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d).$
In light of Theorem 2.4, we need to show that for $\epsilon^{\ast}>0$, $\zeta>0$ sufficiently small and $\delta\in(0,1)$, one has
$\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta\leq h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast}).$ (3.19)
However, we will give a sketch for the proof of (3.19) since it is essentially analogous to that of the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$ presented in subsection 3.1.2.
Write $H=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)$ and assume that $H>0$. Take $\mu^{\ast}\in M_{f}^{e}(X)$ such that
$\underline{h}_{\mu^{\ast}}^{K}(f,3\epsilon^{\ast},\delta)>H-\zeta.$
Then, for sufficiently large $\mathcal{N}\in\mathbb{N}$, there exists a $(\mathcal{N},3\epsilon^{\ast})$-separated subset $S_{\mathcal{N}}$ with
#$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)].$
Now fix such $\mathcal{N}$, and write $\mathcal{C}_{(0,0)}=S_{\mathcal{N}}$, $c_{(0,0)}=\mathcal{N}$, $N_{(0,0)}=1$.
Since $K$ is a connected closed subset, a standard argument shows that there exists a family $\{\mu_{k,i}:k\geq 1,1\leq i\leq Q_{k}\}$ of measures in $K$ such that for all $k\geq 1$, $1\leq i<Q_{k}$,
$K\subset\bigcup_{i=1}^{Q_{k}}B(\mu_{k,i},\frac{1}{k}),\ D(\mu_{k,i},\mu_{k,i+1})<\frac{1}{k},\ \text{and }D(\mu_{k,Q_{k}},\mu_{k+1,1})<\frac{1}{k}.$
Consider the set $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$ and endow $\mathcal{A}=\{(k,0):k\in\mathbb{Z}\ \text{and }k\geq-1\}\cup\mathcal{A}^{\prime}$ with the lexicographic order. Similarly, the sequences $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ are constructed as in subsection 3.1.2.
For the nonempty open set $U$, there exist $a_{0}\in X$ and $\epsilon_{0}\in(0,\epsilon^{\ast})$ with $\overline{B(a_{0},\epsilon_{0})}\subset U$. Set $\mathcal{C}_{(-1,0)}=\{a_{0}\}$, $n_{(-1,0)}=c_{(-1,0)}=N_{(-1,0)}=1$. For $k\geq 1$, let $\mathcal{C}_{(k,0)}$, $n_{(k,0)}$, $c_{(k,0)}$ and $N_{(k,0)}$ be the same as in subsection 3.1.2. For $(k,i)\in\mathcal{A}^{\prime}$, since $G_{\mu_{k,i}}\neq\emptyset$, one can pick $x_{k,i}\in X$ and $n_{(k,i)}\in\mathbb{N}$ such that
$D(\mathcal{E}_{n_{(k,i)}}(x),\mu_{k,i})<\delta_{(k,i)},\ \frac{n_{(k,i)}}{n_{(k,i)}+m_{(k,i)}}>\max\left\{1-\zeta,1-\frac{1}{k}\right\}.$
Set $\mathcal{C}_{(k,i)}=\{x_{k,i}\}$. Now choose a strictly increasing sequence of integers $\{N_{(i,j)}\}_{(k,j)\in\mathcal{A}\ \text{with }j\neq 0}$ with $N_{(1,1)}=1$ and
$\begin{split}&n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}\left[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})\right],\\ &\sum_{(k^{\prime},j^{\prime})\leq(k,j)}\left[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})\right]\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k}.\end{split}$
Based on the data $\mathcal{C}_{(k,j)}$, $N_{(k,j)}$ for $(k,j)\in\mathcal{A}$, similar to the previous subsections, one can construct the intermediate sets $\mathcal{S}_{(k,j)}$, $\mathcal{I}_{(k,j)}$, $\mathbb{F}_{(k,j)}$ to obtain the fractal set
$\mathbb{F}=\bigcap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)}\subset G_{K}^{\omega,U}.$
It is worth mentioning that for any $(k,j)>(0,0)$, #$\#\mathcal{I}_{(k,j)}=\#S_{\mathcal{N}}$. Moreover, for any distinct $x$, $x^{\prime}\in\mathcal{I}_{(0,0)}$, one has
$d_{M_{(0,0)}}(x,x^{\prime})>2\epsilon^{\ast}.$
Recall $M_{(0,0)}=1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}$. Let $y$, $y^{\prime}\in\mathbb{F}$ with $y\in\overline{B}_{M_{(0,0)}}(x,2\epsilon_{(0,0)})$ and $y^{\prime}\in\overline{B}_{M_{(0,0)}}(x^{\prime},2\epsilon_{(0,0)})$. Then one has
$d_{M_{(0,0)}}(y,y^{\prime})>d_{M_{(0,0)}}(x,x^{\prime})-d_{M_{(0,0)}}(x,y)-d_{M_{(0,0)}}(x^{\prime},y^{\prime})>\epsilon^{\ast}.$
Hence, for sufficiently large $\mathcal{N}$, there is a $(1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N},\epsilon^{\ast})$-separated subset in $\mathbb{F}$ with cardinality #$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)]$. Thus
$h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast})\geq\limsup_{\mathcal{N}\to\infty}\frac{\mathcal{N}(H-2\zeta)}{1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}}=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta.$
This implies the inequality (3.19).
3.4 An extension to metric mean dimensions of sup sets and sub sets
In this section, we consider two variants of saturated sets, namely sup sets and sub sets, and establish variational principles for metric mean dimensions of sup sets and sub sets. The main results are presented in Theorems 3.7, 3.8 and 3.9.
In general, given a point $x\in X$, the limit $\lim_{n\to\infty}\mathcal{E}_{n}(x)$ may not exist in the sense of weak-topology. The set of such points is denoted by $I(f)$, known as the historic set [^1] This terminology is due to Rulle’s original paper [Rue01]. of $f$. By Birkhoff ergodic theorem, although the historic set is usually regarded as negligible in the measure-theoretic context, its topological dynamics can be quite rich. Especially for systems with specification-like properties, it can even have full topological entropy [CTS05] and full metric mean dimension [LV21, LL24]. To understand the underlying geometry structure of $I(f)$, we define the following sets based on the framework introduced and further developed by Olsen [O02, O03, O06, O08] and Olsen and Winter [OW03]. More precisely, given a nonempty subset $C\subset M_{f}(X)$, we define the sup set, equ set [^2] In this context, we do not take into account the deformation of $\mathcal{E}_{n}(x)$ via a continuous affine map from $M_{f}(X)$ to a topological vector space $Y$. Hence, we prefer to use the term “saturated set” rather than “equ set”. and sub set of $C$ as
${}^{C}G:$ =$=\{x\in X:V_{f}(x)\supset C\},$
$G_{C}:$ =$=\{x\in X:V_{f}(x)=C\},$
$G^{C}:$ =$=\{x\in X:V_{f}(x)\subset C\},$
respectively.
Note that the equ set is precisely the saturated set associated with $C$ defined previously. More generally, by using two nonempty subsets $C_{1},C_{2}\subset M_{f}(X)$, we define the set
$G(C_{1},C_{2}):=\{x\in X:C_{1}\subset V_{f}(x)\subset C_{2}\},$
which covers the three different notions above in an unified manner.
Now, we are ready to establish variational principles for the three given types of level sets $G^{C},^{C}G,G(C_{1},C_{2})$ by means of the Bowen metric mean dimension.
Theorem 3.7.
Let $(X,d,f)$ be a t.d.s. with specification property. Then for any nonempty subset $C\subseteq M_{f}(X)$,
$\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
We claim that for every $\epsilon>0$,
$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)=\inf_{\mu\in\overline{{\rm Conv}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$ (3.20)
where $\overline{{\rm Conv}}(C)$ denotes the convex closure of set $C$.
Indeed, since $C\subset\overline{{\rm Conv}}(C)$, it suffices to show
$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\inf_{\mu\in\overline{{\rm Conv}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Let $\mu\in\overline{{\rm Conv}}(C)$. We can choose $\mu_{n}=\sum_{j=1}^{k_{n}}\lambda_{n}^{(j)}v_{n}^{(j)}$ with $v_{n}^{(j)}\in C$, $\lambda_{n}^{(j)}>0$ and $\sum_{j=o}^{k_{n}}\lambda_{n}^{(j)}=1$ such that $\mu_{n}\rightarrow\mu$ as $n\to\infty$. The upper semi-continuity of $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ on $M_{f}(X)$ (cf. [FW16, Lemma 2.3, (3)]) implies that
$\limsup_{n\to\infty}~{}{\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu_{n}}(f,\xi)}\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Combining this and $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ is concave in $\mu$, we can deduce that
$\inf_{\nu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\nu}(f,\xi)<\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)+\gamma$
holds for any for $\gamma>0$. This yields the desired inequality and hence (3.20) holds.
Notice that $G_{\overline{{\rm Conv}}(C)}\subset^{C}G$, we have
$\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm Conv}}(C)},d\right)$
[…]
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Conversely, for any $\mu\in C$ and $\epsilon>0$, by Lemma 3.6 one has
$h_{top}^{B}(f,^{C}G,\epsilon)\leq h_{top}^{B}(f,\{x\in X:\mu\in V_{f}(x)\},\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
This shows $h_{top}^{B}(f,^{C}G,\epsilon)\leq\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ for every $\epsilon>0$. ∎
Theorem 3.8.
Let $(X,d,f)$ be a t.d.s. with specification property. Then for any nonempty closed subset $C\subseteq M_{f}(X)$, we have
$\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
For every $\epsilon>0$, by Lemma 3.6 again we have
$h_{top}^{B}(f,G^{C},\epsilon)\leq h_{top}^{B}(f,\{x\in X:V_{f}(x)\cap{C}\not=\emptyset\},\epsilon)\leq\sup_{\mu\in{C}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
By (3.17), there exists a constant $c>0$ such that
$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu},\epsilon)$
holds for all $\mu\in C$ and any sufficiently small $\epsilon>0$. Combining with the fact that $G_{\mu}\subset G^{C}$ holds for each $\mu\in C$, one gets
$\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G^{C},\epsilon).$
Theorem 3.9.
Let $(X,d,f)$ be a t.d.s. with specification property. Suppose that $C_{1}\subseteq M_{f}(X)$ is a nonempty set satisfying that $\overline{{\rm Conv}}(C_{1})$ is a connected component of $C_{2}\subset M_{f}(X)$. Then
$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C_{1}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
Since $C_{1}\subset\overline{{\rm Conv}}(C_{1})\subset C_{2}$, by Theorem 1.2 we have
$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm Conv}}(C_{1})},d\right)$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in\overline{{\rm Conv}}(C_{1})}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C_{1}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Furthermore, we have $\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)\leq\overline{\rm mdim}_{M}^{B}\left(f,^{C_{1}}G,d\right)$. Using Theorem 3.7, we get
$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C_{1}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
This completes the proof. ∎
The Theorems 3.7, 3.8 and 3.9 imply the following corollaries, which allows us to compare the metric mean dimensions of sup sets and sub sets with the saturated sets.
Corollary 3.10.
Let $(X,d,f)$ be a t.d.s. with specification property.
  • (1)
    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)=\overline{\rm mdim}_{M}^{B}\left(f,G_{C},d\right).$
  • (2)
    If $C$ is a nonempty compact convex set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right).$
Proof.
(1) follows by Theorems 1.2 and 3.7; (2) is due to Theorems 1.1 and 3.8. ∎
Corollary 3.11.
Let $(X,d,f)$ be a t.d.s. with specification property.
  • (1)
    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{B}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)\leq\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right).$
  • (2)
    If $C$ is a nonempty compact convex set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right)\leq\overline{\rm mdim}_{M}^{P}\left(f,^{C}G,d\right).$
  • (3)
    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{UC}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{UC}\left(f,G^{C},d\right)=\overline{\rm mdim}_{M}^{UC}\left(f,^{C}G,d\right).$
Proof.
$(1)$ is clear. For $(2)$, it follows from Theorem 1.1 that
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)=\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)\leq\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right).$
On the other hand, one has $h_{top}^{P}(f,G^{C},\epsilon)\leq\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ by Lemma 3.2. Hence, we have $\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right)$.
$(3)$ follows from Theorem 1.4 and the relations $G_{C}\subset G^{C}$ and $G_{C}\subset^{C}G$. ∎
4 Some applications of main results
In this section, we apply the main results to multifractal analysis of level sets and mean Li-Yorke chaos in infinite entropy systems. The main results in this section are presented in Theorems 4.1 and 4.3.
4.1 Variational principle of level sets
In this subsection, we establish variational principles for level sets of continuous functions on $M_{f}(X)$ within the context of an abstract framework.
Given a dynamical system $(X,f)$, assume that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function. We define
$L_{\varphi}=\left[\inf_{\mu\in M_{f}(X)}\varphi(\mu),\ \sup_{\mu\in M_{f}(X)}\varphi(\mu)\right].$
For any $\alpha\in L_{\varphi}$, we define $M_{f}(\varphi,\alpha):=\left\{\mu\in M_{f}(X):\varphi(\mu)=\alpha\right\}$. The continuity of $\varphi$ then implies that $M_{f}(\varphi,\alpha)$ is a nonempty closed subset of $M_{f}(X)$. We define the level set of $\alpha$ w.r.t. $\varphi$ as
$K_{\alpha}:=\left\{x\in X:\inf_{\mu\in V_{f}(x)}\varphi(\mu)=\sup_{\mu\in V_{f}(x)}\varphi(\mu)=\alpha\right\},$
which consists of the set of points of $X$ such that $\varphi$ takes constant value $\alpha$ on $V_{f}(x)$.
Condition A: For every $\alpha\in L_{\varphi}$, we always assume that $M_{f}(\varphi,\alpha)$ is a convex subset.
Several examples of the function $\varphi$ that satisfy Condition A are stated below:
  • (a)
    If $\varphi=\alpha\in\mathbb{R}$ is a constant, then $L_{\varphi}=\{\alpha\}$, $K_{\alpha}=X$, $M_{f}(\varphi,\alpha)=M_{f}(X)$;
  • (b)
    We define
    $\varphi(\mu):=\frac{\int\psi\mathop{}\!\mathrm{d}\mu}{\int\phi\mathop{}\!\mathrm{d}\mu},$
    where $\psi$, $\phi\in C(X)$ and $\int\phi\mathop{}\!\mathrm{d}\mu\neq 0$ for all $\mu\in M_{f}(X)$ (e.g. $\phi>0$).
    Then for any $\alpha\in L_{\varphi}$,
    $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\int\phi\mathop{}\!\mathrm{d}\mu\},$
    $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\int\phi\mathop{}\!\mathrm{d}\mu\};$
  • (c)
    Let $\Phi:=\{\phi_{n}\}$ be an asymptotically additive potential, that is, for any $\delta>0$, there exists $\psi_{\delta}\in C(X)$ such that $\limsup_{n\to\infty}\frac{1}{n}||\phi_{n}-S_{n}\psi_{\delta}||<\delta$. We define
    $\varphi:M_{f}(X)\to\mathbb{R},~{}\mu\mapsto\lim_{n\to\infty}\int\frac{\phi_{n}(x)}{n}\mathop{}\!\mathrm{d}\mu(x).$
    Then $\varphi$ is continuous (cf. [FH10, Lemma A.4]), and $M_{f}(\varphi,\alpha)$ is a convex subset for all $\alpha\in L_{\Phi}$. Specially, let $\psi\in C(X)$ and define $\phi_{n}=S_{n}\psi$, then $\Phi=\{\phi_{n}\}$ is an additive potential, in this case we have
    $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\ \int\psi\mathop{}\!\mathrm{d}\mu=\alpha\},$
    $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
Given a nonempty open subset $U\subset X$, we similarly define the mixed $\alpha$-level set
$K_{\alpha}^{\omega,U}=K_{\alpha}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}.$
Theorem 4.1.
Let $(X,d,f)$ be a t.d.s. with the specification property, and let $U\subset X$ be a nonempty open subset. Assume that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function satisfying Condition A. Then for any $\alpha\in L_{\varphi}$ and $S\in\{B,P\}$,
$\overline{\rm mdim}_{M}^{S}(f,K_{\alpha}^{\omega,U},d)=$ $\overline{\rm mdim}_{M}^{S}(f,K_{\alpha},d)$
$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
[…]
Proof.
Define $G_{\mu}^{\omega,U}:=G_{\mu}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}$. By (3.17), there exists a constant $c>0$ such that
$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu}^{\omega,U},\epsilon)$
holds for all $\mu\in M_{f}(X)$ and sufficiently small $\epsilon>0$. Together with the fact that $G_{\mu}^{\omega,U}\subset K_{\alpha}^{\omega,U}$ holds for any $\mu\in M_{f}(\varphi,\alpha)$, we have
$\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,K_{\alpha}^{\omega,U},\epsilon)$
for all sufficiently small $\epsilon>0$. This implies that
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\overline{\rm mdim}_{M}^{B}(f,K_{\alpha}^{\omega,U},d).$ (4.1)
Recall $G^{M_{f}(\varphi,\alpha)}=\{x\in X:V_{f}(x)\subset M_{f}(\varphi,\alpha)\}$. Fix $x\in X$. Since the statement $\varphi(\mu)=\alpha$ for all $\mu\in V_{f}(x)$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\varphi,\alpha)$, it follows that $K_{\alpha}=G^{M_{f}(\varphi,\alpha)}$. By Lemma 3.2 $(1)$, we get
$\overline{\rm mdim}_{M}^{P}(f,K_{\alpha},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (4.2)
By (4.1), (4.2), and the fact stated in Proposition 2.1 $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)$ for any $Z\subset X$, we complete the proof. ∎
A direct consequence of Theorem 4.1 is the variational principles for the Bowen metric mean dimension of level sets, as derived by Backes and Rodrigues [BR23, Theorem A]. However, we are able to strengthen this result by demonstrating that the variational principle remains valid for the packing metric mean dimension of level sets [^3] Readers can turn to [O03, Tho09, ZC13] and references therein for more examples about the multifractal analysis of level sets using topological pressure..
Corollary 4.2.
Let $(X,d,f)$ be a t.d.s. with the specification property and $\psi$ be a continuous function on $X$. Let
$E_{\alpha}:=\{x\in X:\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha\}$
denote the $\alpha$-level set of $\psi$. If $\alpha\in\mathbb{R}$ is a constant such that $E_{\alpha}\not=\emptyset$, then
$\overline{\rm mdim}_{M}^{B}(f,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(f,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\psi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
[…]
Proof.
[…]
$K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
Fix $x\in X$. Since the statement $\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\psi,\alpha)$, this shows $K_{\alpha}=E_{\alpha}$. We obtain the desired result by Theorem 4.1. ∎
4.2 Linking metric mean dimension and mean Li-Yorke chaotic
In the present subsection, we establish a connection between the concepts of metric mean dimension and mean Li-Yorke chaos within the context of infinite entropy. We provide a quantitative description of the chaotic behavior exhibited by infinite entropy systems. Specifically, we demonstrate that the packing and upper capacity metric mean dimensions of the set of mean Li-Yorke pairs are equal to twice the metric mean dimension of the phase space, as stated in Theorem 4.3.
Given a t.d.s. $(X,d,f)$, this naturally induces a product system $(X\times X,d\times d,f\times f)$, where $X\times X$ is a compact space equipped with the product metric defined as
$d\times d((x_{1},y_{1}),(x_{2},y_{2}))=\max\{d(x_{1},x_{2}),d(y_{1},y_{2})\},$
and the mapping $f\times f:X\times X\rightarrow X\times X$, given by $(f\times f)(x,y):=(fx,fy)$, is continuous.
A pair $(x,y)\in X\times X$ is called a mean Li-Yorke pair if
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)=0,~{}\text{and}\ \limsup_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)>0.$
The set of all mean Li-Yorke pairs is denoted by ${\rm MLY}(f)$. A subset $S$ of $X$ is called a mean Li-Yorke set if any two distinct points in $S$ form a mean Li-Yorke pair. We say that the system $(X,f)$ is mean Li-Yorke chaotic if there is an uncountable mean Li-Yorke set. The well-known fact is that positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]).
Theorem 4.3.
Let $(X,d,f)$ be a t.d.s. with specification property. Then for $S\in\{UC,P\}$ we have
$\overline{\rm mdim}_{M}^{S}(f\times f,{\rm MLY}(f),d\times d)=2~{}\overline{\rm mdim}_{M}(f,X,d)$
$\underline{\rm mdim}_{M}^{S}(f\times f,{\rm MLY}(f),d\times d)=2~{}\underline{\rm mdim}_{M}(f,X,d).$
Proof.
It suffices to show the first equality since the second one is obtained similarly.
We claim that there exists $\mu_{0}\in M_{f\times f}(X\times X)$ such that $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}(x,y)>0$.
Notice that the specification property implies that the system has positive topology entropy. Since positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]), we can choose two distinct points of $x,y\in X$, a sequence $\{n_{i}\}$ of positive integers that converges to $\infty$ as $i\to\infty$, and $\theta>0$ (depending on $x,y$) such that
$\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta.$
Set $\mu_{n_{i}}=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$. Then, in the weak-topology any accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfies the desired claim.
Now take $\nu\in M_{f}(X)$ arbitrarily and put $\mu_{1}:=\phi_{\ast}\nu=\nu\circ\phi^{-1}$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$. It is easy to check that $\mu_{1}\in M_{f\times f}(X\times X)$, and $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$. For sufficiently small $\gamma>0$, let
$c_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)-\gamma\}.$
By Theorem 2.4, we pick a sequence of positive real numbers $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and a sequence of invariant measures $m_{k}\in M_{f\times f}(X\times X)$ such that for each $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)>c_{\gamma}.$ (4.3)
Let $K=\overline{{\rm Conv}}\{\mu_{0},\mu_{1},m_{k},k\geq 1\}$, i.e., the convex closure of the set $\{\mu_{0},\mu_{1}\}\cup\{m_{k}:k\geq 1\}$. Then $K$ is a closed convex subset of $M_{f\times f}(X\times X)$. It is easy to see that if two systems admit the specification property, so does the product system. By [PS07], we have $G_{K}\not=\emptyset$. For any $(x_{0},y_{0})\in G_{K}$, there are two sequences of integers $\{t_{n}^{0}\}_{n\in\mathbb{N}}$ and $\{t_{n}^{1}\}_{n\in\mathbb{N}}$ with $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$, respectively. Hence,
$\lim_{n\to\infty}\frac{1}{t_{n}^{0}}\sum_{i=0}^{t_{n}^{0}-1}d(f^{i}x_{0},f^{i}y_{0})=\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}>0,$
and
$\lim_{n\to\infty}\frac{1}{t_{n}^{1}}\sum_{i=0}^{t_{n}^{1}-1}d(f^{i}x_{0},f^{i}y_{0})=\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0.$
This implies that $G_{K}\subset{\rm MLY}(f)$. Therefore, by Theorem 1.1, one has
$\overline{\rm mdim}_{M}^{P}(f\times f,{\rm MLY}(f),d\times d)\geq$ $\overline{\rm mdim}_{M}^{P}(f\times f,G_{K},d\times d)$
$=$ $\limsup_{\epsilon\to 0}\frac{\sup_{\nu\in K}\inf_{{\rm diam}(\eta,d\times d)<\epsilon}h_{\nu}(f\times f,\eta)}{|\log\epsilon|}$
$\geq$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)}{|\log\epsilon_{k}|}\geq c_{\gamma}.$
Since $\gamma$ is arbitrary, one has
$\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)\leq$ $\overline{\rm mdim}^{P}_{M}(f\times f,{\rm MLY}(f),d\times d)$
$\leq$ $\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)$
$\leq$ $\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d),$
where we used the Proposition 2.1 for the second inequality. Together with $\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)=2~{}\overline{\rm mdim}_{M}(f,X,d)$, we obtain the desired equality. ∎
Finally, we give an example to verify these results presented in Corollary 4.2 and Theorem 4.3.
Example 4.4.
Endow the product space $[0,1]^{\mathbb{Z}}$ with the product metric:
$d((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}})=\sum_{n\in\mathbb{Z}}\frac{|x_{n}-y_{n}|}{2^{|n|}},$
and let $\sigma:[0,1]^{\mathbb{Z}}\to[0,1]^{\mathbb{Z}}$ be the left shift map defined by $\sigma((x_{n})_{n\in\mathbb{Z}})=(x_{n+1})_{n\in\mathbb{Z}}$.
It is shown that the system $([0,1]^{\mathbb{Z}},d,\sigma)$ has the specification property (cf. [DGS76, Proposition 21.2]), and ${\rm{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d)=1$ [LT18].
  • (1)
    Thus, by Theorem 4.3, we have
    $\overline{\rm mdim}_{M}^{P}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=\overline{\rm mdim}_{M}^{UC}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=2.$
  • (2)
    Let $\varphi:[0,1]^{\mathbb{Z}}\to\mathbb{R}$ be a continuous map defined by $\varphi((x_{n})_{n\in\mathbb{Z}})=x_{0}$. Then for any $\alpha\in[0,1]$, by Corollary 4.2 we have
    $\overline{\rm mdim}_{M}^{B}(\sigma,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(\sigma,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{\sigma}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi),$
[…]
4.3 Generic points of full shift over compact metric space
In this subsection, based on the main results, we present some specific formulae for the metric mean dimensions of the set of generic points of full shifts over compact metric spaces.
We begin by reviewing some definitions and results for our statements. Let $(X,d,f)$ be a t.d.s., $\mu\in M_{f}(X)$, $\epsilon>0$ and $x\in X$. We define the Brin-Katok local entropy at $x$ as
$\overline{h}_{\mu}^{BK}(x,\epsilon)=\limsup_{n\to\infty}\frac{-\log\mu(B_{n}(x,\epsilon))}{n},$
and the Brin-Katok local $\epsilon$-entropy of $\mu$ as
$\overline{h}_{\mu}^{BK}(f,\epsilon):=\int\overline{h}_{\mu}^{BK}(x,\epsilon)\mathop{}\!\mathrm{d}\mu,$
respectively. Specially, if $\mu\in M_{f}^{e}(X)$, then $h_{\mu}^{BK}(x,\epsilon)=h_{\mu}^{BK}(f,\epsilon)$ holds for $\mu$-a.e. $x\in X$.
Example 4.5.
Let $K$ be a compact metric space with a metric $d$. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with the following product metric
$d^{\mathbb{Z}}\left((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}}\right)=\sum_{n\in\mathbb{Z}}\frac{d(x_{n},y_{n})}{2^{|n|}}.$
It is well-known that the system $(K^{\mathbb{Z}},\sigma)$ has the specification property. Moreover, it was proved in [VV17, Theorem 5] and [S22, Theorem 15] that,
$\overline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K,d),$ (4.4)
where $\overline{\rm dim}_{B}(K,d)$ is the upper box dimension of $K$ under the metric $d$.
Then, by Corollary 1.5, the metric mean dimensions of the generic points of the full shift over compact metric space are fully determined by the boxing dimension of the underlying space and mean Rényi information dimension of invariant measures.
Theorem 4.6.
Under the setting of Example 4.5, for any $\mu\in M_{\sigma}(K^{\mathbb{Z}})$ one has
$\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K,d)$
$\overline{\rm mdim}_{M}^{B}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi).$
Next, we present two examples of full shifts with the maximal metric mean dimension measure.
Example 4.7.
Consider the Example 4.4 again. Let $\mathcal{L}$ be the Lebesgue measure on $[0,1]$, and define $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ as the product measure on $[0,1]^{\mathbb{Z}}$. Shi proved that $\lim_{\epsilon\to 0}\frac{h_{\mu}^{BK}(\sigma,\epsilon)}{|\log\epsilon|}=1$ [S22, Example 12]. Together with [YCZ25, Theorem 1.1], we have
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi)=1={\rm\overline{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d).$
Then, by Corollary 1.5, we have
${\rm\overline{mdim}}_{M}^{B}(\sigma,G_{\mu},d)={\rm\overline{mdim}}_{M}^{P}(\sigma,G_{\mu},d)={\rm\overline{mdim}}_{M}^{UC}(\sigma,G_{\mu},d)=1.$
Example 4.8.
Endow $\mathbb{R}^{m}$ with the supremum norm $\|\cdot\|_{\mathbb{R}^{m}}$, and let $T_{i}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ $1\leq i\leq l$ be contracting similitudes, that is,
$\|T_{i}x-T_{i}y\|_{\mathbb{R}^{m}}=r_{i}\|x-y\|_{\mathbb{R}^{m}}$
for some $r_{i}\in(0,1)$. One says that the iterated function system $\{T_{i}\}_{1\leq i\leq l}$ satisfies the Open Set Condition if there is a bounded open subset $V\subset\mathbb{R}^{m}$ such that $T_{i}(V)\subset V,\ \forall 1\leq i\leq l$, and $T_{i}(V)\cap T_{j}(V)=\emptyset,\ \text{for }i\neq j.$
Under the OSC condition, it is well known that there are (unique) compact subset $K$, which we call the self-similar set, and a measure $\mu$, which we call the self-similar measure such that
  • (1)
    $K=\cup_{i=1}^{l}T_{i}(K)$;
  • (2)
    ${\rm dim}_{H}(K,d)={\rm dim}_{B}(K,d)=\alpha$, where $\alpha$ is the unique solution of equation
    $\sum_{i=1}^{l}r_{i}^{\alpha}=1,$
    and ${\rm dim}_{H}$ denotes the Hausdorff dimension under the metric $d:=\|\cdot\|_{\mathbb{R}^{m}}$. $\alpha$ is called the similarity dimension of the IFS;
  • (3)
    $\mu=\sum_{i=1}^{l}r_{i}^{\alpha}\cdot\mu\circ T_{i}^{-1}$;
  • (4)
    ${\rm supp}(\mu)=K$.
Many famous fractal sets, for instance, the middle thirds Cantor set, the Sierpinński triangle or gasket, are the attractors of iterated function systems satisfying the Open Set Condition.
Now take the attractor $K$ obtained above as in the Example 4.5. Then the product measure $\nu:=\mu^{\otimes\mathbb{Z}}$ on $K^{\mathbb{Z}}$ is invariant with respect to the left shift $\sigma$.
Theorem 4.9.
Under the setting of Example 4.8, $\nu$ is the maximal metric mean dimension such that the three types of metric mean dimensions of $G_{\nu}$ coincide with the similarity dimension. i.e.,
$\overline{\rm mdim}_{M}^{S}(\sigma,G_{\nu},d^{\mathbb{Z}})=\alpha,$
where $S\in\{B,P,UC\}$
Proof.
Hutchinson [Hut81] showed that for any $z\in K$ and $\epsilon\in(0,1)$, one has
$\mu(B(z,\epsilon))\leq C\epsilon^{\alpha},$ (4.5)
where $C>0$ is a constant independent of the choice of the point  $z\in K$ and  $\epsilon$. For any $x=(x_{k})_{k\in\mathbb{Z}}\in K^{\mathbb{Z}}$ and $\epsilon\in(0,1)$, we define
[…]
Then the Bowen ball $B_{n}(x,\epsilon)$ is contained in $L_{n}(x,\epsilon)$ and $\nu(L_{n}(x,\epsilon))\leq C^{n}\epsilon^{n\alpha}$ by (4.5). Thus, one has
$\nu(B_{n}(x,\epsilon))\leq C^{n}\epsilon^{n\alpha},\ \forall x\in K^{\mathbb{Z}},$
which yields that $\overline{h}_{\nu}^{BK}(\sigma,\epsilon)\geq\alpha|\log\epsilon|-\log C.$ By the Shannon-McMillan-Breiman theorem, the following inequality is valid for any t.d.s. $(X,d,f)$:
$\overline{h}_{\mu}^{BK}(f,\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),\ \forall\mu\in M_{f}(X).$ (4.6)
Hence
[…]
which implies the desired result. ∎
Acknowledgments: The first and second authors were supported by NNSF of China (12101340). The third author was supported by the China Postdoctoral Science Foundation (No. 2024M763856); R.Yang would like to thank Yunping Wang for many helpful communication and the warm hospitality when he was visiting the Ningbo University. The authors also thank Prof. Ercai Chen for many useful comments.