### v4

# Yong Ji1, Junye Li1 and Rui Yang∗2,3

Abstract. From a geometric perspective, we employ metric mean dimension to investigate the set of generic points of invariant measures and saturated sets in infinite entropy systems. For systems with the specification property, we establish certain variational principles for the Bowen and packing metric mean dimensions of saturated sets in terms of Kolmogorov-Sinai $\epsilon$-entropy, and prove that the upper capacity metric mean dimension of saturated sets has full metric mean dimension. Consequently, the Bowen and packing metric mean dimensions of the set of generic points of invariant measures coincide with the mean Rényi information dimension, and the upper capacity metric mean dimension of the set of generic points of invariant measures also has full metric mean dimension.
As applications, for systems with the specification property, we present the qualitative characterization of the metric mean dimensions of level sets, the set of mean Li-Yorke pairs in infinite-entropy systems, and the set of generic points of invariant measures in full shifts over compact metric spaces.
Keywords: multifractal analysis; metric mean dimension; saturated set; generic point; variational principle
1 Introduction
Let $(X,d,f)$ be a topological dynamical system (TDS for short), where $(X,d)$ is a compact metric space, and $f:X\to X$ is a homeomorphism. Let $M(X)$, $M_{f}(X)$, $M_{f}^{e}(X)$ denote the sets of Borel probability measures on $X$ endowed with the weak-topology, $f$-invariant, and $f$-ergodic Borel probability measures on $X$, respectively.
Quantitative analysis of dynamical systems aims to use entropy-like quantities to understand the dynamics of the abstract dynamical systems. Among these entropy-like quantities, measure-theoretic entropy and topological entropy are two topological invariants to characterize the topological complexity of dynamical systems, which are related by the classical variational principle:
$h_{top}(f)=\sup_{\mu\in M_{f}(X)}{h_{\mu}(f)},$
where $h_{top}(f)$ denotes the topological entropy of $X$, and $h_{\mu}(f)$ is the measure-theoretic entropy of $\mu$. To extend the concept of topological entropy to any non-empty subsets, resembling the definition of Hausdorff dimension, Bowen [Bow73] in his pioneering work introduced Bowen topological entropy on subsets. It turns out that the Bowen topological entropy provides a powerful tool for studying the “size” of fractal-like sets and has directly contributed to the development of multifractal analysis in dynamical systems. The multifractal analysis, with its origin dating back to Besicovitch, takes into account the decomposition of the whole phase space into smaller subsets consisting of points with the similar dynamical behaviors. And then, from a geometrical and topological perspectives, we use the fractal dimension and entropy-like quantity to describe the “size” of these subsets. Such a procedure allows us to obtain more information about the dynamics and even recovers the partial dynamics. This phenomenon is also known as multifractal rigidity.
Typically, infinite entropy systems exhibit rather complicated topological complexity. As a result, the concept of topological entropy fails to offer more information about the dynamics of the system. In 2000, Lindenstrauss and Weiss [LW00] introduced the metric mean dimension, which is a dynamical analogue of box dimension. The measure-theoretic counterpart of metric mean dimension, called upper and lower mean Rényi information dimensions, is due to Gutman and Śpiewak[GS21]. The analogous variational principles for metric mean dimension can be found in the references [LT18, GS21, YCZ25, S22]. Given $x\in X$, denote by
$\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}$
the $n$-th empirical measure of $x$, where $\delta_{x}$ is the Dirac measure at $x$. It is well-known that, in the weak-topology, the limit-point set $V_{f}(x)$ of the sequence $\{\mathcal{E}_{n}(x)\}_{n\geq 1}$ is a non-empty compact connected subset of $M_{f}(X)$ . In particular, if $V_{f}(x)=\{\mu\}$ for some $f$-invariant measure $\mu$, then $x$ is called a generic point of $\mu$. Denote by $G_{\mu}$ the set of all generic points of $\mu$. The Birkhoff ergodic theorem shows that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic.
We aim to study the multifractal analysis of the set of generic points in infinite entropy systems. Let us mention some results about the topological entropy of the set of generic points. For every ergodic measure $\mu$, Bowen [Bow73] proved that the Bowen topological entropy of $G_{\mu}$ equals $h_{\mu}(f)$, i.e.,
$h_{top}^{B}(f,G_{\mu})=h_{\mu}(f).$
For the non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$, and there exist the examples showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$. Therefore, Bowen’s result fails for non-ergodic measures. In 2007, Pfister and Sullivan [PS07] proved that Bowen’s equality still holds for non-ergodic measures within a more general framework. More generally, given a non-empty closed connected subset $K\subset M_{f}(X)$, the saturated set $G_{K}$ associated with $K$ is defined by
$G_{K}=\{x\in X:V_{f}(x)=K\}.$
By introducing the $g$-almost product property (e.g. $\beta$-shifts), which is strictly weaker than the specification property, and the uniform separation property (e.g. expansive and more generally asymptotically $h$-expansive systems), they established a variational principle for $G_{K}$:
$h_{top}^{B}(T,G_{K})=\inf_{\mu\in K}h_{\mu}(f).$
Furthermore, under only the assumption of the $g$-almost product property, they [PS07] showed that $h_{top}^{B}(f,G_{\mu})=h_{\mu}(f)$ for all $\mu\in M_{f}(X)$. For systems with the specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] also obtained a (dual) variational principle for packing topological entropy of $G_{K}$:
$h_{top}^{P}(f,G_{K})=\sup_{\mu\in K}h_{\mu}(f).$
Later, for systems with the $g$-almost product property, Hou, Tian and Zhang [HTZ23] proved that the upper capacity topological entropy of the saturated sets with more complicated dynamics has full topological entropy, i.e.,
$h_{top}^{UC}(f,G_{K}^{C})=h_{top}(f),$
where $G_{K}^{C}:=G_{K}\cap\{x\in X:C_{f}(X)\subset\omega_{f}(x)\},$ $\omega_{f}(x):=\bigcap_{n=0}^{\infty}\overline{\bigcup_{k=n}^{\infty}\{f^{k}x\}}$ is the $\omega$-limit set of $x$, and $C_{f}(X)$ is the measure center given by $C_{f}(X):=\overline{\cup_{\mu\in M_{f}(X)}{\rm supp}{(\mu)}}$, ${\rm supp}(\mu)$ is the supported set of $\mu$. Then, in the context of the infinite entropy systems, the aforementioned work suggests the following questions for saturated sets:
Questions:
  • (1)
    For ergodic measures, does the metric mean dimension of the set of generic points coincide with the mean Rényi information dimension?
  • (2)
    For non-ergodic measures, under the assumption of specification-like property, are there certain variational principles for metric mean dimensions of saturated sets? If not, can metric mean dimensions of saturated sets have the full metric mean dimension?
For (1), the authors in [YCZ25] proved that the packing metric mean dimension of the set of generic points equals the mean Rényi information dimension. However, this equality holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula within the context of infinite entropy systems ([YCZ25, §5, Question,(1)]); for (2), since the uniform separation property implies the system has finite topological entropy, extending the work of [PS07, ZCC12, HTZ23] to the framework of infinite entropy systems encounters some nontrivial challenges. Actually, since the asymptotic behavior of orbits ${\rm orb}_{f}(x)=\{f^{n}(x):n\geq 0\}$ is a fundamental topic in ergodic theory, the subsequent statements of main results involve more general (mixed) saturated sets. For simplicity, we write
$G_{K}^{\omega,U}:=G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},$
where $U\subset X$ is a non-empty open subset, $K\subset M_{f}(X)$ is a non-empty closed connected (or convex) subset. If $G_{K}^{\omega,U}\not$$\emptyset$, then there exists a transitive point in the sufficiently “small” open set $U$ such that $V_{f}(x)=K$. For systems with the specification property, we provide a complete answer to question (2).
Our main results clarify that the saturated sets exhibit different dynamical behaviors in terms of different types of metric mean dimensions. There are certain variational principles for the Bowen and packing metric mean dimensions of saturated sets, as presented in Theorems 1.1, 1.2 and 1.3.
Let $\overline{\rm mdim}_{M}^{B}(f,Z,d)$, $\overline{\rm mdim}_{M}^{P}(f,Z,d)$ and $\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ denote the Bowen/packing/upper capacity metric mean dimension of $Z$, respectively.
Theorem 1.1.
Let $(X,d,f)$ be a TDS with the specification property. If $K\subseteq M_{f}(X)$ is a non-empty compact convex set, and $U\subset X$ is a non-empty open set, then
$\overline{\rm mdim}_{M}^{P}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Theorem 1.2.
Let $(X,d,f)$ be a TDS with the specification property. If $K\subseteq M_{f}(X)$ is a non-empty compact connected set, and $U\subset X$ is a non-empty open set, then
$\overline{\rm mdim}_{M}^{B}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Theorem 1.3.
Let $(X,d,f)$ be a TDS with the specification property. If $K\subseteq M_{f}(X)$ is a non-empty compact convex set, then
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\underline{\rm mrid}_{\mu}^{B}(f,K,d),$
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\overline{\rm mrid}_{\mu}^{P}(f,K,d),$
where $\underline{\rm mrid}_{\mu}^{B}(f,K,d)$ and $\overline{\rm mrid}_{\mu}^{P}(f,K,d)$ denotes the Bowen and packing Rényi information dimensions of $\mu$, respectively.
We remark that any convex set in a topological vector space is automatically connected. Although the $\mu$-measure of $G_{\mu}$ is zero for every non-ergodic measure $\mu$, its topological complexity can be sufficiently large, that is, the upper capacity metric mean dimension of saturated sets carries the full metric mean dimension, as stated in Theorem 1.4.
Theorem 1.4.
Let $(X,d,f)$ be a TDS with the specification property. Let $K\subseteq M_{f}(X)$ be a non-empty compact connected set, and $U\subset X$ be a non-empty open set. Then
$\overline{\rm mdim}_{M}^{UC}\left(f,G_{K},d\right)=\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}(f,X,d),$
where $\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ is the upper capacity metric mean dimension of $Z$.
As an immediate consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case occurs when $K\subset M_{f}(X)$ consists of exactly a single point. The following Corollary 1.5 extends the result [YCZ25, Theorem 1.1] to all non-ergodic measures for systems with the specification property.
Corollary 1.5.
Let $(X,d,f)$ be a TDS with the specification property. Then for every $\mu\in M_{f}(X)$,
$\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}^{P}_{M}(f,G_{\mu},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi),$
$\overline{\rm mdim}_{M}^{UC}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
Except Theorem 1.3, all theorems mentioned above are valid for the corresponding lower metric mean dimensions by considering changing the $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$. Besides, it becomes unclear whether the aforementioned results for weaker specification-like property.
The paper is organized as follows. In section 2, we recall the precise definitions of metric mean dimensions and the associated properties, and two types of measure-theoretic $\epsilon$-entropies, and the specification property of TDSs. In section 3, we prove the main results. In section 4, we consider some applications of the main results, and apply it to investigate the level sets of the Birkhoff average of a continuous potential, Li-Yorke chaos in finite entropy systems and the metric mean dimensions of generic points of the full shift over compact metric spaces.
2 Preliminaries
In this section, we recall the concepts of metric mean dimensions in both topological and measure-theoretic situations, and state the precise definition of the specification property in dynamical systems.
2.1 Three types of metric mean dimensions
In this subsection, we review the definitions of upper capacity metric mean dimension of subsets[LW00], Bowen metric mean dimension[LV21, W21] and packing metric mean dimension [YCZ22] of subsets.
Let $(X,d,f)$ be a TDS. For any natural number $n\in\mathbb{N}$, we define the $n$-th Bowen metric $d_{n}$ on $X$ as
$d_{n}(x,y):=\max_{0\leq i\leq n-1}d(f^{i}x,f^{i}y),$
where $x,y\in X$. Then the Bowen open ball and closed ball centered at $x$ with radius $\epsilon$ in the metric $d_{n}$ are given by
$B_{n}(x,\epsilon):$ $=\{y\in X:d_{n}(x,y)<\epsilon\},$
$\overline{B}_{n}(x,\epsilon):$ $=\{y\in X:d_{n}(x,y)\leq\epsilon\}$
respectively.
Let $Z$ be a non-empty subset of $X$. Fix $\epsilon>0$. A set $E\subset Z$ is said to be an $(n,\epsilon)$-separated set of $Z$ if any distinct $x,y\in E$ implies $d_{n}(x,y)>\epsilon$. Denote by $s_{n}(Z,\epsilon)$ the largest cardinality of $(n,\epsilon)$-separated sets of $Z$. The $\epsilon$-upper capacity topological entropy of $Z$ is given by
$h_{top}^{UC}(f,Z,\epsilon)=\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(Z,\epsilon).$
Recall that the upper capacity topological entropy of $Z$ is the limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 0, i.e.,
$h_{top}^{UC}(f,Z)=\lim_{\epsilon\to 0}h_{top}^{UC}(f,Z,\epsilon)=\sup_{\epsilon>0}h_{top}^{UC}(f,Z,\epsilon).$
To get more information about the dynamics of infinite entropy systems, analogous to the definition of box dimension in fractal geometry, we formulate a refined entropy-like quantity using the $\epsilon$-upper capacity topological entropy.
Definition 2.1.
The upper capacity metric mean dimension of $Z$ is defined by
$\overline{\rm mdim}_{M}^{UC}(f,Z,d):=\limsup_{\epsilon\to 0}\frac{h_{top}^{UC}(f,Z,\epsilon)}{|\log\epsilon|}.$
If $Z=X$, the upper capacity metric mean dimension of $X$ is reduced to the metric mean dimension introduced by Lindenstrauss-Weiss [LW00]; in this case, we use the notation $h_{top}(f,X,\epsilon)$ instead of $h_{top}^{UC}(f,X,\epsilon)$, and $\overline{\rm mdim}_{M}(f,X,d)$ instead of $\overline{\rm mdim}_{M}^{UC}(f,X,d)$, respectively. Such a formulation extends the concept of metric mean dimension to any subsets of the phase space, which is not necessarily compact or $f$-invariant. Notice that the (upper capacity) metric mean dimension is defined by measuring how fast the $\epsilon$-upper capacity topological entropy diverges to $\infty$ as $\epsilon\to 0$. That is, for sufficiently small $\epsilon>0$, we may think of
$h_{top}^{UC}(f,Z,\epsilon)\approx\overline{\rm mdim}_{M}^{UC}(f,Z,d)\cdot|\log\epsilon|.$
It is easy to see that any system with finite topological entropy has zero metric mean dimension. This shows that the metric mean dimension is a suitable “measure” for capturing the dynamics of the infinite entropy systems.
The Bowen topological entropy[Bow73] and packing topological entropy[FH12] in dimension theory, which are defined by Carathéodory-Pesin structures, are the analogues of the Hausdorff dimension and packing dimension in fractal geometry. The corresponding dimensional characterizations for metric mean dimension, which have been introduced in [LV21, YCZ22, W21] through Carathéodory-Pesin structures, are the following Bowen and packing metric mean dimensions.
For $Z\subset X$, $s>0$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
$M(Z,s,N,\epsilon)=\inf\left\{\sum_{i\in I}\exp(-n_{i}s)\right\},$
where the infimum is taken over all finite or countable families $\{B_{n_{i}}(x_{i},\epsilon)\}_{i\in I}$ such that $Z\subset\cup_{i\in I}B_{n_{i}}(x_{i},\epsilon)$ with $x_{i}\in X,n_{i}\geq N$ for all $i\in I$.
The quantity $M(Z,s,N,\epsilon)$ is non-decreasing as $N$ increases. So the limit
$M(Z,s,\epsilon)=\lim_{N\to\infty}M(Z,s,N,\epsilon)$
exists. There is a critical value of the parameter $s$, which we call $\epsilon$-Bowen topological entropy, such that $M(Z,s,\epsilon)$ jumps from $\infty$ to 0, i.e.,
$h_{top}^{B}(f,Z,\epsilon)$ $=\inf\{s>0:M(Z,s,\epsilon)=0\}$
$=\sup\{s>0:M(Z,s,\epsilon)=\infty\}.$
Definition 2.2.
The Bowen upper metric mean dimension of $Z$ is defined by
$\overline{\rm mdim}^{B}_{M}(f,Z,d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{B}(f,Z,\epsilon)}{|\log\epsilon|}.$
For $Z\subseteq X$, $s>0$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
$P(Z,s,N,\epsilon)=\sup\left\{\sum_{i\in I}\exp(-sn_{i})\right\},$
where the supremum is taken over all finite or countable pairwise disjoint families $\{\overline{B}_{n_{i}}(x_{i},\epsilon)\}_{i\in I}$ with $x_{i}\in Z$, $n_{i}\geq N$ for all $i\in I$.
Since the quantity $P(Z,s,N,\epsilon)$ is non-increasing as $N$ increases, and thus the following limit exists:
$P(Z,s,\epsilon)=\lim_{N\to\infty}P(Z,s,N,\epsilon).$
Observe that the quantity $P(Z,s,\epsilon)$ does not satisfy the countable additive property with respect to $Z$. So the modifications are needed to obtain this property:
$\mathcal{P}(Z,s,\epsilon):=\inf\left\{\sum_{i=1}^{\infty}P(Z_{i},s,\epsilon):\bigcup_{i=1}^{\infty}Z_{i}\supseteq Z\right\}.$
There is a critical value of the parameter $s$, which we call $\epsilon$-packing topological entropy, such that $\mathcal{P}(Z,s,\epsilon)$ jumps from $\infty$ to 0. That is, the critical value is given by
$h_{top}^{P}(f,Z,\epsilon)$ $=\inf\{s>0:\mathcal{P}(Z,s,\epsilon)=0\}$
$=\sup\{s>0:\mathcal{P}(Z,s,\epsilon)=\infty\}.$
Definition 2.3.
The packing upper metric mean dimension of $Z$ is defined by
$\overline{\rm mdim}_{M}^{P}(f,Z,d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{P}(f,Z,\epsilon)}{|\log\epsilon|}.$
One can replace $\limsup_{\epsilon\to 0}$ by $\liminf_{\epsilon\to 0}$ in above definitions to obtain the corresponding three types of lower metric mean dimensions on subsets, which are denoted by $\underline{\rm mdim}_{M}^{S}(f,Z,d)$ with $S\in\{UC,B,P\}$. If $\underline{\rm mdim}_{M}^{S}(f,Z,d)=\overline{\rm mdim}_{M}^{S}(f,Z,d)$, we call the common value ${\rm mdim}_{M}^{S}(f,Z,d)$ the (capacity/Bowen/packing) metric mean dimension of $Z$.
We summarize some fundamental properties of metric mean dimensions derived in [YCZ22, Proposition 3.4] and [YCZ23, Proposition 2.8].
Proposition 2.1.
$(1)$ If $Z_{1}\subset Z_{2}\subset X$, then for each $S\in\{UC,B,P\}$,
$\overline{\rm mdim}_{M}^{S}(f,Z_{1},d)\leq\overline{\rm mdim}_{M}^{S}(f,Z_{2},d).$
$(2)$ If $Z=\bigcup_{i=1}^{\infty}Z_{i}$, then for every $\epsilon>0$,
$h_{top}^{B}(f,Z,\epsilon)$ $=\sup_{i\geq 1}h_{top}^{B}(f,Z_{i},\epsilon),$
$h_{top}^{P}(f,Z,\epsilon)$ $=\sup_{i\geq 1}h_{top}^{P}(f,Z_{i},\epsilon),$
and $h_{top}^{UC}(f,Z,\epsilon)=\max_{1\leq i\leq N}h_{top}^{UC}(f,Z_{i},\epsilon)$ if $Z=\bigcup_{i=1}^{N}Z_{i}$.
$(3)$ If $Z=\bigcup_{i=1}^{N}Z_{i}$, then for each $S\in\{UC,B,P\}$,
$\overline{\rm mdim}_{M}^{S}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}_{M}^{S}(f,Z_{i},d).$
$(3)$ If $Z$ is a non-empty subset of $X$, then for every $0<\epsilon<1$, we have
$h_{top}^{B}(f,Z,3\epsilon)\leq h_{top}^{P}(f,Z,\epsilon)\leq h_{top}^{UC}(f,Z,\epsilon).$
Consequently, $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)\leq\overline{\rm mdim}_{M}^{UC}(f,Z,d)$.
Furthermore, if $Z$ is $f$-invariant and compact, then the three types of metric mean dimensions coincide.
Given a certain subset of phase space, in general it is not easy to calculate its precise Bowen and packing metric mean dimensions. The Entropy Distribution Principles allow us to get a lower bound for $\epsilon$-Bowen and $\epsilon$-packing topological entropies of certain fractal-like sets, and hence yields the lower bounds of metric mean dimension of the fractal-like sets.
The Entropy Distribution Principles of $\epsilon$-Bowen and packing topological entropies are stated as follows:
Lemma 2.1.
[BR23, Lemma 13] Let $(X,d,f)$ be a TDS and $\epsilon>0$. Suppose that $Z$ is a Borel set of $X$ and $0<s<\infty$. If there exist $\mu\in M(X)$ and a constant $C>0$ such that $\mu(Z)>0$ and $\mu(B_{n}(x,\epsilon))\leq Ce^{-ns}$ for every each Bowen ball $B_{n}(x,\epsilon)\cap Z\neq\emptyset$, then
$h_{top}^{B}(f,Z,\epsilon)\geq s.$
Lemma 2.2.
Let $(X,d,f)$ be a TDS and $\epsilon>0$. Suppose that $Z$ is a Borel set of $X$ and $0<s<\infty$. If there exist $\mu\in M(X)$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ such that $\mu(Z)>0$ and $\mu(B_{n_{i}}(x,\epsilon))\leq Ce^{-n_{i}s}$ for each $i$ and for any $x\in Z$, then
$h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s.$
Proof.
By definition of the $\epsilon$-packing topological entropy, it is enough to prove that for every sufficiently small $\delta>0$,
$P\left(E,s-\delta,\frac{\epsilon}{6}\right)=\infty$
for any Borel subset $E\subset Z$ with $\mu(E)>0$.
For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of $E$. Then, by $5r$-covering lemma [Mat95, Theorem 2.1], there exists a countable pairwise disjoint sub-collection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, where $J$ is a index set, such that $E\subset\bigcup_{j\in J}B_{n_{i}}(x_{j},\epsilon).$ In particular, the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint. Hence, one has
$P\left(E,s-\delta,n_{i},\frac{\epsilon}{6}\right)\geq\sum_{j\in J}\exp[-n_{i}(s-\delta)]\geq\frac{e^{n_{i}\delta}}{C}\sum_{j\in J}\mu(B_{n_{i}}(x_{j},\epsilon))\geq\frac{e^{n_{i}\delta}}{C}\mu(E).$
Letting $i\to\infty$, we have $P(E,s-\delta,\frac{\epsilon}{6})=\infty$. This shows that $\mathcal{P}(Z,s-\delta,\frac{\epsilon}{6})=\infty$, and hence $h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s$ by letting $\delta\to 0$. ∎
2.2 Measure-theoretic $\epsilon$-entropy of invariant measures
In this subsection, we recall the definitions of Kolmogorov-Sinai $\epsilon$-entropy and Katok $\epsilon$-entropy of invariant measures.
Let $C(X)$ denote the space of real-valued continuous functions on $X$, equipped with the supremum norm $||\cdot||$. By virtue of the Riesz representation theorem for compact metric spaces, we are able to endow $M(X)$ with the weak-topology, which is metrizable by the following metric:
$D(\mu,\nu):=\sum_{k=1}^{\infty}\frac{|\int\phi_{k}\mathop{}\!\mathrm{d}\mu-\int\phi_{k}\mathop{}\!\mathrm{d}\nu|}{2^{k}},$ (2.1)
where $\{\phi_{k}\}$ is a countable family of dense subsets of $C(X)$ taking values in $[0,1]$. Thus, the diameter of $M(X)$ in the sense of the metric $D$ is always less than $1$, and for fixed $\mu,\nu\in M(X)$, both $D(\mu,\cdot)$ and $D(\cdot,\nu)$ are convex functions on $M(X)$. Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in the weak-topology.
Given $\mu\in M_{f}(X)$ and a finite Borel partition $\xi=\left\{C_{1},\cdots,C_{k}\right\}$, the partition entropy of $\xi$ w.r.t. $\mu$ is defined by $H_{\mu}(\xi)=-\sum_{i=1}^{k}\mu(C_{i})\log(\mu(C_{i})).$ We define the measure-theoretic entropy of $\mu$ w.r.t. $\xi$ as
$h_{\mu}(f,\xi)=\lim_{n\to+\infty}\frac{1}{n}H_{\mu}(\xi^{n}),$
Denote by $\xi^{n}:=\vee_{i=0}^{n-1}f^{-i}\xi$ is the $n$-th join of the preimage partitions $\xi,f^{-1}\xi,\ldots,f^{-(n-1)}\xi$ on $X$. The measure-theoretic entropy of $\mu$ is defined as
$h_{\mu}(f)=\sup_{\xi}h_{\mu}(f,\xi),$
where the supremum ranges over all finite Borel partitions of $X$.
For topological dynamical systems, the measure-theoretic counterpart of metric mean dimension is the measure-theoretic metric mean dimension of invariant measures. It measures the divergent rate of the limits of the measure-theoretic $\epsilon$-entropies such as Brin-Katok $\epsilon$-entropy, Katok $\epsilon$-entropy, Kolmogorov-Sinai $\epsilon$-entropy, and the others. A systemic study about the measure-theoretic $\epsilon$-entropies is stated in [YCZ25]. Here, we only recall two types of the measure-theoretic $\epsilon$-entropies.
Let $\mu\in M_{f}(X)$ and $\epsilon>0$.
  • (a)
    The Kolmogorov-Sinai $\epsilon$-entropy [GS21] of $\mu$ is given by
    $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
    where the infimum is taken over all finite Borel partitions of $X$ with diameter less than $\epsilon$.
  • (b)
    For $\delta\in(0,1)$ and $n\in\mathbb{N}$, we put
    $N_{\mu}(n,\epsilon,\delta):=\min\{\#E:E\subset X~\emph{and}~\mu(\cup_{x\in E}B_{n}(x,\epsilon))>1-\delta\}.$
    The Katok $\epsilon$-entropy of $\mu$ [Kat80] is defined by
    $\underline{h}_{\mu}^{K}(f,\epsilon,\delta)=\liminf_{n\to\infty}\frac{1}{n}\log N_{\mu}(n,\epsilon,\delta).$
The Katok $\epsilon$-entropy and Kolmogorov-Sinai $\epsilon$-entropy are related by the following inequalities:
Lemma 2.3.
[S22, Lemma 4.1] Let $(X,d,f)$ be a TDS and $\mu\in M_{f}^{e}(X)$, If $\mathcal{U}$ is a finite open cover of $X$ with the diameter ${\rm diam}(\mathcal{U})\leq 4\epsilon$ and the Lebesgue number ${\rm Leb}(\mathcal{U})\geq\epsilon$, then for any $\delta\in(0,1)$,
$\underline{h}_{\mu}^{K}(f,4\epsilon,\delta)\leq\inf_{\xi\succ\mathcal{U}}h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta),$
where the infimum $\xi$ ranges over all finite Borel partitions of $X$ refining than $\mathcal{U}$.
Consequently, for every $\epsilon>0$, one has $\underline{h}_{\mu}^{K}(f,16\epsilon,\delta)\leq\inf_{{\rm diam}(\xi)<4\epsilon}\limits h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta).$
We state the variational principles between the metric mean dimension and the above two types of measure-theoretic $\epsilon$-entropies.
Theorem 2.4.
(cf. [YCZ25, Theorems 1.2 and 1.3]) Let $(X,d,f)$ be a TDS. Then for any $\delta\in(0,1)$,
$\overline{\rm mdim}_{M}(f,X,d)$ $=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(X)}F(\mu,\epsilon),$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}^{e}(X)}F(\mu,\epsilon),$
where $F(\mu,\epsilon)\in\{\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi),\underline{h}^{K}_{\mu}(f,\epsilon,\delta)\}$.
The results are also valid for $\underline{\rm mdim}_{M}(f,X,d)$ by changing $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.
2.3 Specification property of TDSs
In general, it is difficult to find a real orbit with the desired dynamical behavior, while one can get an approximate orbit of a point that $\epsilon$-shadows any specified finite orbit-segments, which is known as the specification property in TDSs.
Definition 2.4.
One says that a TDS $(X,d,f)$ has the specification property if for any $\epsilon>0$, there is a positive integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists a point $x\in X$ such that
$d(f^{p}x_{j},f^{p+a_{j}}x)<\epsilon$
for any $p=0,\cdots,b_{j}-a_{j}$, $j=1,\cdots,k$.
In other words, the point $x$ $\epsilon$-shadows the $k$ pieces of finite orbits of $x_{j}$:
$\{x_{j},f(x_{j}),\ldots,f^{b_{j}-a_{j}}(x_{j})\},\ j=1,\ldots,k.$
For instance, every topological mixing locally maximal hyperbolic set, and the full shift on $0,1]^{\mathbb{Z}}$ have the specification property. More examples of infinite entropy systems with the specification property can be found in [BR23, Section IV].
A point $x\in X$ is called almost periodic if for any open neighborhood $U$ of $x$, there exists $N\in\mathbb{N}$ such that for every $n\in\mathbb{N}$, there exists an integer $k$ with $n\leq k\leq n+N$ for which $f^{k}(x)\in U$. Denote by $AP(X)$ the set of all almost periodic points of $X$.
Proposition 2.2.
[HTZ23, Proposition 2.11] If a TDS $(X,d,f)$ has the specification property, then the almost periodic set AP(X) is dense in $X$.
3 Proofs of main results
In this section, we prove the main results of this paper.
3.1 Packing metric mean dimension of saturated sets
To prove Theorem 1.1, we divide its proof into two parts.
3.1.1 Upper bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K},d)$
Lemma 3.1.
[Y23, Lemma 3.1] Let $(X,d,f)$ be a TDS and $\epsilon>0$. Let $\{F_{n}\}$ be a sequence of non-empty subsets of $X$, and let $E_{n}$ be an $(n,\epsilon)$-separated subset of $F_{n}$ with the largest cardinality. We define
$v_{n}:=\frac{1}{\#E_{n}}\sum_{x\in E_{n}}\mathcal{E}_{n}(x).$
Then for any limit point $\mu$ of the sequence $\{\nu_{n}\}$, one has
$\limsup_{n\to\infty}\frac{1}{n}\log\#E_{n}\leq\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
It essentially follows from the proof of the upper bound for the classical variational principle (cf.[Wal82, Theorem 8.6]). ∎
Lemma 3.2.
Let $(X,d,f)$ be a TDS, and let $K\subset M_{f}(X)$ be a non-empty closed convex subset. Then for any $\epsilon>0$,
$h_{top}^{P}(f,G_{K},\epsilon)\leq h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where $G^{K}:=\{x\in X:V_{f}(x)\subset K\}.$
Consequently, for every $\mu\in M_{f}(X)$ and $\epsilon>0$,
$h_{top}^{P}(f,G_{\mu},\epsilon)\leq\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
It suffices to show the inequality
$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
For $n\in\mathbb{N}$ and $\delta>0$, we put
$\mathcal{R}_{K}(\delta,k):$ $=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta),~\forall n\geq k\right\},$
$\mathcal{Q}_{K}(\delta,n):$ $=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)\right\},$
and define
$\Theta(K,\epsilon):=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta,n),\epsilon),$
where $s_{n}(\mathcal{Q}_{K}(\delta,n),\epsilon)$ is the largest cardinality of $(n,\epsilon)$-separated sets of $\mathcal{Q}_{K}(\delta,n)$.
For every $x\in G^{K}$, we have $V_{f}(x)\subset B(K,\delta)$ for every $\delta>0$. Then the cardinality of $n$ such that $\mathcal{E}_{n}(x)\in M_{f}(X)\backslash B(K,\delta)$ must be finite. This shows that $x\in\mathcal{R}_{K}(\delta,k)$ for some $k$. By $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}_{K}(\delta,k)$ for every $\delta>0$, it follows from Proposition 2.1 that
$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{k}h_{top}^{P}(f,\mathcal{R}_{K}(\delta,k),\epsilon)\leq\sup_{k}h_{top}^{UC}(f,\mathcal{R}_{K}(\delta,k),\epsilon).$ (3.1)
For any $\tau>0$, there exists $\delta_{0}>0$ such that
$\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta_{0},n),\epsilon)<\Theta(K,\epsilon)+\tau.$
For any $k\geq 1$, it holds that $\mathcal{R}_{K}(\delta_{0},k)\subset\mathcal{Q}_{K}(\delta_{0},n)$ for any $n\geq k$. Thus, we have
$h_{top}^{UC}(f,\mathcal{R}_{K}(\delta_{0},k),\epsilon)\leq\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta_{0},n),\epsilon)<\Theta(K,\epsilon)+\tau.$
Using (3.1) and letting $\tau\to 0$, we conclude that $h_{top}^{P}(f,G^{K},\epsilon)\leq\Theta(K,\epsilon)$.
Take a strictly decreasing sequence $\{\delta_{k}\}$ with $\delta_{k}\to 0$ and a strictly increasing sequence $\{n_{k}\}$ with $n_{k}\to\infty$ such that
$s_{n_{k}}(\mathcal{Q}_{K}(\delta_{k},n_{k}),\epsilon)>\exp[n_{k}(\Theta(K,\epsilon)-\gamma)].$
Let $F_{k}$ be a $(n_{k},\epsilon)$-separated subset of $\mathcal{R}_{K}(\delta_{k},n_{k})$ with the largest cardinality. We set
$\mu_{k}:=\frac{1}{\#F_{k}}\sum_{x\in F_{k}}\mathcal{E}_{n_{k}}(x).$
Then $\mu_{k}\in B(K,\delta_{k})$ since $K$ is convex. Now choose a subsequence $\{n_{k_{j}}\}$ such that $\mu_{k_{j}}\to\mu$ as $j\to\infty$. Then $\mu\in K$ since $K$ is a closed. By Lemma 3.1, we have
$\Theta(K,\epsilon)-\gamma\leq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log s_{n_{k_{j}}}(\mathcal{Q}_{k_{j}}(\delta_{k_{j}},n_{k_{j}}),\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
This implies that $h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$. ∎
Thus, by Lemma 3.2 the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K},d)$ is bounded above by
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.2)
3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
The challenging direction is how to estimate the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$. This shall be done by constructing a geometric Moran fractal set that is contained in $G_{K}^{\omega,U}$.
For any sufficiently small $\epsilon^{\ast}>0$, we prove that
$\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$ (3.3)
Then the above inequality yields the desired lower bound.
Without loss of generality, one may assume $H:=\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)>0$. Let $\zeta>0$ and then take $\mu^{\ast}\in K$ such that
$\inf_{\mathrm{diam}(\xi)<20\epsilon^{\ast}}h_{\mu^{\ast}}(f,\xi)>H-\zeta.$ (3.4)
Since $K$ is a connected and compact set, we choose a sequence $\{\mu_{(k,i)}:k\in\mathbb{N},1\leq i\leq Q_{k}\}\subseteq K$ such that
$K\subset\bigcup_{i=1}^{Q_{k}}B(\mu_{(k,i)},\frac{1}{k}),\ D(\mu_{(k,i)},\mu_{(k,i+1)})$ $\leq\frac{1}{k},\ 1\leq i<Q_{k}$
$D(\mu_{(k,Q_{k})},\mu_{(k+1,1)})$ $\leq\frac{1}{k},\ \mu_{(k,Q_{k})}=\mu^{\ast}.$
Now endow the set $\mathcal{A}=\{(0,0)\}\cup\{(k,j):k\geq 1,0\leq j\leq Q_{k}\}$ with the lexicographic order. Denote by $\iota(k,j)$ the position of $(k,j)\in\mathcal{A}$. For $(k,j)\in\mathcal{A}$, we shall use $(k,j)^{\ast}$ and $(k,j)_{\ast}$ to represent the one before and the one after $(k,j)$, respectively.
Let $\{\delta_{(k,j)}\}$ be a strictly decreasing sequence indexed by $\mathcal{A}$ with $\delta_{(0,0)}<\zeta$, $\delta_{(k,j)}\to 0$ as $(k,j)\to\infty$. We shall simply write $\delta_{k}=\delta_{(k,Q_{k})}$. For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$ be chosen as in the definition of the specification property.
For any non-empty open set $U$, there exist $0<\tilde{\epsilon}<\epsilon^{*}$ and a point $a_{0}\in X$ such that ${\overline{B}(a_{0},\tilde{\epsilon})}\subseteq U$. Set
$\mathcal{C}_{(0,0)}=\{a_{0}\},\ n_{(0,0)}=1,\ N_{(0,0)}=1.$
For any $k\in\mathbb{N}$, there exists $\Delta_{k}=\{a_{1}^{k},\ldots,a_{q_{k}}^{k}\}\subset AP(X)$, which is $\epsilon_{(k,0)}$-dense in $X$ by Proposition 2.2. By the specification property, there exists $a_{k}\in X$ such that
$d\left(f^{(i-1)m_{(k,0)}+i-1}(a_{k}),a_{i}^{k}\right)<\epsilon_{(k,0)},\ i=1,2,\ldots,q_{k}.$
Write $\mathcal{C}_{(k,0)}=\{a_{k}\}$, $n_{(k,0)}=q_{k}+(q_{k}-1)m_{(k,0)}$ and $N_{(k,0)}=1$.
Let $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$ and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$. The following measure approximation of $\mu^{*}$ is a slight modification of [LL24, Lemma 2.1], so we omit the proof.
Lemma 3.3.
For each $k\in\mathbb{N}$, there exists $\nu_{k}\in M_{f}(X)$ satisfying
$(1)$ $v_{k}=\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\nu_{i}^{k}$, where $\nu_{i}^{k}\in M_{f}^{e}(X)$, $\lambda_{i}^{k}\in\mathbb{Q}^{+}$ for every $1\leq i\leq j_{k}$, and $\sum_{i=1}^{j_{k}}\lambda_{i}^{k}=1$;
$(2)$ one has
$\inf_{\xi\succ\mathcal{U}}h_{\mu^{\ast}}(f,\xi)\leq\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)+\zeta;$ (3.5)
$(3)$ $D(\mu^{\ast},\nu_{k})<\delta_{k}$, where $D$ is the compatible metric on $M(X)$ given in (2.1).
Since $\nu_{i}^{k}$ is ergodic, by Birkhoff ergodic theorem there exists $N\in\mathbb{N}$ such that the set
$Y_{i}(N)=\left\{x\in X:D(\mathcal{E}_{n}(x),\nu_{i}^{k})<\delta_{k},\ \forall n>N\right\}$
has $\nu_{i}^{k}$-measure greater than $1-\zeta$ for all $1\leq i\leq j_{k}$. Applying Lemma 2.3, one has
$\underline{h}_{\nu_{i}^{k}}^{K}(f,20\epsilon^{\ast},\zeta)\leq\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)\leq\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)<\infty.$
Now take an integer $n_{k}^{\prime}\in\mathbb{N}$ such that
$(1)$ $\lambda_{i}^{k}n_{k}^{\prime}$ is an integer for all $1\leq i\leq j_{k}$, and
$\frac{n_{k}^{\prime}}{n_{k}}>\max\left\{1-\zeta,1-\frac{1}{k}\right\},$ (3.6)
where $n_{k}=n_{(k,Q_{k})}=n_{k}^{\prime}+(j_{k}-1)m_{(k,Q_{k})}$.
$(2)$ there exists a subset $E_{k,i}$ that is a $(\lambda_{i}^{k}n_{k}^{\prime},5\epsilon^{\ast})$-separated set of $Y_{i}(N)$ satisfying
$E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right],$
where the second inequality is derived from Lemma 2.3.
$(3)$ for every $x\in E_{k,i}$,
$D(\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x),\nu_{i}^{k})<\delta_{k}.$
By the specification property, for $k\geq 1$ and each $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})\in E_{k,1}\times E_{k,2}\times\cdots\times E_{k,j_{k}}$, there exists $y=y(\underline{x}_{k})\in X$ such that
$d\left(f^{T_{i}+t_{i}}(y),f^{t_{i}}(x_{i})\right)<\epsilon_{(k,Q_{k})}<\frac{\epsilon^{\ast}}{2^{k+2}},\ 1\leq i\leq j_{k},\ 0\leq t_{i}\leq\lambda_{i}^{k}n_{k}^{\prime}-1,$
where $T_{1}=0$ and $T_{i}=\lambda_{1}^{k}n_{k}^{\prime}+\cdots+\lambda_{i-1}^{k}n_{k}^{\prime}+(i-1)m_{k}$ for $2\leq i\leq j_{k}$. For any two distinct $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})$, $\underline{x}_{k}^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{j_{k}}^{\prime})$ in $E_{k,1}\times\cdots\times E_{k,j_{k}}$, one has $y(\underline{x}_{k})\neq y^{\prime}(\underline{x}_{k}^{\prime})$. Indeed, if $x_{i}\neq x_{i}^{\prime}$ for some $1\leq i\leq j_{k}$, then
$5\epsilon^{\ast}<d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(x_{i},x_{i}^{\prime}\right)\leq d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(x_{i},f^{T_{i}}y\right)+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y,f^{T_{i}}y^{\prime}\right)+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y^{\prime},x_{i}^{\prime}\right)$
$<\frac{\epsilon^{\ast}}{2^{k+1}}+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y,f^{T_{i}}y^{\prime}\right),$
which implies $d_{n_{k}}(y,y^{\prime})>4\epsilon^{\ast}$.
Let
$\mathcal{C}_{(k,Q_{k})}=\left\{y(x_{1},x_{2},\ldots,x_{j_{k}}):\ (x_{1},x_{2},\ldots,x_{j_{k}})\in E_{k,1}\times E_{k,2}\times\cdots\times E_{k,j_{k}}\right\}.$
Then $\mathcal{C}_{(k,Q_{k})}$ is an $(n_{k},4\epsilon^{\ast})$-separated set and
$\#\mathcal{C}_{(k,Q_{k})}$ $=\Pi_{i=1}^{j_{k}}\#E_{k,i}\geq\Pi_{i=1}^{j_{k}}\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right]$
[…]
${\geq}\exp\left[n_{k}(1-\zeta)\left(H-3\zeta\right)\right],$ (3.7)
where the last inequality holds since ${\rm diam}(\xi)<20\epsilon^{\ast}$ and (3.4), (3.6).
For $n\in\mathbb{N}$ and $\varphi\in C(X)$, we write $S_{n}\varphi=\sum_{l=0}^{n-1}\varphi\circ f^{i}$ as the $n$-th Birkhoff sum of $\varphi$. Then for $y\in\mathcal{C}_{(k,Q_{k})}$ and $\varphi\in C(X)$ taking values in $[0,1]$,
$\left|\int\varphi\mathop{}\!\mathrm{d}\mathcal{E}_{n_{k}}(y)-\int\varphi\mathop{}\!\mathrm{d}\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y))\right)\right|$
$\leq$ $\left|\frac{1}{n_{k}}\sum_{i=1}^{j_{k}}S_{\lambda_{i}^{k}n_{k}^{\prime}}\varphi(f^{T_{i}}(y))-\frac{1}{n_{k}^{\prime}}\sum_{i=1}^{j_{k}}S_{\lambda_{i}^{k}n_{k}^{\prime}}\varphi(f^{T_{i}}(y))\right|+\frac{n_{k}-n_{k}^{\prime}}{n_{k}}\|\varphi\|$
$\leq$ $\left[\left|\frac{1}{n_{k}^{\prime}}-\frac{1}{n_{k}}\right|\cdot n_{k}^{\prime}+\frac{n_{k}-n_{k}^{\prime}}{n_{k}}\right]\cdot\|\varphi\|<\frac{2}{k}\|\varphi\|,$
which shows that $D(\mathcal{E}_{n_{k}}(y),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}f^{T_{i}}(y))<\frac{2}{k}$. Since the map $x\mapsto\delta_{x}$ is uniformly continuous, we may assume that for any $x,y\in X$ with $d(x,y)<\frac{\epsilon^{\ast}}{2^{k+2}}$ implies $D(\delta_{x},\delta_{y})<\frac{1}{k}$. Since
$D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i}),\nu_{k}\right)<\delta_{k},\ \ D\left(\nu_{k},\mu^{\ast}\right)<\delta_{k},$
one has that for any $y\in\mathcal{C}_{(k,Q_{k})}$,
$D\left(\mathcal{E}_{n_{k}}(y),\mu^{\ast}\right)\leq\frac{2}{k}+D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i})\right)+2\delta_{k}<\frac{3}{k}+2\delta_{k}.$ (3.8)
Until now, by virtue of the specification property, we have obtained two sets $\mathcal{C}_{(k,0)}$ and $\mathcal{C}_{(k,Q_{k})}$ for each $k\geq 1$. We proceed to consider the cases for $1\leq j<Q_{k}$.
Since the system has the specification property, the generic point $G_{\mu}\neq\emptyset$ holds for all $\mu\in M_{f}(X)$ [PS07]. Using this fact for any $k\geq 1$ and $1\leq j<Q_{k}$, there exist $x_{k,j}\in G_{\mu_{(k,j)}}$ and $n_{(k,j)}\in\mathbb{N}$ such that
$D(\mathcal{E}_{n_{(k,j)}}(x_{k,j}),\mu_{(k,j)})<\delta_{(k,j)},$
and
$\frac{n_{(k,j)}}{n_{(k,j)}+m_{(k,j)}}>1-\frac{1}{k}.$
Now set $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$.
With these data $\{\mathcal{C}_{(k,j)}\}$ and $\{n_{(k,j)}\}$, we are ready to proceed the next construction:
Step $1$. Construct the intermediate sets $\mathcal{S}_{(k,j)}$.
Recall that we have defined $N_{(k,0)}=1$ for any $k\geq 0$. Now choose a strictly increasing sequence of integers $\{N_{(k,j)}\}_{(k,j)\in\mathcal{A},~j\neq 0}$ such that $N_{(1,1)}=1$ and
$\begin{split}\sum_{0\leq k^{\prime}\leq k+1}(n_{(k^{\prime},0)}+m_{(k^{\prime},0)})+n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}&\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})],\\ \sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})]&\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k},\\ n_{k}N_{(k,Q_{k})}&\geq(1-\zeta)\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}].\end{split}$ (3.9)
For $(k,0)\in\mathcal{A}$, we set $\mathcal{S}_{(k,0)}:=\mathcal{C}_{(k,0)}$ and $c_{k,0}:=n_{(k,0)}$. We divide the following two cases to define the rest $\mathcal{S}_{(k,j)}$ indexed by $j\neq 0$.
Case $1$. $1\leq j<Q_{k}$.
Recall $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$. By the specification property again, there exists $y_{k,j}\in X$ such that the pieces of orbits $\{x_{k,j},\cdots,f^{n_{(k,j)-1}}(x_{(k,j)})\}$ repeating $N_{(k,j)}$ times, can be $\epsilon_{(k,j)}$-traced by $y_{(k,j)}$ with the gap $m_{(k,j)}$. Now we set $\mathcal{S}_{(k,j)}=\{y_{k,j}\}$, and $c_{k,j}=n_{(k,j)}N_{(k,j)}+(N_{(k,j)}-1)m_{(k,j)}$.
Case $2$. $j=Q_{k}$.
Enumerate the points in $\mathcal{C}_{(k,Q_{k})}$ as
$\mathcal{C}_{(k,Q_{k})}=\{y_{i}^{k}:i=1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}.$
For any $(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}$, let $y(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in X$ be a point given by the specification property, such that the orbit $\epsilon_{(k,Q_{k})}$-shadows, with the gap $m_{(k,Q_{k})}$, the pieces of orbits $\{y_{i_{l}}^{k},f(y_{i_{l}}^{k}),\ldots,f^{n_{k}-1}(y_{i_{l}}^{k})\}$, $l=1,2,\ldots,N_{(k,Q_{k})}$. We write
$\mathcal{S}_{(k,Q_{k})}=\left\{y=y(i_{1},\ldots,i_{N_{(k,Q_{k})}}):(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}\right\}$
and $c_{k,Q_{k}}=n_{k}\cdot N_{(k,Q_{k})}+(N_{(k,Q_{k})}-1)m_{(k,Q_{k})}$. Since $\mathcal{C}_{(k,Q_{k})}$ is an $(n_{k},4\epsilon^{\ast})$-separated set, this implies that $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set. Then, by (3.1.2) and (3.9), the cardinality of $\mathcal{S}_{(k,Q_{k})}$ is bounded blew by
$\mathcal{S}_{(k,Q_{k})}=\#\mathcal{C}_{(k,Q_{k})}^{N_{(k,Q_{k})}}\geq\exp\left[\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}]\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$ (3.10)
Similar to the arguments in (3.8), one has
[…]
and
$D(\mathcal{E}_{c_{k,Q_{k}}}(y),\mu^{\ast})<\frac{5}{k}+2\delta_{k},\ \forall\ y\in\mathcal{S}_{(k,Q_{k})}.$
Hence, noting that $\mu_{(k,Q_{k})}=\mu^{*}$, in an unified manner we always have
[…]
Step $2$. Construct the $\mathbb{F}_{(k,j)}$ and the Moran set $\mathbb{F}$.
This step can be regarded as an analogue of the construction of the middle-third Cantor set of $[0,1]$.
Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$. Suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$, and then we will describe how to construct $\mathcal{I}_{(k,j)}$. Let
$M_{(k,j)}$ $=M_{(k,j)^{\ast}}+m_{(k,j)}+c_{k,j};$
$M_{(k,j)}^{i}$ $=M_{(k,j)^{\ast}}+i(m_{(k,j)}+n_{(k,j)}),\ 0\leq i\leq N_{(k,j)}.$
Note that $M_{(k,j)}^{N_{(k,j)}}=M_{(k,j)}$ and the number of $\mathcal{S}_{(k,Q_{k})}$ given in (3.10) has the estimation:
$\mathcal{S}_{(k,Q_{k})}\geq\exp\left[(M_{(k,Q_{k})}-M_{(k-1,Q_{k-1})})(1-\zeta)^{2}(H-3\zeta)\right].$ (3.12)
For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, by the specification property let $z=z(x,y)\in X$ be some point such that
$d_{M_{(k,j)^{\ast}}}(x,z)<\epsilon_{(k,j)},\ d_{c_{k,j}}(y,f^{M_{(k,j)^{\ast}}+m_{(k,j)}}(z))<\epsilon_{(k,j)}.$
We set $\mathcal{I}_{(k,j)}=\{z(x,y):x\in\mathcal{I}_{(k,j)^{\ast}},y\in\mathcal{S}_{(k,j)}\}$. For $x\in\mathcal{I}_{(k,Q_{k})^{\ast}}$ and two distinct $y,y^{\prime}\in\mathcal{S}_{(k,Q_{k})}$, since $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set, one has
$d_{M_{(k,Q_{k})}}(z(x,y),z(x,y^{\prime}))>2\epsilon^{\ast},$
which leads to
$\#\mathcal{I}_{(k,j)}=\#\mathcal{I}_{(k,j)^{\ast}}\cdot\#\mathcal{S}_{(k,j)}=\Pi_{(k^{\prime},j^{\prime})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},j^{\prime})}=\Pi_{(k^{\prime},Q_{k^{\prime}})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}.$ (3.13)
Moreover, according to (3.11), for $z\in\mathcal{I}_{(k,j)}$ with $j\neq 0$, we have
$D(\mathcal{E}_{c_{k,j}}(f^{M_{(k,j)^{\ast}}}(z)),\mu_{(k,j)})<\frac{6}{k}+2\delta_{(k,j)}.$ (3.14)
Now set
$\mathbb{F}_{(k,j)}=\bigcup_{x\in\mathcal{I}_{(k,j)}}\overline{B}_{M_{(k,j)}}(x,2\epsilon_{(k,j)}).$
Then for each $(k,j)\in\mathcal{A}$, the following statements are satisfied:
$(1)$ $\overline{B}_{M_{(k,j)}}(x,2\epsilon_{(k,j)})\cap\overline{B}_{M_{(k,j)}}(x^{\prime},2\epsilon_{(k,j)})=\emptyset$ holds for any distinct $x,x^{\prime}\in\mathcal{I}_{(k,j)}$;
$(2)$ if $z\in\mathcal{I}_{(k,j)}$ descends from $x\in\mathcal{I}_{(k,j)^{\ast}}$, one has
$\overline{B}_{M_{(k,j)}}(z,2\epsilon_{(k,j)})\subset\overline{B}_{M_{(k,j)^{\ast}}}(x,2\epsilon_{(k,j)^{\ast}}).$
This implies that $\{\mathbb{F}_{(k,j)}:(k,j)\in\mathcal{A}\}$ is a sequence of decreasing non-empty closed subsets of $X$. Let
$\mathbb{F}=\bigcap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)}=\bigcap_{k\in\mathbb{N}}\mathbb{F}_{(k,Q_{k})}.$
Lemma 3.4.
The fractal set $\mathbb{F}$ is contained in $G_{K}^{\omega,U}$, i.e.,
$\mathbb{F}\subset G_{K}^{\omega,U}=G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\}.$
Proof of Lemma 3.4.
By the construction of $\mathbb{F}$, we know
$\mathbb{F}\subset\mathbb{F}_{(0,0)}=\overline{B}_{n(0,0)}(a_{0},\frac{\epsilon^{\ast}}{8})\subset{\overline{B}(a_{0},\tilde{\epsilon})}\subset U.$
For $p\in\mathbb{F}$, we have $p\in\mathbb{F}_{(k,0)}$ for any $k\geq 1$. Hence, for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there exists $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<4\epsilon_{(k,0)}$. This implies that the orbit of $p$ is $5\epsilon_{(k,0)}$-dense in $X$. By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have
$\mathbb{F}\subset\{x\in X:\omega_{f}(x)=X\}.$
Let $p\in\mathbb{F}$. For sufficiently large $n\in\mathbb{N}$, to estimate $D(\mathcal{E}_{n}(p),K)$ we split into four cases to discuss.
Case $(1)$: There exists $k>1$ such that $M_{(k+1,0)}\leq n<M_{(k+1,1)}^{1}$.
Then there is some $z\in\mathcal{I}_{(k,Q_{k})}$ such that $p\in\overline{B}_{M_{(k,Q_{k})}}(z,\frac{\epsilon_{(k,Q_{k})}}{2})$, and
$D(\mathcal{E}_{n}(p),\mu^{\ast})\leq$ $\frac{M_{(k,Q_{k})^{\ast}}}{n}D(\mathcal{E}_{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})+\frac{n-M_{(k,Q_{k})}}{n}D(\mathcal{E}_{n-M_{(k,Q_{k})}}(f^{M_{(k,Q_{k})}}(p)),\mu^{\ast})$
$+\frac{c_{k,Q_{k}}+m_{(k,Q_{k})}}{n}D(\mathcal{E}_{c_{k,Q_{k}}+m_{(k,Q_{k})}}(f^{M_{(k,Q_{k})^{\ast}}}(p)),\mu^{\ast}).$
Noticing that both $\frac{M_{(k,Q_{k})^{\ast}}}{n}$ and $\frac{n-M_{(k,Q_{k})}}{n}$ are smaller than $\frac{1}{k}$, and using formula (3.14), we have $D(\mathcal{E}_{c_{k,Q_{k}}+m_{(k,Q_{k})}}f^{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})<\frac{7}{k}+2\delta_{(k,Q_{k})}$. Thus,
$D(\mathcal{E}_{n}(p),K)\leq\frac{9}{k}+2\delta_{(k,Q_{k})}.$
Case $(2)$: There exists $(k,j)\in\mathcal{A}$ with $j\neq 0$ such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
Then there is some $z\in\mathcal{I}_{(k,j)}$ such that $p\in\overline{B}_{M_{(k,j)}}(z,\frac{\epsilon_{(k,j)}}{2})$, and
$D(\mathcal{E}_{n}(p),\mu_{(k,j)})\leq$ $\frac{M_{(k,j)^{\ast}}}{n}D(\mathcal{E}_{M_{(k,j)^{\ast}}}(p),\mu_{(k,j)})+\frac{n-M_{(k,j)}}{n}D(\mathcal{E}_{n-M_{(k,j)}}(f^{M_{(k,j)}}(p)),\mu_{(k,j)})$
$+\frac{c_{k,j}+m_{(k,j)}}{n}D(\mathcal{E}_{c_{k,j}+m_{(k,j)}}(f^{M_{(k,j)^{\ast}}}(p)),\mu_{(k,j)})$
The similar argument as in Case $(1)$ yields that
$D(\mathcal{E}_{n}(p),K)\leq\frac{9}{k}+2\delta_{(k,j)}.$
Case $(3)$: There exist $(k,1)\in\mathcal{A}$ and $i>1$ such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
One has
$D($ $\mathcal{E}_{n}(p),\mu_{(k,1)})\leq\frac{M_{(k-1,Q_{k-1})^{\ast}}}{n}D(\mathcal{E}_{M_{(k-1,Q_{k-1})^{\ast}}}(p),\mu_{(k,1)})$
$+\frac{c_{k-1,Q_{k-1}}+m_{(k-1,Q_{k-1})}}{n}D(\mathcal{E}_{c_{k-1,Q_{k-1}}+m_{(k-1,Q_{k-1})}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})$
$+\frac{c_{k,0}+m_{(k,0)}}{n}D(\mathcal{E}_{c_{k,0}+m_{(k,0)}}(f^{M_{(k-1,Q_{k-1})}}(p)),\mu_{(k,1)})$
$+\frac{n_{(k,1)}+m_{(k,1)}}{n}\sum_{l=0}^{i-1}D(\mathcal{E}_{n_{(k,1)}+m_{(k,1)}}(f^{M_{(k,1)}^{l}}(p)),\mu_{(k,1)})$
$+\frac{n-M_{(k,1)}^{i}}{n}D(\mathcal{E}_{n-M_{(k,1)}^{i}}(f^{M_{(k,1)}^{i}}(p)),\mu_{(k,1)}).$
Note that
$\frac{M_{(k-1,Q_{k-1})^{\ast}}}{n}<\frac{1}{k-1},\ \frac{c_{k,0}+m_{(k,0)}}{n}<\frac{1}{k},\ \frac{n-M_{(k,1)}^{i}}{n}<\frac{1}{k},$
$D(\mathcal{E}_{n_{(k,1)}+m_{(k,1)}}(f^{M_{(k,1)}^{l}}(p)),\mu_{(k,1)})\leq\frac{5}{k}+2\delta_{(k,1)},$
and $D(\mathcal{E}_{c_{k-1,Q_{k-1}}+m_{(k-1,Q_{k-1})}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})$ is smaller than
$D(\mathcal{E}_{c_{k-1,Q_{k-1}}+m_{(k-1,Q_{k-1})}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu^{\ast})+D(\mu^{\ast},\mu_{(k,1)})\leq\frac{8}{k-1}+2\delta_{(k-1,Q_{k-1})}.$
This gives us
$D(\mathcal{E}_{n}(p),K)\leq\frac{16}{k-1}+4\delta_{(k-1,0)}.$
Case $(4)$: There exist $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$ such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
In this case, using the expression
$\mathcal{E}_{n}(z)$ $=\frac{M_{(k,j-1)^{\ast}}}{n}\mathcal{E}_{M_{(k,j-1)^{\ast}}}(p)+\frac{c_{k,j-1}+m_{(k,j-1)}}{n}\mathcal{E}_{c_{k,j-1}+m_{(k,j-1)}}(f^{M_{(k,j-1)^{\ast}}}(p))$
$+\frac{n_{(k,j)}+m_{(k,j)}}{n}\sum_{l=0}^{i-1}\mathcal{E}_{n_{(k,j)}+m_{(k,j)}}(f^{M_{(k,j)}^{l}}(p))+\frac{n-M_{(k,1)}^{i}}{n}\mathcal{E}_{n-M_{(k,1)}^{i}}(f^{M_{(k,1)}^{i}}(p)),$
and repeating the similar argument as above, one can obtain
$D(\mathcal{E}_{n}(p),K)\leq\frac{15}{k}+4\delta_{(k,j)}.$
To sum up, whatever case we consider, by letting $n\to\infty$ we get $V_{f}(p)\subset K$. Based on the above arguments, one can also show that the sequences $\{\mathcal{E}_{n}(p)\}_{n\in\mathbb{N}}$ and $\{\mu_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ have the same limit points. Notice that $K=\overline{\{\mu_{(k^{\prime},j^{\prime})}:(k^{\prime},j^{\prime})>(k,j)\}}$ for any fixed $(k,j)\in\mathcal{A}$. Consequently, we have $V_{f}(p)=K$, and hence $\mathbb{F}\subset G_{K}$. ∎
Step $3$. Construct the measure $\rho$ supported on $\mathbb{F}$.
Now we define a probability measure $\rho$ on $\mathbb{F}$ satisfying the $\rho$-measure of some appropriate dynamical balls decays exponentially fast along the sequence $\{M_{k}\}$, where $M_{k}:=M_{(k,Q_{k})}$. Then we can invoke the Entropy Distribution Principle stated in Lemma 2.2 to obtain a lower bound of $\epsilon$-packing topological entropy of $\mathbb{F}$.
Define
$\rho_{k}=\frac{1}{\#\mathcal{I}_{(k,Q_{k})}}\sum_{x\in\mathcal{I}_{(k,Q_{k})}}\delta_{x}.$
For simplicity, we write $\mathcal{I}_{k}$, $\mathbb{F}_{k}$ instead of $\mathcal{I}_{(k,Q_{k})}$, $\mathbb{F}_{(k,Q_{k})}$ respectively. Since $M(X)$ is compact, we may assume that $\rho_{k}\to\rho\in M(X)$ as $k\to\infty$. It is clear that $\rho_{k}(\mathbb{F}_{k})=1$ for every $k$, and for any $k>k_{0}$, one has $\rho_{k}(\mathbb{F}_{k_{0}})=1$ since $\mathbb{F}_{k}\subset\mathbb{F}_{k_{0}}$. Thus, $\rho(\mathbb{F}_{k_{0}})\geq\limsup_{k\to\infty}\rho_{k}(\mathbb{F}_{k_{0}})=1$. This implies that $\rho(\mathbb{F})=1$. Recall that $M_{(0,0)}=n_{(0,0)}=1$. By (3.12) and (3.13), we have
$\mathcal{I}_{k}\geq\exp\left[(M_{k}-1)\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
Lemma 3.5.
Let $k_{0}\geq 1$, $z\in X$ and $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ be a Bowen ball that intersects $\mathbb{F}$. Then, for any $k>k_{0}$, one has
$\rho_{k}\left(B_{M_{k_{0}}}\left(z,\frac{\epsilon^{\ast}}{2}\right)\right)\leq\frac{1}{\#\mathcal{I}_{k_{0}}}.$
Proof of Lemma 3.5.
Since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}\not$$\emptyset$, the set $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}_{k_{0}}$ is also non-empty by the construction of $\mathbb{F}$ and $(\mathbb{F}_{k_{0}}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))=1$. Notice that $\mathbb{F}_{k_{0}+1}$ descends from $\mathbb{F}_{k_{0}}$. Then one has
$(\mathbb{F}_{k_{0}+1}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))\leq\#\mathcal{S}_{(k_{0}+1,Q_{k_{0}+1})},$
and hence
$\rho_{k_{0}+1}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{\#\mathcal{S}_{(k_{0}+1,Q_{k_{0}+1})}}{\#\mathcal{I}_{k_{0}+1}}=\frac{1}{\#\mathcal{I}_{k_{0}}}.$
In general, for any $k>k_{0}$, one has
$\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{\Pi_{k_{0}+1\leq k^{\prime}\leq k}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}}{\#\mathcal{I}_{k}}=\frac{1}{\#\mathcal{I}_{k_{0}}}.$
By Lemma 3.5, for sufficiently large $k_{0}$ and any Bowen ball $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ with $z\in\mathbb{F}$, since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ is open, one has
$\rho\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)$ $\leq\liminf_{k\to\infty}\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{1}{\#\mathcal{I}_{k_{0}}}$
$\leq\exp\left[-(M_{k_{0}}-1)(1-\zeta)^{2}\cdot(H-3\zeta)\right]$
$\leq\exp\left[-M_{k_{0}}(1-\zeta)^{3}\cdot(H-3\zeta)\right]$
Using Lemma 2.2, one has
$(1-\zeta)^{3}\cdot(H-3\zeta)\leq h_{top}^{P}(f,\mathbb{F},\frac{\epsilon^{\ast}}{12})\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$
The arbitrariness of $\zeta$ implies that
$\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$
Dividing in both sides of the above inequality by $|\log\epsilon^{\ast}|$, and then taking $\limsup_{\epsilon^{\ast}\to 0}$, we have
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d).$ (3.15)
Finally, by (3.2) and (3.15) we get
$\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
which proves Theorem 1.1.
3.2 Bowen metric mean dimension of saturated sets
In this subsection, we prove Theorems 1.2 and 1.3.
3.2.1 Upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$
The following lemma allows us to get an upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$.
Lemma 3.6.
[Y23, Lemma 3.3,(3)] Let $(X,d,f)$ be a TDS, and let $K\subset M_{f}(X)$ be a non-empty closed subset.
  • (1)
    For any $\epsilon>0$, we have
    $h_{top}^{B}(f,\{x\in X:V_{f}(x)\cap K\not=\emptyset\},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
  • (2)
    For any $\epsilon>0$, we have
    $h_{top}^{B}(f,G_{K},\epsilon)\leq\inf_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Thus, by Lemma 3.6 we have
$\overline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.16)
3.2.2 Lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
The proof for the lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$ is similar to that in subsection 3.1.2. So we only give a sketch for the proof.
For any sufficiently small $\epsilon^{\ast}>0$, we show
$\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2}).$ (3.17)
Fix $\zeta>0$ and pick a sequence $\{\mu_{(k,i)}:k\geq 1,1\leq i\leq Q_{k}\}\subset K$ of measures in $K$ such that
$K\subset\cup_{i=1}^{Q_{k}}B(\mu_{(k,i)},\frac{1}{k}),\ D(\mu_{(k,i)},\mu_{(k,i+1)})<\frac{1}{k},\ D(\mu_{(k,Q_{k})},\mu_{(k+1,1)})<\frac{1}{k}.$
Put $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$, $\mathcal{A}=\{(k,0):k\geq 0\}\cup\mathcal{A}^{\prime}$, and endow these sets with the lexicographic order again. Using the fixed notation in subsection 3.1.2, we have the data $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ and $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$.
Assume that $H:=\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)>0$. Then one has
$\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu_{(k,i)}}(f,\xi)>H-\zeta,\ \forall(k,i)\in\mathcal{A}^{\prime}.$
Let $\mathcal{U}$ be an open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$, and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$. For $(k,i)\in\mathcal{A}^{\prime}$, by Lemma 3.3 let $\nu_{(k,i)}=\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\nu_{l}^{k,i}$ be the finite convex combination with rational coefficients of $f$-ergodic probability measures such that
$D(\mu_{(k,i)},\nu_{(k,i)})<\delta_{(k,i)},\ \ \inf_{\xi\succ\mathcal{U}}h_{\mu_{(k,i)}}(f,\xi)\leq\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)+\zeta.$
Repeating the previous argument as in subsection 3.1.2, we choose an integer $n_{(k,i)}^{\prime}$ such that the following conditions are satisfied:
$(1)$ for any $1\leq l\leq p_{k,i}$, $\lambda_{k,i,l}n_{(k,i)}^{\prime}$ is an integer and
$\frac{n_{(k,i)}^{\prime}}{n_{(k,i)}+m_{(k,i)}}>\max\{1-\zeta,1-\frac{1}{k}\},$
where $n_{(k,i)}:=n_{(k,i)}^{\prime}+(p_{k,i}-1)m_{(k,i)}$;
$(2)$ there exist subsets $E_{k,i,l}$, $1\leq l\leq p_{k,i}$, which are $(\lambda_{k,i,l}n_{(k,i)}^{\prime},5\epsilon^{\ast})$-separated satisfying
$E_{k,i,l}\geq\exp\left[\lambda_{k,i,l}n_{(k,i)}^{\prime}\cdot\left(\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)-\zeta\right)\right];$
$(3)$ for any $x\in E_{k,i,l}$,
$D(\mathcal{E}_{\lambda_{k,i,l}n_{(k,i)}^{\prime}}(x),\nu_{l}^{k,i})<\delta_{(k,i)}.$
Let
$\mathcal{C}_{(k,i)}=\left\{y=y(x_{1},\ldots,x_{p_{k,i}}):(x_{1},\ldots,x_{p_{k,i}})\in E_{k,i,1}\times\cdots\times E_{k,i,p_{k,i}}\right\},$
where $y(x_{1},\ldots,x_{p_{k,i}})$ is the point that $\epsilon_{(k,i)}$-shadows the pieces of orbits
$\{x_{l},\ldots,f^{\lambda_{k,i,l}n_{(k,i)}^{\prime}}(x_{l})\},\ 1\leq l\leq p_{k,i},$
with gap $m_{(k,i)}$. Then $\mathcal{C}_{(k,i)}$ is a $(n_{(k,i)},4\epsilon^{\ast})$-separated set with
$\mathcal{C}_{(k,i)}\geq\exp\left[n_{(k,i)}(1-\zeta)\cdot(H-2\zeta)\right].$
For $k\geq 0$, let $\mathcal{C}_{(k,0)}$, $n_{(k,0)}$, $N_{(k,0)}$ be defined as in subsection 3.1.2. Choose a strictly increasing sequence $\{N_{(k,i)}\}_{(k,i)\in\mathcal{A}^{\prime}}$ such that $N_{(1,1)}=1$ and
$\begin{split}\sum_{0\leq k^{\prime}\leq k+1}(n_{(k^{\prime},0)}+m_{(k^{\prime},0)})+n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}&\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})],\\ \sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})]&\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k}.\end{split}$
Set $\mathcal{S}_{(k,0)}=\mathcal{C}_{(k,0)}$, $c_{k,0}=n_{(k,0)}$ for $k\geq 0$. For every $(k,i)\in\mathcal{A}^{\prime}$ and any tuple of points $(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})\in\mathcal{C}_{(k,i)}^{N_{(k,i)}}$, there exists a point $z=z(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})$ that $\epsilon_{(k,i)}$-shadows the pieces of orbits
$\{y_{k,i,l},f(y_{k,i,l}),\ldots,f^{n_{(k,i)}-1}(y_{k,i,l})\},\ l=1,\cdots,N_{(k,i)},$
with the same gap $m_{(k,i)}$. Denote by $\mathcal{S}_{(k,i)}$ the set of such shadowing points. It is a $(c_{k,i},3\epsilon^{\ast})$-separated set, where $c_{k,i}=n_{(k,i)}N_{(k,i)}+(N_{(k,i)}-1)m_{(k,i)}$.
Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$. Suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$. Let $M_{(k,j)}=M_{(k,j)^{\ast}}+m_{(k,j)}+c_{k,j}$. For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, let $z=z(x,y)$ be the shadowing point such that $d_{M_{(k,j)^{\ast}}}(z,x)<\epsilon_{(k,j)}$ and
$d_{c_{k,j}}(y,f^{M_{(k,j)^{\ast}}+m_{(k,j)}}(z))<\epsilon_{(k,j)}.$
We set $\mathcal{I}_{(k,j)}=\{z=z(x,y):x\in\mathcal{I}_{(k,j)^{\ast}},y\in\mathcal{S}_{(k,j)}\}$, and then define
$\mathbb{F}=\cap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)},$
where $\mathbb{F}_{(k,j)}=\cup_{z\in\mathcal{I}_{(k,j)}}\overline{B}_{M_{(k,j)}}(z,2\epsilon_{(k,j)})$. One can verify that $\mathbb{F}\subset G_{K}^{\omega,U}$.
We proceed to define a sequence of measures $\{\rho_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ by
$\rho_{(k,j)}:=\frac{1}{\#\mathcal{I}_{(k,j)}}\sum_{x\in\mathcal{I}_{(k,j)}}\delta_{x}.$
Without loss of generality, we can assume that $\rho_{(k,j)}\to\rho$. Furthermore, one can check that $\rho(\mathbb{F})=1$, and for every sufficiently large $n$ and $B_{n}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}$, it holds that
$\rho(B_{n}(z,\frac{\epsilon^{\ast}}{2}))\leq\exp\left[-n(1-\zeta)^{2}(H-2\zeta)\right].$
Then, by the Lemma 2.1 we have
$(1-\zeta)^{2}(H-2\zeta)\leq h_{top}^{B}(f,\mathbb{F},\frac{\epsilon^{\ast}}{2})\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2}).$
Letting $\zeta\to 0$ gives us
$\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2})$
Therefore, by (3.16) and (3.17) we get
$\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{B}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{\mathrm{diam}\xi<\epsilon}h_{\mu}(f,\xi).$
3.2.3 Proof of Theorem 1.3
The more satisfactory results would be to exchange the order of $\limsup_{\epsilon\to 0}$ and $\sup_{\mu\in K}$ (or $\inf_{\mu\in K}$) of Theorems 1.1 and 1.2. However, up to now, the authors remain unaware of whether there exist any counter-examples among infinite entropy systems with the specification property for which the following equalities hold strictly, i.e.,
$\sup_{\mu\in K}\left\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\right\}<\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d),$
$\inf_{\mu\in K}\left\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\right\}<\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d).$
Inspired by the definition of mean Rényi information dimension [GS21], we define two variants of mean Rényi information dimension of invariant measures.
Let $K$ be a closed set of $M_{f}(X)$. For any $\mu\in K$, we use $M_{K}(\mu)$ to denote the space of sequences of invariant measures within $K$ which converge to $\mu$ in the weak-topology.
Definition 3.1.
For any $\mu\in K$, we define the Bowen and packing Rényi information dimensions of $\mu$ as
$\underline{\rm mrid}_{\mu}^{B}(f,K,d):$ $=\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\},$
$\overline{\rm mrid}_{\mu}^{P}(f,K,d):$ $=\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\},$
respectively.
With the help of Bowen and packing Rényi information dimensions, we establish the variational principles for Bowen and packing metric mean dimensions of saturated sets, whose form are close to the ones of Bowen and packing topological entropies of saturated sets [PS07, ZCC12].
Now we are ready to prove Theorem 1.3. We first prove the variational principle:
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\underline{\rm mrid}_{\mu}^{B}(f,K,d).$
Fix $\mu\in K$. For every $(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)$, by Lemma 3.6 it holds that
$h_{top}^{B}(f,G_{K},\epsilon)\leq\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
This shows that
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)$
and hence $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\inf_{\mu\in K}\limits\underline{\rm mrid}_{\mu}^{B}(f,K,d)$.
To show the reverse inequality
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\geq\inf_{\mu\in K}\limits\underline{\rm mrid}_{\mu}^{B}(f,K,d),$
we assume that $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)$ is finite; otherwise there is nothing left to prove. Let $\gamma>0$. By Theorem 1.2, we can choose a strictly decreasing sequence $\epsilon_{k}\to 0$ as $k\to 0$ and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for any $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi)<\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$ (3.18)
Without loss of generality, we assume that $\mu_{k}\rightarrow\mu$ as $k\to\infty$. Then $\mu\in K$ since $K$ is closed. For every $\epsilon>0$, if $\epsilon\in(\epsilon_{k+1},\epsilon_{k}]$ for some $k\geq 1$, we define $\nu_{\epsilon}:=\epsilon_{k}$. Then $(\nu_{\epsilon})\in M_{K}(\mu)$. Using (3.18), we have
$\underline{\rm mrid}_{\mu}^{B}(f,K,d)\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
Letting $\gamma\to 0$, we have
$\inf_{\mu\in K}\underline{\rm mrid}_{\mu}^{B}(f,K,d)\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d).$
Next we show the variational principle:
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\limits\overline{\rm mrid}_{\mu}^{P}(f,K,d).$
Using (3.3), there exists a constant $c>0$ for every sufficiently small $\epsilon>0$,
$\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{P}(f,G_{K},c\epsilon).$
This yields the inequality
$\sup_{\mu\in K}\overline{\rm mrid}_{\mu}^{P}(f,K,d)\leq\overline{\rm mdim}_{M}^{P}(f,G_{K},d).$
On the other hand, let $\gamma>0$ and put $M_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{P}(f,G_{K},d)-\gamma\}.$ By Theorem 1.1, we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real numbers that converges to 0 and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for sufficiently large $k$,
$M_{\gamma}<\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi).$
By the similar arguments, we have
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)\leq\sup_{\mu\in K}\overline{\rm mrid}_{\mu}^{P}(f,K,d).$
3.3 Upper capacity metric mean dimension of saturated sets
In this subsection, we give the proof of Theorem 1.4.
By Proposition 2.1, it is clear that
$\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d)\leq\overline{\rm mdim}_{M}^{UC}(f,G_{K},d)\leq\overline{\rm mdim}_{M}(f,X,d)$
is clear. Thus, it suffices to show
$\overline{\rm mdim}_{M}(f,X,d)\leq\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d).$
In light of Theorem 2.4, we need to show that for every sufficiently small $\epsilon^{\ast}>0$, $\zeta>0$ and $\delta\in(0,1)$, one has
$\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta\leq h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast}).$ (3.19)
Here, we will give a sketch for the proof of (3.19) since it is essentially analogous to that of the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$ presented in subsection 3.1.2.
Write $H=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)$ and assume that $H>0$. Take $\mu^{\ast}\in M_{f}^{e}(X)$ such that
$\underline{h}_{\mu^{\ast}}^{K}(f,3\epsilon^{\ast},\delta)>H-\zeta.$
Then, for sufficiently large $\mathcal{N}\in\mathbb{N}$, there exists a $(\mathcal{N},3\epsilon^{\ast})$-separated subset $S_{\mathcal{N}}$ with
$S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)].$
Since $K$ is a connected closed subset, a standard argument shows that there exists a family $\{\mu_{k,i}:k\geq 1,1\leq i\leq Q_{k}\}$ of measures in $K$ such that for all $k\geq 1$, $1\leq i<Q_{k}$,
$K\subset\bigcup_{i=1}^{Q_{k}}B(\mu_{k,i},\frac{1}{k}),\ D(\mu_{k,i},\mu_{k,i+1})<\frac{1}{k},\ \text{and }D(\mu_{k,Q_{k}},\mu_{k+1,1})<\frac{1}{k}.$
Consider the set $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$ and endow $\mathcal{A}=\{(k,0):k\in\mathbb{Z}\ \text{and }k\geq-1\}\cup\mathcal{A}^{\prime}$ with the lexicographic order. Similarly, the sequences $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ are constructed as in subsection 3.1.2.
For the non-empty open set $U$, there exist $a_{0}\in X$ and $\epsilon_{0}\in(0,\epsilon^{\ast})$ with ${\overline{B}(a_{0},\epsilon_{0})}\subset U$. Set $\mathcal{C}_{(-1,0)}=\{a_{0}\}$, $n_{(-1,0)}=c_{(-1,0)}=N_{(-1,0)}=1$, and $\mathcal{C}_{(0,0)}=S_{\mathcal{N}}$, $c_{(0,0)}=\mathcal{N}$, $N_{(0,0)}=1$. For $k\geq 1$, let $\mathcal{C}_{(k,0)}$, $n_{(k,0)}$, $c_{(k,0)}$ and $N_{(k,0)}$ be the same as in subsection 3.1.2. For $(k,i)\in\mathcal{A}^{\prime}$, since $G_{\mu_{k,i}}\neq\emptyset$, one can pick $x_{k,i}\in X$ and $n_{(k,i)}\in\mathbb{N}$ such that
$D(\mathcal{E}_{n_{(k,i)}}(x),\mu_{k,i})<\delta_{(k,i)},\ \frac{n_{(k,i)}}{n_{(k,i)}+m_{(k,i)}}>\max\left\{1-\zeta,1-\frac{1}{k}\right\}.$
Set $\mathcal{C}_{(k,i)}=\{x_{k,i}\}$. Now choose a strictly increasing sequence of integers $\{N_{(i,j)}\}_{(k,j)\in\mathcal{A}\ \text{with }j\neq 0}$ with $N_{(1,1)}=1$ and
$\begin{split}&n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}\left[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})\right],\\ &\sum_{(k^{\prime},j^{\prime})\leq(k,j)}\left[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})\right]\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k}.\end{split}$
Based on the data $\mathcal{C}_{(k,j)}$, $N_{(k,j)}$ for $(k,j)\in\mathcal{A}$, similar to the previous subsections, one can construct the intermediate sets $\mathcal{S}_{(k,j)}$, $\mathcal{I}_{(k,j)}$, $\mathbb{F}_{(k,j)}$ to obtain a fractal set
$\mathbb{F}=\bigcap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)}\subset G_{K}^{\omega,U}.$
It is worth mentioning that for any $(k,j)>(0,0)$, $\mathcal{I}_{(k,j)}=\#S_{\mathcal{N}}$. Moreover, for any distinct $x$, $x^{\prime}\in\mathcal{I}_{(0,0)}$, one has
$d_{M_{(0,0)}}(x,x^{\prime})>2\epsilon^{\ast}.$
Recall $M_{(0,0)}=1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}$. Let $y$, $y^{\prime}\in\mathbb{F}$ with $y\in\overline{B}_{M_{(0,0)}}(x,2\epsilon_{(0,0)})$ and $y^{\prime}\in\overline{B}_{M_{(0,0)}}(x^{\prime},2\epsilon_{(0,0)})$. Then one has
$d_{M_{(0,0)}}(y,y^{\prime})>d_{M_{(0,0)}}(x,x^{\prime})-d_{M_{(0,0)}}(x,y)-d_{M_{(0,0)}}(x^{\prime},y^{\prime})>\epsilon^{\ast}.$
Hence, $\mathbb{F}$ is a $(1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N},\epsilon^{\ast})$-separated subset with $S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)]$. Thus,
$h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast})\geq\limsup_{\mathcal{N}\to\infty}\frac{\mathcal{N}(H-2\zeta)}{1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}}=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta.$
This implies the inequality (3.19) by letting $\zeta\to 0$.
3.4 Metric mean dimensions of sup sets and sub sets
In this section, we consider two variants of saturated sets, namely sup sets and sub sets, and establish variational principles for metric mean dimensions of sup sets and sub sets.
In general, given a point $x\in X$, the limit $\lim_{n\to\infty}\mathcal{E}_{n}(x)$ may not exist in the weak-topology. The set of such points is denoted by $I(f)$, known as the historic set [^1] This terminology is due to Rulle’s original paper [Rue01]. of $f$. To understand the underlying geometry structure of $I(f)$, we define the following sets based on the framework introduced and further developed by Olsen [O02, O03, O06, O08] and Olsen and Winter [OW03]. More precisely, given a non-empty subset $C\subset M_{f}(X)$, we define the sup set and sub set of $C$ as
$^{C}G:$ $=\{x\in X:V_{f}(x)\supset C\},$
$G^{C}:$ $=\{x\in X:V_{f}(x)\subset C\},$
respectively. More generally, for two non-empty subsets $C_{1},C_{2}\subset M_{f}(X)$, we can define
$G(C_{1},C_{2}):=\{x\in X:C_{1}\subset V_{f}(x)\subset C_{2}\},$
to recover the above notions in an unified manner.
Now, we are ready to establish variational principles for $G^{C},^{C}G$.
Theorem 3.7.
Let $(X,d,f)$ be a TDS with the specification property. Then for any non-empty subset $C\subseteq M_{f}(X)$,
$\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
We claim that for every $\epsilon>0$,
$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)=\inf_{\mu\in\overline{{\rm co}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$ (3.20)
where $\overline{{\rm co}}(C)$ denotes the closure of the convex hull of $C$.
Indeed, since $C\subset\overline{{\rm co}}(C)$, it suffices to show
$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\inf_{\mu\in\overline{{\rm co}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Let $\mu\in\overline{{\rm co}}(C)$. Choose $\mu_{n}=\sum_{j=1}^{k_{n}}\lambda_{n}^{(j)}v_{n}^{(j)}$ with $v_{n}^{(j)}\in C$, $0<\lambda_{n}^{(j)}<1$ and $\sum_{j=0}^{k_{n}}\lambda_{n}^{(j)}=1$ such that $\mu_{n}\rightarrow\mu$ as $n\to\infty$. The upper semi-continuity of $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ on $M_{f}(X)$ (cf. [FW16, Lemma 2.3, (3)]) implies that
$\limsup_{n\to\infty}~{\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu_{n}}(f,\xi)}\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Since $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ is a concave function on $M_{f}(X)$, for every $\gamma>0$ we can deduce that
$\inf_{\nu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\nu}(f,\xi)<\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)+\gamma.$
This yields the desired inequality (3.20).
Notice that $G_{\overline{{\rm co}}(C)}\subset^{C}G$, we have
$\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm co}}(C)},d\right)$
[…]
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Fix $\mu\in C$ and $\epsilon>0$. By Lemma 3.6, one has
$h_{top}^{B}(f,^{C}G,\epsilon)\leq h_{top}^{B}(f,\{x\in X:\mu\in V_{f}(x)\},\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
and hence $h_{top}^{B}(f,^{C}G,\epsilon)\leq\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$. ∎
Theorem 3.8.
Let $(X,d,f)$ be a TDS with the specification property. Then for any non-empty closed subset $C\subseteq M_{f}(X)$, we have
$\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
For every $\epsilon>0$, by Lemma 3.6 we have
$h_{top}^{B}(f,G^{C},\epsilon)\leq h_{top}^{B}(f,\{x\in X:V_{f}(x)\cap{C}\not=\emptyset\},\epsilon)\leq\sup_{\mu\in{C}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
By (3.17), there exists a constant $c>0$ such that for all $\mu\in C$ and any sufficiently small $\epsilon>0$,
$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu},\epsilon).$
Combining this inequality with the fact that $G_{\mu}\subset G^{C}$ for each $\mu\in C$, we have
$\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G^{C},\epsilon).$
Theorem 3.9.
Let $(X,d,f)$ be a TDS with the specification property. Let $C_{1}\subseteq M_{f}(X)$ be a non-empty set and $C_{1}\subseteq M_{f}(X)$. If $\overline{{\rm co}}(C_{1})$ is a connected component of $C_{2}$, then
$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C_{1}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
Since $C_{1}\subset\overline{{\rm co}}(C_{1})\subset C_{2}$, by Theorem 1.2 we have
$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm co}}(C_{1})},d\right)$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in\overline{{\rm co}}(C_{1})}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C_{1}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Furthermore, we have $\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)\leq\overline{\rm mdim}_{M}^{B}\left(f,^{C_{1}}G,d\right)$. Using Theorem 3.7, we get
$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C_{1}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
This completes the proof. ∎
The Theorems 3.7, 3.8 and 3.9 imply the following corollaries, which allows us to compare the metric mean dimensions of sup sets and sub sets with the saturated sets.
Corollary 3.10.
Let $(X,d,f)$ be a TDS with the specification property.
  • (1)
    If $C$ is a non-empty compact connected set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)=\overline{\rm mdim}_{M}^{B}\left(f,G_{C},d\right).$
  • (2)
    If $C$ is a non-empty compact convex set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right).$
Proof.
(1) follows from Theorems 1.2 and 3.7; (2) is due to Theorems 1.1 and 3.8. ∎
Corollary 3.11.
Let $(X,d,f)$ be a TDS with the specification property.
  • (1)
    If $C$ is a non-empty compact connected set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{B}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)\leq\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right).$
  • (2)
    If $C$ is a non-empty compact convex set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right)\leq\overline{\rm mdim}_{M}^{P}\left(f,^{C}G,d\right).$
  • (3)
    If $C$ is a non-empty compact connected set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{UC}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{UC}\left(f,G^{C},d\right)=\overline{\rm mdim}_{M}^{UC}\left(f,^{C}G,d\right).$
Proof.
$(1)$ is clear.
$(2)$. It follows from Theorem 1.1 that
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)=\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)\leq\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right).$
On the other hand, one has $h_{top}^{P}(f,G^{C},\epsilon)\leq\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ by Lemma 3.2. Hence, we have $\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right)$.
$(3)$. It follows from Theorem 1.4 and the relations $G_{C}\subset G^{C}$ and $G_{C}\subset^{C}G$. ∎
4 The applications of main results
In this section, we apply the main results to study the metric mean dimension of level sets, mean Li-Yorke chaos in infinite entropy systems and the metric mean dimension of the set of generic points of full shifts over compact metric spaces.
4.1 Metric mean dimensions of level sets
Let $(X,d,f)$ be a TDS and $\psi$ be a continuous function on $X$. If $\mu$ is a $f$-invariant probability measure on $X$, the Birkhoff ergodic theorem verifies that for $\mu$-a.e. $x\in X$, the limit
$\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)$
exists.
However, given $\alpha\in\mathbb{R}$ it does not give more information about the “size” of the set of the points of $X$ that the time average of $\psi$ equals $\alpha$. For this reason, the $\alpha$-level set of $\psi$ is defined by
$K_{\alpha}:=\{x\in X:\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha\}.$
In [BR23], the authors proved that if $K_{\alpha}$ is non-empty and the TDS $(X,d,f)$ has the specification property, then
$\overline{\rm mdim}_{M}^{B}(f,K_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\psi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where $M_{f}(\psi,\alpha):=\{\mu\in M_{f}(X):\int\psi d\mu=\alpha\}$.
Actually, we are able to prove this variational principle by an alternative approach and give new characterization for the (packing/upper capacity) metric mean dimension of level sets.
Theorem 4.1.
Let $(X,d,f)$ be a TDS with the specification property and $\psi$ be a continuous function on $X$. If $\alpha\in\mathbb{R}$ is a constant such that $K_{\alpha}\not$$\emptyset$, then
$\overline{\rm mdim}_{M}^{B}(f,K_{\alpha},d)$ $=\overline{\rm mdim}_{M}^{P}(f,K_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\psi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
$\overline{\rm mdim}_{M}^{UC}(f,K_{\alpha},d)$ $=\overline{\rm mdim}_{M}(f,X,d).$
The corresponding result are also valid for the three types of lower metric mean dimensions by changing $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.
Proof.
If $K_{\alpha}\not$$\emptyset$, then $M_{f}(\psi,\alpha)$ is a non-empty convex set. For $x\in X$, the statement $\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\psi,\alpha)$. Recall $G^{M_{f}(\varphi,\alpha)}=\{x\in X:V_{f}(x)\subset M_{f}(\varphi,\alpha)\}$. It follows that $K_{\alpha}=G^{M_{f}(\varphi,\alpha)}$. By Lemma 3.2 $(1)$, we get
$\overline{\rm mdim}_{M}^{P}(f,K_{\alpha},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (4.1)
By inequality (3.17), there exists a constant $c>0$ such that for every $\mu\in M_{f}(X)$ and sufficiently small $\epsilon>0$,
$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu},\epsilon).$
Since $G_{\mu}\subset K_{\alpha}$ for any $\mu\in M_{f}(\varphi,\alpha)$, we have
$\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,K_{\alpha},\epsilon),$
which implies that
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\overline{\rm mdim}_{M}^{B}(f,K_{\alpha},d).$ (4.2)
By (4.1), (4.2) and $\overline{\rm mdim}_{M}^{B}(f,K_{\alpha},d)\leq\overline{\rm mdim}_{M}^{P}(f,K_{\alpha},d)$, we get the first equalities.
Fix $\mu\in M_{f}(\psi,\alpha)$. Since the specification property implies $G_{\mu}\not$$\emptyset$ [PS07], we have $G_{\mu}\subset K_{\alpha}$. By Corollary 1.5, we obtain $\overline{\rm mdim}_{M}^{UC}(f,K_{\alpha},d)=\overline{\rm mdim}_{M}(f,X,d)$. ∎
Example 4.2.
Endow the product space $0,1]^{\mathbb{Z}}$ with the product metric:
$d((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}})=\sum_{n\in\mathbb{Z}}\frac{|x_{n}-y_{n}|}{2^{|n|}},$
and let $\sigma:[0,1]^{\mathbb{Z}}\to[0,1]^{\mathbb{Z}}$ be the left shift map defined by $\sigma((x_{n})_{n\in\mathbb{Z}})=(x_{n+1})_{n\in\mathbb{Z}}$.
It is shown that the system $([0,1]^{\mathbb{Z}},d,\sigma)$ has the specification property (cf. [DGS76, Proposition 21.2]), and ${\rm{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d)=1$ [LT18, E. Example]. Let $\varphi:[0,1]^{\mathbb{Z}}\to\mathbb{R}$ be a continuous map given by $\varphi((x_{n})_{n\in\mathbb{Z}})=x_{0}$. Then for any $\alpha\in[0,1]$, by Theorem 4.1 we have
$\overline{\rm mdim}_{M}^{B}(\sigma,K_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(\sigma,K_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{\sigma}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi),$
where $K_{\alpha}:=\{x\in[0,1]^{\mathbb{Z}}:\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}x_{j}=\alpha\}$, and $M_{\sigma}(\varphi,\alpha):=\{\mu\in M_{\sigma}([0,1]^{\mathbb{Z}}):\int\varphi d\mu=\alpha\}$. Furthermore, we have
$\overline{\rm mdim}_{M}^{UC}(\sigma,K_{\alpha},d)=1.$
4.2 Mean Li-Yorke chaos in infinite entropy systems
In this subsection, we provide a quantitative result involving the chaotic behavior of infinite entropy systems. Specifically, we prove that the (packing/upper capacity) metric mean dimension of the set of mean Li-Yorke pairs is equal to twice the metric mean dimension of the phase space.
Given a TDS $(X,d,f)$, this naturally induces a product system $(X\times X,d\times d,f\times f)$, where $X\times X$ is a compact space equipped with the product metric:
$d\times d((x_{1},y_{1}),(x_{2},y_{2}))=\max\{d(x_{1},x_{2}),d(y_{1},y_{2})\},$
and $(f\times f)(x,y):=(fx,fy)$. A pair $(x,y)\in X\times X$ is called a mean Li-Yorke pair if
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)=0,~\text{and}\ \limsup_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)>0.$
Namely, the orbits of $x$ and $y$ is sufficiently close to each other for sufficiently $n$, but they may be apart from each other for infinitely many time.
Denote by ${\rm MLY}(f)$ the set of all mean Li-Yorke pairs. A subset $S\subset X$ is called a mean Li-Yorke set if any two distinct points in $S$ form a mean Li-Yorke pair. The system $(X,f)$ is mean Li-Yorke chaotic if there is an uncountable mean Li-Yorke set. A well-known fact is that positive entropy systems are mean Li-Yorke chaotic [Dow14].
Theorem 4.3.
Let $(X,d,f)$ be a TDS with the specification property. Then for $S\in\{UC,P\}$, we have
$\overline{\rm mdim}_{M}^{S}(f\times f,{\rm MLY}(f),d\times d)=2~\overline{\rm mdim}_{M}(f,X,d)$
$\underline{\rm mdim}_{M}^{S}(f\times f,{\rm MLY}(f),d\times d)=2~\underline{\rm mdim}_{M}(f,X,d).$
Proof.
It suffices to prove the first equality since the second one can be obtained in a similar manner.
We claim that there exists $\mu_{0}\in M_{f\times f}(X\times X)$ such that
$\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}(x,y)>0.$
Notice that the specification property implies that the system has positive topology entropy. Since positive entropy systems are mean Li-Yorke chaotic [Dow14], we choose two distinct points $x,y\in X$, a sequence $\{n_{i}\}$ of positive integers that converges to $\infty$ as $i\to\infty$, and $\theta>0$ (depending on $x,y$) such that $\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta.$ Put $\mu_{n_{i}}:=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$. Then, in the weak-topology, any accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfies the desired claim.
Fix $\nu\in M_{f}(X)$ and put $\mu_{1}:=\nu\circ\phi^{-1}$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$. It is easy to see that $\mu_{1}\in M_{f\times f}(X\times X)$ and $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$. Fix $\gamma>0$ and let
$c_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d)-\gamma\}.$
By Theorem 2.4, we choose a sequence of positive real numbers $\epsilon_{k}\to 0$ as $k\to\infty$ and a sequence $\{m_{k}\}$ of $f\times f$-invariant measures such that for every $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)>c_{\gamma}.$
Let
$K=\overline{{\rm co}}\{\mu_{0},\mu_{1},m_{k},k\geq 1\}.$
Then $K$ is a closed convex subset of $M_{f\times f}(X\times X)$. Since $(X\times X,d\times dmf\times f)$ has the specification property, by [PS07] we have $G_{K}\not$$\emptyset$. For any $(x_{0},y_{0})\in G_{K}$, there exists two sequences of integers $\{t_{n}^{0}\}$ and $\{t_{n}^{1}\}$ such that $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$, respectively. Hence,
$\lim_{n\to\infty}\frac{1}{t_{n}^{0}}\sum_{i=0}^{t_{n}^{0}-1}d(f^{i}x_{0},f^{i}y_{0})=\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}>0,$
and
$\lim_{n\to\infty}\frac{1}{t_{n}^{1}}\sum_{i=0}^{t_{n}^{1}-1}d(f^{i}x_{0},f^{i}y_{0})=\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0.$
This yields that $G_{K}\subset{\rm MLY}(f)$. Therefore, by Theorem 1.1 one has
$\overline{\rm mdim}_{M}^{P}(f\times f,{\rm MLY}(f),d\times d)\geq$ $\overline{\rm mdim}_{M}^{P}(f\times f,G_{K},d\times d)$
$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\nu\in K}\inf_{{\rm diam}(\eta,d\times d)<\epsilon}h_{\nu}(f\times f,\eta)$
$\geq$ $\limsup_{k\to\infty}\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)\geq c_{\gamma}.$
Letting $\gamma\to 0$, one has
$\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d)\leq$ $\overline{\rm mdim}^{P}_{M}(f\times f,{\rm MLY}(f),d\times d)$
$\leq$ $\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)$
$\leq$ $\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d)$
$=$ $2~\overline{\rm mdim}_{M}(f,X,d).$
This completes the proof. ∎
For instance, under the setting of Example 4.2 we have
$\overline{\rm mdim}_{M}^{P}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=\overline{\rm mdim}_{M}^{UC}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=2.$
4.3 Generic points of full shift over compact metric space
For the full shifts over compact metric spaces, we present the precise formulae for the metric mean dimension of the set of generic points of invariant measures.
Let $(K,d_{K})$ be a compact metric space. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with the following product metric
$d\left((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}}\right)=\sum_{n\in\mathbb{Z}}\frac{d_{K}(x_{n},y_{n})}{2^{|n|}}.$
It is well-known that the system $(K^{\mathbb{Z}},d,\sigma)$ has the specification property. It was proved in [VV17, Theorem 5] that,
$\overline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d)$ $=\overline{\rm dim}_{B}(K,d_{K}),$ (4.3)
$\underline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d)$ $=\underline{\rm dim}_{B}(K,d_{K}),$ (4.4)
where $\overline{\rm dim}_{B}(K,d)$ and $\underline{\rm dim}_{B}(K,d)$ denote the upper and lower box dimensions of $K$ under the metric $d$, respectively.
By Corollary 1.5, for the full shift over compact metric spaces the metric mean dimensions of the set of generic points of invariant measures are fully determined by the box dimension of the underlying space and mean Rényi information dimension of invariant measures.
Theorem 4.4.
Let $(K^{\mathbb{Z}},d,\sigma)$ be the full shift over $(K,d_{K})$. Then for any $\mu\in M_{\sigma}(K^{\mathbb{Z}})$,
$\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K,d)$
$\overline{\rm mdim}_{M}^{B}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi).$
Next, for certain full shift over self-similar set we show the metric mean dimension of the product of the self-similar measure is closely related to the similarity dimension of IFS.
Let $m\geq 1$ be a positive integer and endow $\mathbb{R}^{m}$ with the supremum norm $\|\cdot\|_{\mathbb{R}^{m}}$. Assume that $T_{i}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ is contracting similitude for $1\leq i\leq l$, i.e., for some $r_{i}\in(0,1)$,
$\|T_{i}x-T_{i}y\|_{\mathbb{R}^{m}}=r_{i}\|x-y\|_{\mathbb{R}^{m}}.$
The iterated function system (IFS for short) $\{T_{i}\}_{1\leq i\leq l}$ satisfies the Open Set Condition if there exists a bounded open subset $V\subset\mathbb{R}^{m}$ such that $T_{i}(V)\subset V,\ \forall 1\leq i\leq l$, and $T_{i}(V)\cap T_{j}(V)=\emptyset~\text{for}~i\neq j.$ For instance, the middle-third Cantor set, the Sierpinński triangle or gasket, are the attractors of iterated function systems satisfying the Open Set Condition. Under the OSC condition, it is well known that there are (unique) compact subset $K$, which we call the self-similar set, and a measure $\mu$, which we call the self-similar measure, such that
  • (1)
    $K=\cup_{i=1}^{l}T_{i}(K)$;
  • (2)
    ${\rm dim}_{H}(K,d)={\rm dim}_{B}(K,d)=\alpha$, and $\alpha$ is the unique solution of equation
    $\sum_{i=1}^{l}r_{i}^{\alpha}=1,$
    where ${\rm dim}_{H}$ denotes the Hausdorff dimension under the metric $d:=\|\cdot\|_{\mathbb{R}^{m}}$. The root $\alpha$ is called the similarity dimension of the IFS;
  • (3)
    $\mu=\sum_{i=1}^{l}r_{i}^{\alpha}\cdot\mu\circ T_{i}^{-1}$ and ${\rm supp}(\mu)=K$.
Now take the attractor $K$ obtained as above and the product measure $\nu:=\mu^{\otimes\mathbb{Z}}$ on $K^{\mathbb{Z}}$.
Theorem 4.5.
The three types of metric mean dimensions of $G_{\nu}$ coincide with the similarity dimension. i.e.,
$\overline{\rm mdim}_{M}^{S}(\sigma,G_{\nu},d^{\mathbb{Z}})=\alpha=\limsup_{\epsilon\to 0}\frac{\inf_{{\rm diam}(\xi)<\epsilon}h_{\nu}(\sigma,\xi)}{|\log\epsilon|},$
where $S\in\{B,P,UC\}$.
Proof.
In [Hut81], Hutchinson proved that there exists a constant $C>0$ such that for any $z\in K$ and $\epsilon\in(0,1)$, one has
$\mu(B(z,\epsilon))\leq C\epsilon^{\alpha}.$ (4.5)
Fix $\epsilon\in(0,1)$. For any $x=(x_{k})_{k\in\mathbb{Z}}\in K^{\mathbb{Z}}$, we define
$L_{n}(x,\epsilon)=\left\{(y_{k})_{k\in\mathbb{Z}}\in K^{\mathbb{Z}}:\|x_{k}-y_{k}\|_{\mathbb{R}^{m}}<\epsilon,\ 0\leq k\leq n-1\right\}.$
Then $B_{n}(x,\epsilon)\subset L_{n}(x,\epsilon)$. Since $\nu(L_{n}(x,\epsilon))\leq C^{n}\epsilon^{n\alpha}$ by (4.5), one has
$\nu(B_{n}(x,\epsilon))\leq C^{n}\epsilon^{n\alpha},\ \forall x\in K^{\mathbb{Z}},$
which yields that the Brin-Katok local $\epsilon$-entropy of $\nu$ is bounded blew by
$\overline{h}_{\nu}^{BK}(\sigma,\epsilon):=\int\limsup_{n\to\infty}\frac{-\log\mu(B_{n}(x,\epsilon))}{n}d\nu\geq\alpha|\log\epsilon|-\log C.$
Applying the Shannon-McMillan-Breiman theorem, the following inequality is valid for any TDS $(X,d,f)$:
$\overline{h}_{\mu}^{BK}(f,\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),\ \forall\mu\in M_{f}(X).$ (4.6)
Hence, we have
[…]
This implies the desired result. ∎
Acknowledgments: The first and second authors were supported by the National Natural Science Foundation of China (No. 12101340). The third author was supported by the China Postdoctoral Science Foundation (No. 2024M763856) and the Postdoctoral Fellowship Program of CPSF (No. GZC20252040); Rui Yang would like to thank Dr. Yunping Wang for many helpful discussions and the hospitality extended to him during his visit to Ningbo University. Special thanks also go to Prof. Ercai Chen for his many useful comments.